### v1

# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

Abstract
Information from related source studies can often enhance the findings of a target study. However, the distribution shift between target and source studies can severely impact the efficiency of knowledge transfer. In the high-dimensional regression setting, existing transfer approaches mainly focus on the parameter shift. In this paper, we focus on the high-dimensional quantile regression with knowledge transfer under three types of distribution shift: parameter shift, covariate shift, and residual shift. We propose a novel transferable set and a new transfer framework to address the above three discrepancies. Non-asymptotic estimation error bounds and source detection consistency are established to validate the availability and superiority of our method in the presence of distribution shift. Additionally, an orthogonal debiased approach is proposed for statistical inference with knowledge transfer, leading to sharper asymptotic results. Extensive simulation results as well as real data applications further demonstrate the effectiveness of our proposed procedure.
Keywords: transfer learning; high-dimensional quantile regression; distribution shift; transferable set; orthogonal debiasing
1 Introduction
Previous experiences can offer valuable insights for learning new tasks. Transfer learning, the technique of improving the performance of target learners on target domains by leveraging knowledge from different but related source domains (Zhuang et al.,, 2021), has shown significant success across various application scenarios. These applications range from machine learning problems such as natural language processing (Ruder et al.,, 2019) and computer vision (Shao et al.,, 2015), to science problems such as protein representation (Fenoy et al.,, 2022) and drug discovery (Turki et al.,, 2017), owing to the increasing richness of data sources. However, our understanding of transfer learning from a statistical perspective remains incomplete, with the fundamental challenge being the detection and utilization of valuable information within heterogeneous source studies.
In this paper, we focus on the transfer learning for quantile regression in the context of high-dimensional setting. Quantile regression (QR) models the impact of covariates on the conditional response distribution and has achieved great success in many applications. For QR in various areas, see Koenker, (2017) for a detailed review. One of QR’s most appealing features is its ability to describe the impact of covariates not only on the mean but also on the tails of the response distribution with robustness to heteroscedastic errors. Given potential non-negligible divergence between target and source distributions, QR can provide a more comprehensive assessment of what can be transferred from the source studies. Moreover, we focus on the regime of high-dimensional models, since the error bounds of high-dimensional estimators are heavily limited by the sample size. This motivates us to explore multiple external datasets to improve the performance of high-dimensional estimators on the target task.
To set up the framework, suppose we have a target dataset $((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}$ and the parameter of interest is the quantile of the response $y_{i}^{(0)}$ conditional on covariates $\boldsymbol{x}_{i}^{(0)}\in\mathbb{R}^{p}$ at a given quantile level $\tau\in(0,1)$, denoted by $\boldsymbol{\beta^{*}}$ (sometimes we use $\boldsymbol{w}^{(0)}$ for unity). Also, we have $K$ independent source datasets $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$, where the corresponding quantile coefficients are denoted by $\boldsymbol{w}^{(k)},k=1,\dots,K$. Denote the positive covariance matrix $\boldsymbol{\Sigma}^{(k)}:=\mathbb{E}[\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]$ for each study. In the high-dimensional setting, the number of covariates $p$ is very large, possibly larger than the overall sample size $N:=\sum_{k=0}^{K}n_{k}$. Assume a linear QR model, which is
$y_{i}^{(k)}=(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{w}^{(k)}+\epsilon_{i}^{(k)},\qquad i=1,\dots,n_{k},\quad k=0,\dots,K,$ (1)
where each model residual $\epsilon_{i}^{(k)}$ satisfies $\mathbb{P}(\epsilon_{i}^{(k)}\leq 0|\boldsymbol{x}_{i}^{(k)})=\tau$ with conditional density function $f^{(k)}(\cdot|\boldsymbol{x}_{i}^{(k)})$. Assume the target parameter $\boldsymbol{\beta^{*}}$ is sparse, where the number of non-zero coefficients satisfies $\|\boldsymbol{\beta^{*}}\|_{0}=s\ll\min\{n_{0},p\}$. Also, we define the contrast vector $\boldsymbol{\delta}^{(k)}:=\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$ (obviously $\boldsymbol{\delta}^{(0)}=0$). Our goal is to estimate the target parameter $\boldsymbol{\beta^{*}}$, with the help of abundant external data.
Three types of distribution shift. For the QR model (1) specified on all studies, there may exist three types of distribution shifts between the target and source domains:
  1. 1.
    parameter shift: $\boldsymbol{\beta^{*}}\neq\boldsymbol{w}^{(k)}$, i.e., the model coefficients differ.
  2. 2.
    residual shift: $P(\epsilon_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})\neq P(\epsilon_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$, i.e., the conditional distributions of model residuals differ.
  3. 3.
    covariate shift: $P(\boldsymbol{x}_{i}^{(0)})\neq P(\boldsymbol{x}_{i}^{(k)})$, i.e., the marginal covariate distributions differ.
Both parameter shift and residual shift can be seen as the posterior/label shift, meaning that the conditional distribution of responses differs between $P(y_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})$ and $P(y_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$. Many examples have revealed that transferring source studies under distribution shift may pose challenges to the final performance and lead to the negative transfer (Wang et al.,, 2019; Zhang et al.,, 2022). Hence, the core issue of this paper is:
How to identify the transferable information from different source studies under distribution (covariate/parameter/residual) shift? How to conduct valid prediction and inference for high-dimensional quantile regression with knowledge transfer?
Methodologically, transfer learning falls within the field of multi-source data analysis, a prominent topic in contemporary statistics. Numerous approaches have been developed for multi-source settings, varying by task type, data category, and other factors. These include federated learning (McMahan et al.,, 2017), distributed learning (Jordan et al.,, 2019; Fan et al.,, 2023), fusion learning (Shen et al.,, 2020; Li and Luedtke,, 2023), multi-task learning (Duan and Wang,, 2023; Knight and Duan,, 2023), and transfer learning (Pan and Yang,, 2009; Zhuang et al.,, 2021), each with its unique characteristics and structures. For transfer learning, it aims to leverage information from source domains to improve the performance of the target task, with the core challenge being how to extract valuable source information while addressing data heterogeneity. Transfer learning has been studied across a variety of contexts, including linear regression (Li et al.,, 2022), generalized linear regression (Tian and Feng,, 2023), nonparametric classification (Cai and Wei,, 2021), nonparametric regression (Cai and Pu,, 2024), graphical models (Li et al.,, 2023), and causal inference (Wei et al.,, 2023).
The main focus of this paper lies on the transfer learning for high-dimensional linear models, where the goal is estimation and inference for the target parameter. For parameter estimation, a seminal and widely adopted approach is the two-step framework introduced in Li et al., (2022). In this method, the oracle transferable set $A_{h}:=\{\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h\}$ is defined, where source studies with a small contrast of parameters are considered informative. The two-step procedure first combines the data from $A_{h}$ to obtain a pooled estimate, then debiases this estimate using the target data. Tian and Feng, (2023) then extended this approach to generalized linear regression, providing a consistent source detection algorithm based on cross-validated loss comparison. Huang et al., (2023) and Zhang and Zhu, (2023) further developed direct extensions of the two-step framework and source detection procedure for high-dimensional quantile regression and smoothed quantile regression, respectively. Among the limited literature for statistical inference with knowledge transfer, Tian and Feng, (2023), Li et al., (2024), and Huang et al., (2023) applied the one-step correction technique (Zhang and Zhang,, 2014; van de Geer et al.,, 2014) to debias the previous transfer estimate on the target study.
As defined by the transferable set $A_{h}$, the two-step approach only addresses parameter shift, leaving it susceptible to covariate and residual shifts across multiple sources. However, studies in machine learning have shown that it is essential to account for both covariate and label shifts in heterogeneous data settings, such as image recognition Park et al., (2023) and disease diagnosis Chen et al., (2022). Additionally, due to these shifts, prior debiased transfer estimators achieved only $\sqrt{n_{0}}$-normality, providing no improvement over the original target baseline. To reduce the impact of covariate shift, Li et al., (2024) proposed jointly estimating the $\boldsymbol{w}^{(k)}$ of each study using $\ell_{1}$-minimization, Zhao et al., (2023) utilized a residual importance-weighted algorithm to make better use of source observations, and He et al., (2024) added a fused-regularizer to the pooling step of the two-step approach. While these works addressed covariate shift and introduced remedies, none of them account for residual shift. However, several studies suggest that addressing complex label shift issues requires a distributional approach, such as density estimation (Lipton et al.,, 2018; Garg et al.,, 2020; Lee et al.,, 2024), rather than simply assuming a linear relationship and comparing coefficients. In our context of quantile regression, the distributional information around each quantile differs across sources, making it essential to handle residual shift to accurately capture the information at each quantile of the target response. To demonstrate how severe residual shift can lead to negative transfer, we provide a motivating example through simulations of existing transfer methods.
The motivating example follows a similar setup to the simulation in Zhang and Zhu, (2023). We set $p=500,s=10,n_{0}=200,\boldsymbol{\beta^{*}}=(\mathbf{0.5}_{s},\mathbf{0}_{p-s})^{\top},\boldsymbol{x}_{i}^{(0)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$ with $\boldsymbol{\Sigma}_{\boldsymbol{x}}=(0.7^{|i-j|})_{1\leq i,j\leq p}$. We consider only one source study with sample size $n_{1}$ varying from 100 to 500, and we assume no parameter or covariate shift, i.e., $\boldsymbol{\delta}^{(1)}=0,\boldsymbol{x}_{i}^{(1)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$. The key change lies in the residual distributions, where $\epsilon_{i}^{(0)}$ is standard normal, but $\epsilon_{i}^{(1)}$ follows one of the four types of distributions: 1) standard normal: $\mathcal{N}(0,1)$; 2) Cauchy distribution: $\mathcal{C}(0,3)$; 3) mixed Gaussian: $z\mathcal{N}(-3,0.5)+(1-z)\mathcal{N}(3,0.5)$, $z\sim\mathcal{B}ernoulli(\tau)$; 4) noisy normal: $\mathcal{N}(0,5^{2})$. All residuals are generated with necessary translational adjustments for the $\tau$-th quantile at zero.
It is worth noting that in this setting, the source study is included in the oracle transferable set by the previous definition (Li et al.,, 2022; Tian and Feng,, 2023). We compare four methods for estimating the target parameter $\boldsymbol{\beta^{*}}$ at $0.2$-th quantile, which are: 1) L1-QR: $\ell_{1}$-penalized quantile regression (Belloni and Chernozhukov,, 2011) on the target data; 2) Pooling: $\ell_{1}$-penalized quantile regression on the entire dataset; 3) TransQR_2step: two-step transfer for quantile regression (Huang et al.,, 2023); 4) TransSQR_2step: two-step transfer for smoothed quantile regression (Zhang and Zhu,, 2023). We repeat these methods 50 times and report the average of $\ell_{2}$-estimation errors in Figure 1.1.
Figure 1.1: Average $\ell_{2}$-estimation errors of four methods under residual shift
As shown in Figure 1.1, it is clear that, apart from the normal setting which demonstrates the anticipated transfer benefit, all other settings reveal negative transfer effects. Pooling, TransQR_2step, and TransSQR_2step perform significantly worse than the baseline L1-QR on the target data, strongly suggesting that assessing information transferability based solely on parameter similarity is insufficient in cases involving residual shift. This observation raises an interesting question: how can we effectively measure and control residual shift? Note that the adverse impact of residual shift fundamentally arises from information loss around the target quantile. So, what constitutes a meaningful representation of local distributional information? The answer is density. Motivated by this, we propose a novel transferable set:
$\mathcal{C}_{h}=\{k:1\leq k\leq K,\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\mathbb{E}[f^{(0)}(0|\boldsymbol{x}_{i}^{(0)})]\leq h_{2}\mathbb{E}[f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})]\},$
where $h_{1},h_{2}$ represent oracle yet unknown transferability levels. This set is designed to filter out source studies characterized by large parameter discrepancies and small densities at the $\tau$-th quantile, thus enabling the simultaneous management of parameter and residual shifts. We will further elaborate on the rationale behind the construction in Section 2. Using this transferable set, we develop a whole transfer learning framework to address parameter, covariate, and residual shifts concurrently within the high-dimensional quantile regression domain. Our contributions are:
  • We introduce the concept of residual shift in the multi-source setting and define a new transferable set to handle both parameter and residual shifts, offering deeper insights into managing distribution shift under knowledge transfer. For parameter estimation, we extend the constrained $\ell_{1}$-minimization algorithm (Li et al.,, 2024) to quantile regression, which remains robust in the presence of covariate shift. Furthermore, we develop a corresponding source detection procedure and integrate it into a unified one-shot estimation algorithm.
  • In theory, we derive non-asymptotic $\ell_{1}/\ell_{2}$-error bounds for our transfer estimators, broadening existing results to account for multiple types of distribution shifts, not just parameter shift. The resulting convergence rate surpasses that of single-task error bounds (Belloni and Chernozhukov,, 2011; Wang and He,, 2024), underscoring the advantage of our approach, especially when divergent sources are involved. The detection consistency of our source screening procedure is also established.
  • A debiased transfer approach based on Neyman’s orthogonality (Belloni et al.,, 2019) is proposed for statistical inference on quantile coefficients with knowledge transfer. Our method avoids the estimation of high-dimensional precision matrices and allows non-identical data distributions, which helps to alleviate the trouble by covariate and residual shift under knowledge transfer. To the best of our knowledge, our method is the first to debias on informative sources to enhance inference efficiency. By integrating valuable source studies, our approach achieves sharper asymptotic rate than previous results, leading to more efficient debiased estimates and shorter confidence intervals.
The rest of the article is organized as follows. In Section 2, we propose our transfer learning approach under distribution shift. We will introduce our transferable set and transfer framework. Non-asymptotic error bounds are revealed subsequently. Section 3 describes the source detection procedure, accompanied by consistency guarantees. In Section 4, we present the orthogonal debiased approach for statistical inference with knowledge transfer, along with the corresponding asymptotic theory. Section 5 reports simulation results on parameter estimation and statistical inference, highlighting the advantages of our methods. Finally, in Section 6 we apply our methods to the Genotype-Tissue Expression (GTEx) dataset, predicting brain tissue gene expression levels by leveraging information from other tissues.
2 Transfer Learning under Distribution Shift
We begin this section with some necessary notations. For two sequences of positive numbers $\left\{a_{n}\right\}$ and $\left\{b_{n}\right\}$, we write $a_{n}\lesssim b_{n}$ if $a_{n}\leq cb_{n}$ for some universal constant $c\in(0,\infty)$, and $a_{n}\gtrsim b_{n}$ if $a_{n}\geq c^{\prime}b_{n}$ for some universal constant $c^{\prime}\in(0,\infty)$. We say $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. For any $q\in[0,\infty]$ and vector $\boldsymbol{x}=(x_{1},\cdots,x_{p})\in\mathbb{R}^{p}$, we write $\|\boldsymbol{x}\|_{q}$ for its $l_{q}$-norm. Let $a\wedge b$ and $a\vee b$ denote the minimum and maximum of $a$ and $b$, respectively. For a vector $\boldsymbol{v}\in\mathbb{R}^{p}$ and a subset $S\subseteq\{1,2,\dots,p\}$, we use $\boldsymbol{v}_{S}$ to denote the restriction of vector $\boldsymbol{v}$ to the index set $S$. The sub-Gaussian norm of a random variable $u\in\mathbb{R}$ is $\|u\|_{\psi_{2}}=\inf\{t>0:\mathbb{E}[\exp(u^{2}/t^{2})]\leq 2\}$ and the sub-Gaussian norm of a random vector $\boldsymbol{U}\in\mathbb{R}^{n}$ is $\|\boldsymbol{U}\|_{\psi_{2}}=\sup_{\|\boldsymbol{v}\|_{2}=1,\boldsymbol{v}\in\mathbb{R}^{n}}\|\langle\boldsymbol{U},\boldsymbol{v}\rangle\|_{\psi_{2}}$. The sub-exponential norm of a random variable $u\in\mathbb{R}$ is $\|u\|_{\psi_{1}}=\inf\{t>0:\mathbb{E}[|u|/t]\leq 2\}$ and the sub-exponential norm of a random vector $\boldsymbol{U}\in\mathbb{R}^{n}$ is $\|\boldsymbol{U}\|_{\psi_{1}}=\sup_{\|\boldsymbol{v}\|_{2}=1,\boldsymbol{v}\in\mathbb{R}^{n}}\|\langle\boldsymbol{U},\boldsymbol{v}\rangle\|_{\psi_{1}}$. The expectation symbol $\mathbb{E}[(\cdot)_{i}^{(k)}]:=\mathbb{E}_{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})}[(\cdot)_{i}^{(k)}]$ means obtaining expectation with the joint distribution $P(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})$.
2.1 Methodology of Tackling Three Types of Distribution Shift
Parameter Shift. Naturally we should consider a source study as informative if its parameter contrast $\boldsymbol{\delta}^{(k)}$ is relatively small. We choose the $\ell_{1}$-sparsity to quantify the informative level, as in many practical applications, parameter shift typically occurs across multiple dimensions without the overall magnitude growing too fast. Formally, the screening criterion is to reject source studies with $\|\boldsymbol{\delta}^{(k)}\|_{1}$ larger than an oracle transferring level, which has been widely adopted in the previous transfer work (Li et al.,, 2022; Tian and Feng,, 2023).
Residual Shift. Let’s start from the simplest low-dimensional linear model to understand how to tackle the residual shift. Suppose the target and source task as
$\left\{\begin{array}[]{l}\text{target model: }~{}y_{i}^{(0)}=(\boldsymbol{x}_{i}^{(0)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(0)},~{}i=1,\dots,n,~{}\epsilon_{i}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2}),\\ \text{source model: }~{}y_{i}^{(1)}=(\boldsymbol{x}_{i}^{(1)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(1)},~{}i=1,\dots,n,~{}\epsilon_{i}^{(1)}\sim\mathcal{N}(0,\sigma_{1}^{2}).\end{array}\right.$
Denote the covariance matrix $\Sigma_{x}:=\mathbb{E}[\boldsymbol{x}_{i}^{(0)}(\boldsymbol{x}_{i}^{(0)})^{\top}]=\mathbb{E}[\boldsymbol{x}_{i}^{(1)}(\boldsymbol{x}_{i}^{(1)})^{\top}]$. Based on the knowledge of linear models, the target and pooling OLS estimator $\widehat{\boldsymbol{\beta}}_{tar},\widehat{\boldsymbol{\beta}}_{pool}$ satisfy:
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{\sigma_{0}^{2}}{n}\Sigma_{x}^{-1}\right),\quad\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{\sigma_{0}^{2}+\sigma_{1}^{2}}{4n}\Sigma_{x}^{-1}\right).$ (2)
If $\sigma_{1}^{2}$ is much larger than $\sigma_{0}^{2}$, i.e., the source study is much more noisy, then the pooling estimator will be much less efficient compared to the target one. From this perspective, addressing the residual shift is actually to avoid the increasing noise caused by incorporating divergent source studies.
Moving to the low-dimensional quantile regression with the same setting, we utilize the traditional quantile loss $\rho_{\tau}(x):=x(\tau-\mathbb{I}(x\leq 0))$ on the target data and the pooling data respectively. Based on the large sample theory of $M$-estimators (Van der Vaart,, 2000) and necessary regular conditions, it is easy to verify that
[…]
where $f^{(0)}(\cdot),f^{(1)}(\cdot)$ are density functions of residuals $\epsilon_{i}^{(0)},\epsilon_{i}^{(1)}$ respectively. When $f_{1}(0)\leq(\sqrt{2}-1)f_{0}(0)$, $\widehat{\boldsymbol{\beta}}_{pool}$ is less efficient than $\widehat{\boldsymbol{\beta}}_{tar}$. It can be seen that the residual density values $f^{(0)}(0)$ and $f^{(1)}(0)$ serve a similar role for QR estimators as residual variances $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ do for OLS estimators, both indicating the quality of information available for estimation. In fact, this relationship was evident in our motivating example, where unreliable source studies had considerably lower density values around the $\tau$-th quantile compared to the target study. When a source study has a very low density near the $\tau$-th quantile, its QR estimator often performs poorly due to the noisy information, making it unlikely that transferring from such a study would be beneficial. With this in mind, we propose another screening criterion to discard source studies with residual densities below a specified oracle threshold, as they are unlikely to contribute valuable information beyond introducing noise. For simplicity, we denote $f^{(k)}_{i}:=f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$ as the conditional density of residual $\epsilon_{i}^{(k)}$ at zero, hereinafter referred to as the $\tau$-th density.
Remark 1.
Rejecting sources with lower $\tau$-th density $f^{(k)}_{i}$, i.e., leveraging sources with higher $\tau$-th density, can also be explained from the linear regression example (2). If a source study has a negligible error term, i.e., $\sigma_{1}^{2}=0$, then it means accurately identifying the target parameter, leading to a wonderful transferable study. That is exactly why we choose this measure in the transferable set to address the residual shift. Additionally, with balanced study sizes, minimizing the estimator’s mean square error (MSE) benefits from a smaller contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ to reduce bias, while a larger $\tau$-th density helps to control variance. We will illustrate this point through theoretical results later in this section.
Transferable Set. From the analysis above, we need to control the transferring level for both parameter contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and $\tau$-th density $f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$, in order to achieve efficient knowledge transfer in the presence of parameter and residual shift. Hence we propose the oracle transferable set $\mathcal{C}_{h}$ as defined in the introduction:
$\mathcal{C}_{h}=\left\{k:\quad 1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{\mathbb{E}[f^{(0)}_{i}]}{\mathbb{E}[f^{(k)}_{i}]}\leq h_{2}\right\},$ (3)
where $h_{1},h_{2}$ are oracle transferring levels revealing the magnitude of parameter and residual shifts respectively. One level $h_{1}$ controls the parameter similarity, while the other level $h_{2}$ regulates the expectation of source $\tau$-th density $\mathbb{E}[f^{(k)}_{i}]$ compared to the target $\mathbb{E}[f^{(0)}_{i}]$, facilitating the exclusion of heterogeneous source studies. Both requirements in the new transferable set (3) are actually elucidated in the earlier text. To guarantee the improvement of transfer learning, we only leverage the information from source studies included in this set $\mathcal{C}_{h}$. Strictly speaking, both transferring levels should be related to the sample size, since the impact caused by source studies differs with varying sample sizes under the same degree of shift. However, the impact from sample sizes is too hard to derive, where each level’s optimal choice depends on all target and source studies’ characteristics, as shown in the theoretical analysis later. Thus for simplicity, we set two universal oracle levels in (3) with implicitly assuming balanced sample sizes.
Covariate Shift. The negative effect of covariate shift is mainly brought by the divergence of covariance matrices $\boldsymbol{\Sigma}^{(k)}$ between target and source studies. As mentioned by He et al., (2024), in the two-step transfer approach with the transferable set $\mathcal{C}_{h}$, the pooling estimator in the first step comes with an asymptotic bias $\boldsymbol{\delta}_{Pool}:=(\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)})^{-1}\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)}\boldsymbol{\delta}^{(k)}$. This bias can make the contrast $\boldsymbol{\delta}^{(k)}$ amplified by the factor $C_{\boldsymbol{\Sigma}}:=1+|\mathcal{C}_{h}|\max_{k\in\mathcal{C}_{h}}\sup_{1\leq j\leq p}\|e_{j}^{\top}(\boldsymbol{\Sigma}^{(k)}-\boldsymbol{\Sigma}^{(0)})(\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)})^{-1}\|_{1}$, which may diverge when $\boldsymbol{\Sigma}^{(k)}$s are dissimilar. To address this problem, here we generalize the constrained optimization algorithm proposed by Li et al., (2024) to the field of quantile regression, for jointly estimating the target parameter $\boldsymbol{\beta^{*}}$ and contrast vectors $\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}$. Our transfer framework is formalized as follows:
$\left(\widehat{\boldsymbol{\beta}},\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}\right)=\underset{\boldsymbol{\beta},\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}}{\arg\min}\lambda_{\boldsymbol{\beta}}\left\|\boldsymbol{\beta}\right\|_{1}+\sum_{k\in\mathcal{C}_{h}}\lambda_{k}\|\boldsymbol{\delta}^{(k)}\|_{1},$ (4)
subject to $\text{subject to }\left\{\begin{array}[]{l}\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k},\quad\text{for }k\in\{0\}\cup\mathcal{C}_{h},\\ \|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}},\end{array}\right.$
where $S_{n_{k}}(\boldsymbol{\beta}):=\partial L^{(k)}(\boldsymbol{\beta})/\partial\boldsymbol{\beta}$ is the subgradient of the quantile loss function $L^{(k)}(\boldsymbol{\beta}):=\sum_{i=1}^{n_{k}}\rho_{\tau}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{\beta})$ for each study $k\in\{0\}\cup\mathcal{C}_{h}$, $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$ are tuning parameters. It is important to note that the $\ell_{1}$-sparsity regularizer for each contrast $\boldsymbol{\delta}^{(k)}$ allows us to be free from the similarity restriction of covariance matrices, hence removing the $C_{\boldsymbol{\Sigma}}$ factor and achieving robustness under covariance heterogeneity (Li et al.,, 2024; He et al.,, 2024). We will further reveal this property in the subsequent theoretical analysis.
Remark 2.
Here we give some explanations for our transfer framework. The objective function in (4) encourages sparse solutions of $\widehat{\boldsymbol{\beta}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}$. The constraint $\|S_{n_{0}}(\boldsymbol{\beta})\|_{\infty}\leq\lambda_{0}$ is inherited from the target model, imposing that $\boldsymbol{\beta}$ should be identified as the true parameter $\boldsymbol{\beta^{*}}$ in the target task. The constraint $\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k}$ comes from the $k$-th source model, imposing that $\boldsymbol{\delta}^{(k)}$ should be identified as the true contrast $\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$. The last constraint $\|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}}$ aggregates target study and studies in the oracle transferable set $\mathcal{C}_{h}$, revealing that our estimation leverages information from valuable sources. In contrast to the two-step transfer framework (Li et al.,, 2022; Tian and Feng,, 2023), our algorithm can simultaneously estimate $\boldsymbol{\beta^{*}}$ and $\boldsymbol{\delta}^{(k)}$, providing more information for the downstream inference task.
Remark 3.
Zhao et al., (2023) proposed a residual importance weighted transfer approach to tackle distribution shift for high-dimensional linear regression. Specifically, they set the oracle weights as $\omega_{k}=f^{(0)}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{\beta^{*}})~{}/~{}f^{(k)}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\boldsymbol{\beta^{*}}+\boldsymbol{\delta}^{(k)}))$, then conduct the weighted LASSO to correctly identify $\boldsymbol{\beta^{*}}$ from the pooled data. Such weighting construction may still lose efficacy due to residual shift. If the source is abnormal like the motivating example, then the corresponding weight will approach infinity. However, we have shown that this source study ought to be discarded. Therefore, their approach may not be applicable to scenarios involving residual shift in our context.
To solve the constrained $\ell_{1}$-minimization problem (4), we approximate (4) with a joint $\ell_{1}$-penalized quantile regression ($\ell_{1}$-QR) form, leveraging the equivalence between Lasso and Dantzig selectors Candes and Tao, (2007); Bickel et al., (2009):
$\left(\widehat{\boldsymbol{\beta}},\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}\right)=\underset{\boldsymbol{\beta},\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}}{\arg\min}\left\{\sum_{k\in\{0\}\cup\mathcal{C}_{h}}L^{(k)}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})+\lambda_{\boldsymbol{\beta}}\|\boldsymbol{\beta}\|_{1}+\sum_{k\in\mathcal{C}_{h}}\lambda_{k}\|\boldsymbol{\delta}^{(k)}\|_{1}\right\}.$ (5)
This formulation is similar to the co-training step in He et al., (2024), with replacing $\boldsymbol{\delta}^{(k)}$ by $\boldsymbol{w}^{(k)}-\boldsymbol{\beta}$. As suggested by Li et al., (2024), the algorithm for solving (5) proceeds naturally using iterative alternating direction methods. At a high level, our method works as follows. We first obtain an initial estimator $\widehat{\boldsymbol{\beta}}_{0}$ by $\ell_{1}$-QR solely on the target data (or the estimator $\widehat{\boldsymbol{\beta}}_{t-1}$ from the last iteration). Then for each source $k\in\mathcal{C}_{h}$ we subtract $(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}_{t-1}$ from the response $y_{i}^{(k)}$ and obtain the estimator $\widehat{\boldsymbol{\delta}}^{(k)}_{t}$ by $\ell_{1}$-QR on the refined source dataset. After that, we take $\{\widehat{\boldsymbol{\delta}}^{(k)}_{t}\}_{k\in\mathcal{C}_{h}}$ into the objective function (5), solve for the estimator $\widehat{\boldsymbol{\beta}}_{t}$, and begin the next round. The iteration stops when estimators converge, where the detailed procedure is summarized in Algorithm 1. We defer the practical choice of tuning parameters $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$ to the simulation section.
Input: Target data $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\mathcal{C}_{h}}$, tuning parameters $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$, iteration step $T$.
Compute the initial estimate $\widehat{\boldsymbol{\beta}}_{0}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}},\lambda_{0}]$.
for $t=1:T$ do
  1. 1.
    Compute $\widehat{\boldsymbol{\delta}}^{(k)}_{t}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}_{t-1})_{i=1}^{n_{k}},\lambda_{k}],k\in\mathcal{C}_{h}$;
  2. 2.
    Solve the pooled $\ell_{1}$-QR problem
    $\widehat{\boldsymbol{\beta}}_{t}=\underset{\boldsymbol{\beta}}{\arg\min}\left\{\sum_{k\in\{0\}\cup\mathcal{C}_{h}}L^{(k)}(\boldsymbol{\beta}+\widehat{\boldsymbol{\delta}}^{(k)}_{t})+\lambda_{\boldsymbol{\beta}}\left\|\boldsymbol{\beta}\right\|_{1}\right\}.$
Output: $\{\widehat{\boldsymbol{\beta}},\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}\}:=\{\widehat{\boldsymbol{\beta}}_{T},\{\widehat{\boldsymbol{\delta}}^{(k)}_{T}\}_{k\in\mathcal{C}_{h}}\}$.
Algorithm 1 Parameter estimation with oracle transferable set $\mathcal{C}_{h}$
2.2 Theory on Estimation Error
Now we study the non-asymptotic convergence rate of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ obtained by the transfer approach (4) with transferable set $\mathcal{C}$, where $\mathcal{C}$ is an arbitrary prespecified set for theoretical generality. Under the high-dimensional sparse setting, we characterize the geometric structure of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\}_{k=1}^{K}$ through the restrictive sets (Belloni and Chernozhukov,, 2011; Negahban et al.,, 2012; Wang and He,, 2024), which are denoted by
for $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ with $\ell_{0}$-sparsity: $\Gamma_{H}=\left\{v\in\mathbb{R}^{p}:\left\|v_{S^{c}}\right\|_{1}\leq\left\|v_{S}\right\|_{1}\right\},$
for $\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}$ with $\ell_{1}$-sparsity: $\Gamma_{W}^{(k)}=\{v\in\mathbb{R}^{p}:\|v_{S_{a}^{(k)c}}\|_{1}\leq\|v_{S_{a}^{(k)}}\|_{1}+2\|\boldsymbol{\delta}^{(k)}_{S_{a}^{(k)c}}\|_{1}\},$
where $S=\{j:\boldsymbol{\beta}^{*}_{j}\neq 0,1\leq j\leq p\},S_{a}^{(k)}=\{j:|\boldsymbol{\delta}^{(k)}_{j}|>a,1\leq j\leq p\}$ with a threshold $a>0$. The sets $\{\Gamma_{W}^{(k)}\}_{k=1}^{K}$ depends on $a$, but we omit the dependence in notation for simplicity. $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\}_{k=1}^{K}$ will be shown to lie in the sets $\Gamma_{H}$ and $\{\Gamma_{W}^{(k)}\}_{k=1}^{K}$ with high probability respectively in the supplementary materials. Conditions below constitute a set of basic assumptions for establishing the statistical properties, where the index $k$ refers to each study $k\in\{0,\dots,K\}$.
Condition 1.
The conditional density value at the $\tau$-th quantile $f_{i}^{(k)}=f_{i}^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$ is positive, continuously differentiable, and uniformly upper bounded. Also, there exist a universal positive constant $b_{0}$ such that the derivatives $f_{i}^{(k)}\text{}^{\prime}(t|\boldsymbol{x}_{i}^{(k)})$ exist and are uniformly bounded for $|t|\leq b_{0}$.
Condition 2.
There exist universal positive constants $m_{1},m_{2},m_{u}$ such that
$\min\left\{\frac{v^{T}\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v}{\|v\|_{2}^{2}},~{}\frac{\|\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v\|_{\infty}}{\|v\|_{\infty}}\right\}$ $\geq m_{1}\mathbb{E}[f_{i}^{(k)}],$
$\max\left\{\frac{v^{T}\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v}{\|v\|_{2}^{2}},~{}\frac{\|\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v\|_{\infty}}{\|v\|_{\infty}}\right\}$ $\leq m_{2}\mathbb{E}[f_{i}^{(k)}],$
for any non-zero vector $v\in\Gamma_{W}^{(k)}$ ($v\in\Gamma_{H}$ when $k=0$). Also, the spectral norm of covariance matrix satisfies $\lambda_{\max}(\Sigma^{(k)})\leq m_{u}$.
Condition 3.
Each covariate $\boldsymbol{x}_{i}^{(k)}$ is a mean-zero sub-Gaussian random vector with sub-Gaussian norm uniformly bounded by a positive constant $m_{g}$.
Condition 4.
There exists $a_{k}>0$ and corresponding index set $S_{a_{k}}^{(k)}$ such that $\boldsymbol{\delta}^{(k)}=\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)}}+\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)c}}$ with $\|\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)}}\|_{0}\leq s_{k},\|\boldsymbol{\delta}^{(k)}_{S_{a_{k}}^{(k)c}}\|_{1}\leq\eta_{k}$.
Local boundness and smoothness of conditional density functions in Condition 1 are very familiar in QR literatures. Here a uniform lower bound of $f_{i}^{(k)}$ is not assumed, since we want to reveal the effect of residual shift in the error bounds. Condition 2 assumes a modified version of restricted eignevalue condition (RSC) with incorporating the $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$. The reason for placing $\mathbb{E}[f_{i}^{(k)}]$ on the right-hand side is to technically derive error bounds involving each study’s $\tau$-th density. Note that if the study model is homogeneous, then this condition will return to the original RSC form. Condition 3 needs covariates following sub-Gaussian distributions, which relaxes the former requirement of bounded design (Tian and Feng,, 2023; Huang et al.,, 2023). As for Condition 4, we impose an approximate sparse structure on $\boldsymbol{\delta}^{(k)}$ with the sparse support $S_{a_{k}}^{(k)}$ to be different across source studies, which is also similarly assumed in Zhang and Zhu, (2023) and Huang et al., (2023).
Theorem 1.
(Convergence rate for $\widehat{\boldsymbol{\beta}}$ with known transferable set $\mathcal{C}$)
Assume that Conditions 1-4 is satisfied. Define $n_{\mathcal{C}}=n_{0}+\sum_{k\in\mathcal{C}}n_{k}$. Let $\widehat{\boldsymbol{\beta}}$ is obtained from (4) with transferable set $\mathcal{C}$. For each study $k\in\{0\}\cup\mathcal{C}$, let
$\lambda_{\boldsymbol{\beta}}=c_{\lambda}\sqrt{n_{\mathcal{C}}\log p},\qquad\lambda_{k}=c_{\lambda}n_{k}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}},$
where $c_{\lambda}$ is a large enough constant. Let $r_{k}=s_{k}\sqrt{\log p/(n_{k}\wedge n_{0})}+\eta_{k}$, if
$\frac{s\log p}{n_{0}^{1/4}\sqrt{n_{k}}}+\frac{s(\log p)^{3/2}}{n_{k}}+r_{k}\sqrt{\frac{\log p}{\sqrt{n_{k}\wedge n_{0}}}}=o\left(\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\right)$ (6)
holds for each source study $k\in\mathcal{C}$, then for any constant $\varepsilon>0$, with probability at least $1-\varepsilon-(|\mathcal{C}|+1)A_{1}\exp(-A_{2}\log p)$, we have that
$\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{2}$ $\leq C_{1}\frac{\sqrt{n_{\mathcal{C}}s\log p}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}+C_{2}\sqrt{\frac{\sum_{k\in\mathcal{C}}n_{k}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}},$ (7)
$\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{1}$ $\leq C_{1}\frac{s\sqrt{n_{\mathcal{C}}\log p}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}+C_{2}\sqrt{\frac{\sum_{k\in\mathcal{C}}sn_{k}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}},$ (8)
where $A_{1},A_{2},C_{1},C_{2}$ are positive constants and their detailed formation are shown in the supplementary materials.
Remark 4.
The assumption (6) is a technical requirement to bound the infinite norms of subgradients in our proofs. Note that $r_{k}$ is actually the $\ell_{1}$-error bound of each $\widehat{\boldsymbol{\delta}}^{(k)}$ under soft sparsity. A sufficient condition for (6) is to let $s^{6}(\log p)^{2}=o(n_{k})$ and $r_{k}=o(n_{k}^{-1/4})$ for each $k\in\mathcal{C}$, which is not harsh in the high-dimensional setting. Compared to Theorem 1 in Li et al., (2024) under the similar transfer framework, we relaxed their target size assumption $|\mathcal{C}|n_{0}\ll n_{\mathcal{C}}$, and our choice for $\lambda_{\boldsymbol{\beta}}$ frees the dependence on the unknown parameters $s$ and $h_{1}$. Moreover, our analysis allows the size of transferable set $|\mathcal{C}|$ goes to infinity.
The $\ell_{2}/\ell_{1}$-error bounds of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ in (7)/(8) contain two parts. The first part can be seen as the inherent estimation error $\sqrt{s\log p/n_{\mathcal{C}}}$ accompanied by the variance term $\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[f_{i}^{(k)}]$ from the residual shift. The second part represents the bias term involving each contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ from source’s parameter shift weighted by the sample size $n_{k}$, and also includes the variance term. The error bounds reveal that to improve the transfer performance, source studies need to possess small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ as well as large $\mathbb{E}[f_{i}^{(k)}]$, strongly confirming the rationality of our screening criteria (3). Due to the weighted summation form, the optimal informative level of $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and $\mathbb{E}[f_{i}^{(k)}]$ is nearly impossible to derive, since the impact of each source depends on its sample size proportion $n_{k}/n_{\mathcal{C}}$ and other sources’ shift level, let alone the implicit dynamic choice of set $\mathcal{C}$. That’s why we offer the compromised universal levels $h_{1},h_{2}$ in our transferable set $\mathcal{C}_{h}$. For better understanding of our Theorem 1, here we give several quantitative explanations under some simplified settings, taking the $\ell_{2}$ error bound (7) as the example.
  • (Parameter Shift) If all $\tau$-th densities $\mathbb{E}[f_{i}^{(k)}]$ are neglected, then the result (7) can be simplified to
    $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{\mathcal{C}}}}+\sqrt{\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}}.$
    Parallel to the previous transfer theory, the first term $\sqrt{s\log p/n_{\mathcal{C}}}$ enjoys a sharper rate than the single-task $\ell_{2}$-error $\sqrt{s\log p/n_{0}}$ (Belloni and Chernozhukov,, 2011), which indeed reflects the superiority of transfer learning. The second term is a weighted bias term by parameter shift, and if $n_{0}=O(n_{k}),k\in\mathcal{C}$, this term can be simplified to $(\log p/n_{0})^{1/4}\sqrt{\max_{k\in\mathcal{C}}\|\boldsymbol{\delta}^{(k)}\|_{1}}$, corresponding to the result in Theorem 1 of Li et al., (2024). When the contrast magnitude $\max_{k\in\mathcal{C}}\|\boldsymbol{\delta}^{(k)}\|_{1}=o(s\sqrt{\log p/n_{0}})$, i.e., smaller than the order of single-task $\ell_{1}$-error, the transfer estimator will improve the target task performance.
  • (Residual Shift) To our knowledge, this non-asymptotic result containing $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ is new to the transfer literature, since we keep the density items along our proving process. For the first term, we can give a simple example to show that why residual shift could impact the transfer performance. Comparing the single-task $\ell_{2}$-error to the error (7) with only one source study, if the source study has huge residual shift such that $\mathbb{E}[f_{i}^{(1)}]<\frac{\sqrt{n_{0}}(\sqrt{n_{0}+n_{1}}-\sqrt{n_{0}})}{n_{1}}\mathbb{E}[f_{i}^{(0)}]$, then incorporating this source study will lead to negative transfer. In the second term, for simplicity assuming $n_{0}=O(n_{k}),k\in\mathcal{C}$, the error changes to
    $\sqrt{\sqrt{\frac{\log p}{n_{0}}}\cdot\frac{\sum_{k\in\mathcal{C}}n_{k}\|\boldsymbol{\delta}^{(k)}\|_{1}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}}.$
    The first part $\sqrt{\log p/n_{0}}$ is inherited from the target study, since we need the target information to estimate each $\boldsymbol{\delta}^{(k)}$. The second part clearly reveals the effect of both parameter/residual shift weighted by sample sizes, which again corroborates the validity of the construction of our transferable set.
  • (Covariate Shift) Our framework can relax the elusive conditions about the covariance matrices across studies (Assumption 4 in Tian and Feng, (2023), Condition 4 in Zhang and Zhu, (2023), Assumption 4 in Huang et al., (2023)). In contrast, the two-step transfer framework may amplify the error bounds by a factor $C_{\boldsymbol{\Sigma}}$, which may can diverge when the covariance matrices $\boldsymbol{\Sigma}^{(k)},k\in\mathcal{C}$ are dissimilar to $\boldsymbol{\Sigma}^{(0)}$. To put it succinctly, our results neither assume the similarity of $\boldsymbol{\Sigma}^{(k)}$, nor involve the factor $C_{\boldsymbol{\Sigma}}$ mentioned above, hence achieving better performance under covariate shift.
Corollary 1.
Under Conditions 1-4 and assumptions in Theorem 1, take the oracle transferable set $\mathcal{C}=\mathcal{C}_{h}$ and suppose $n_{0}=O(n_{k}),k\in\mathcal{C}_{h}$. Then for any constant $\varepsilon>0$, with probability at least $1-\varepsilon-(|\mathcal{C}_{h}|+1)A_{1}\exp(-A_{2}\log p)$, we can obtain that
$\|\widehat{\boldsymbol{\beta}}_{ora}-\boldsymbol{\beta^{*}}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{\mathcal{C}_{h}}}}h_{2}+(\frac{\log p}{n_{0}})^{\frac{1}{4}}\sqrt{h_{1}h_{2}},$
As shown in Corollary 1, the terms $h_{1}$ and $h_{2}$ regulate the degree of parameter and residual shift, respectively, striking a balance between performance improvement and the risk of negative transfer. For the first term, a sufficient condition to avoid negative transfer is to ensure that $h_{2}=o(\sqrt{n_{\mathcal{C}_{h}}/n_{0}})$, which indicates that the minimum $\tau$-th density in the oracle transferable set should exceed $\sqrt{n_{0}/n_{\mathcal{C}_{h}}}\mathbb{E}[f_{i}^{(0)}]$. The second term highlights the nuanced relationship between the parameter shift $h_{1}$ (measured by $\max_{k\in\mathcal{C}_{h}}\|\boldsymbol{\delta}^{(k)}\|_{1}$) and the residual shift $h_{2}$ (measured by $\max_{k\in\mathcal{C}_{h}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]$), with a sufficient condition being $h_{1}h_{2}=o(s\sqrt{\log p/n_{0}})$. This generalizes the previous result $(\log p/n_{0})^{1/4}\sqrt{h_{1}}$ in Li et al., (2024), allowing for better error control by managing both parameter and residual shift.
3 Transferable Set Detection
Since the oracle transferable set $\mathcal{C}_{h}$ may be unknown in practice, here we propose a detection procedure to screen out the source studies satisfying the requirements in (3). We can first get an initial estimate $\widetilde{\boldsymbol{\beta}}$ through $\ell_{1}$-QR on the target study, then remove the part accounted by $\widetilde{\boldsymbol{\beta}}$ on each source study and conduct $\ell_{1}$-QR to obtain the contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$. With contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$, screening studies with small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ is relatively simple. We can directly calculate each $\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}$ and set a threshold to reject the dissimilar studies. As implied by Corollary 1, $h_{1}$ should be no more than the single-task error, here we set the threshold as $t_{1}\sqrt{\log p/n_{0}}$, where $t_{1}$ is a thresholding parameter. The first screening set is then defined as $\widehat{\mathcal{C}}_{1}=\{k:\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}\leq t_{1}\sqrt{\log p/n_{0}}\}$. We will show that under some regularity conditions, appropriate choice of $t_{1}$ will lead to consistent selection conforming with the requirement $\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1}$.
The second requirement of the oracle transferable set $\mathcal{C}_{h}$ is to reject source studies with tiny expectation of $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ compared to the target $\mathbb{E}[f_{i}^{(0)}]$. We employ the approxmation technique with the Powell bandwidth $b_{k}$ (Koenker,, 2005) to estimate each $\mathbb{E}[f_{i}^{(k)}]$. Specifically, for each study we calculate
$\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathbb{I}\left(|y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})|\leq b_{k}\right)/2b_{k},$
where the practical choice of bandwidth $b_{k}$ will be recommended in the simulation section. Denote the second screening set $\widehat{\mathcal{C}}_{2}=\{k:\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})\geq t_{2}\widetilde{f}^{(0)}(\widetilde{\boldsymbol{\beta}})\}$ with the thresholding parameter $t_{2}$. We will show that under some regularity conditions, appropriate choice of $t_{2}$ will lead to consistent selection conforming with the requirement $\mathbb{E}[f^{(0)}_{i}]/\mathbb{E}[f^{(k)}_{i}]\leq h_{2}$. The final detected transferable set is $\widehat{\mathcal{C}}=\widehat{\mathcal{C}}_{1}\cap\widehat{\mathcal{C}}_{2}$. Note that the calculation of estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ is exactly the first round of Algorithm 1 for all source studies. Also in our simulations, Algorithm 1 usually converges in one step, hence here we propose a unified one-shot algorithm for transferable set detection and parameter prediction, which is shown in Algorithm 2.
Input: Target data $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$, tuning parameters $\lambda_{\boldsymbol{\beta}}$ and $\{\lambda_{k}\}_{0\leq k\leq K}$, bandwidth parameters $\{b_{k}\}_{0\leq k\leq K}$.
  1. 1.
    Compute $\widetilde{\boldsymbol{\beta}}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}},\lambda_{0}]$.
  2. 2.
    Compute each $\widetilde{\boldsymbol{\delta}}^{(k)}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widetilde{\boldsymbol{\beta}})_{i=1}^{n_{k}},\lambda_{k}],1\leq k\leq K$.
  3. 3.
    Compute $\widehat{\mathcal{C}}_{1}=\{k:\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}\leq t_{1}\sqrt{\log p/n_{0}}\}$ and $\widehat{\mathcal{C}}_{2}=\{k:\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})\geq t_{2}\widetilde{f}^{(0)}(\widetilde{\boldsymbol{\beta}})\}$, where $\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathbb{I}(|y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})|\leq b_{k})/2b_{k}$.
  4. 4.
    Determine the transferable set $\widehat{\mathcal{C}}=\widehat{\mathcal{C}}_{1}\cap\widehat{\mathcal{C}}_{2}$, then solve the pooled $\ell_{1}$-QR problem
    $\widehat{\boldsymbol{\beta}}=\underset{\boldsymbol{\beta}}{\arg\min}\left\{\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}L^{(k)}(\boldsymbol{\beta}+\widetilde{\boldsymbol{\delta}}^{(k)})+\lambda_{\boldsymbol{\beta}}\left\|\boldsymbol{\beta}\right\|_{1}\right\}.$
Output: $\widehat{\boldsymbol{\beta}}$.
Algorithm 2 Parameter prediction with transferable set detection
We next investigate the theoretical properties of Algorithm 2, which determines the transferable set $\widehat{\mathcal{C}}$. The key ingredient is to establish the detection consistency of $\widehat{\mathcal{C}}$ towards the oracle but unknown transferable set $\mathcal{C}_{h}$.
Condition 5.
For each source study $k$, let $\Omega_{1}^{(k)}:=C_{3}\frac{s_{k}}{\mathbb{E}[f_{i}^{(k)}]}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}+2\eta_{k}$, where $\eta_{k}$ is defined in Condition 4, and $C_{3}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{1}}:=\{k\in\{1,\dots,K\}:\|\boldsymbol{\delta}^{(k)}\|\leq h_{1}\}$. $\mathcal{C}_{h_{1}}$ satisfies that
$\inf_{k\in\mathcal{C}_{h_{1}}^{c}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}-\Omega_{1}^{(k)}\right)>\sup_{k\in\mathcal{C}_{h_{1}}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}+\Omega_{1}^{(k)}\right).$
Condition 6.
For each study $k$, let
$\Omega_{2}^{(k)}:=C_{4}\left(\sqrt{\frac{s+s_{k}}{b_{k}n_{k}}}+\sqrt{\frac{\eta_{k}}{b_{k}n_{k}}}(\frac{\log p}{n_{k}\wedge n_{0}})^{-1/4}+\sqrt{\frac{\log p}{n_{k}}}+s\sqrt{\frac{\log p}{n_{0}}}+r_{k}\right),$
where $r_{k}$ is defined in Theorem 1 ($r_{0}=0$), and $C_{4}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{2}}:=\{k\in\{1,\dots,K\}:\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]\leq h_{2}\}$. $\mathcal{C}_{h_{2}}$ satisfies that
$\frac{\inf_{k\in\mathcal{C}_{h_{2}}}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)}}>\frac{\sup_{k\in\mathcal{C}_{h_{2}}^{c}}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)}}.$
Remark 5.
Condition 5 and 6 assume a significant gap of quantile coefficients and $\tau$-th densities to ensure the identifiability of transferable source studies satisfying the oracle requirement in (3). In fact, $\Omega_{1}^{(k)},\Omega_{2}^{(k)}$ are estimation error bounds for $\|\widetilde{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\|_{1}$ and $|\widetilde{f}^{(k)}-\mathbb{E}[f_{i}^{(k)}]|$ respectively. Similar conditions are also assumed in Tian and Feng, (2023); Zhang and Zhu, (2023); Huang et al., (2023). Moreover, since we don’t adopt the detection method by comparing loss functions, we relax the restrictions on the similarity of covariance matrices (for example Assumption 6 in Huang et al., (2023)).
Theorem 2.
Assuming Conditions 1-6 and assumptions in Theorem 1, choose the thresholding parameters $t_{1},t_{2}$ in Algorithm 2 as
$\inf_{k\in\mathcal{C}_{h_{1}}^{c}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}-\Omega_{1}^{(k)}\right)/\sqrt{\frac{\log p}{n_{0}}}$ $\geq t_{1}\geq\sup_{k\in\mathcal{C}_{h_{1}}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}+\Omega_{1}^{(k)}\right)/\sqrt{\frac{\log p}{n_{0}}},$
$\frac{\inf_{k\in\mathcal{C}_{h_{2}}}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)}}$ $\geq t_{2}\geq\frac{\sup_{k\in\mathcal{C}_{h_{2}}^{c}}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)}}.$
Then for the detected transferable set $\widehat{\mathcal{C}}$ in Algorithm 2, with any constant $\varepsilon>0$,
$\mathbb{P}(\widehat{\mathcal{C}}=\mathcal{C}_{h})\geq 1-\varepsilon-2(K+1)A_{1}\exp(-A_{2}\log p).$
Consequently, our practical $\widehat{\boldsymbol{\beta}}$ obtained from Algorithm 2 can enjoy the same $\ell_{1}/\ell_{2}$-convergence rates in Corollary 1.
4 Statistical Inference with Knowledge Transfer
4.1 Debiased Estimator based on Neyman Orthogonality
We now turn to the statistical inference for the target parameter, $\boldsymbol{\beta^{*}}$, focusing specifically on the first component, $\beta^{*}_{1}$, without loss of generality. Previous studies (Tian and Feng,, 2023; Li et al.,, 2024; Huang et al.,, 2023) have addressed this inference task building on the principles of debiased lasso (Zhang and Zhang,, 2014; van de Geer et al.,, 2014; Javanmard and Montanari,, 2014). In these approaches, the primary method involves estimating the precision matrix of the loss function and applying a one-step correction to the $\ell_{1}$-penalized estimate. Unfortunately, these methods only achieve $\sqrt{n_{0}}$-normality for the debiased estimators, offering little improvement over debiasing based solely on the target study. The key reason for this limitation is that their debiasing procedures need the covariance matrix to be identical for all samples, a condition that is violated when covariate shift occurs. Also, they do not offer any measures to adjust for the parameter contrast $\boldsymbol{\delta}^{(k)}$. As a result, the $\sqrt{n_{0}}$-normality is suboptimal, especially given the wealth of transferable information from source studies. In an ideal scenario where $\boldsymbol{\delta}^{(k)}=0$ and the residuals $\epsilon_{i}^{(k)}$ follow a normal distribution $\mathcal{N}(0,1)$, meaning all source studies are as informative as the target study, we should attain a $\sqrt{N}$-normality by pooling all studies together. The challenge, then, is how to effectively leverage the information from these transferable source studies to enhance inference efficiency, particularly under distributional shifts.
To address this challenge, we propose a novel approach motivated by the Neyman orthogonal idea (Belloni et al.,, 2019). Before introducing the whole algorithm, we use the $k$-th source study as a methodological warm-up. Denote $\boldsymbol{v}_{-1}$ as the last $p-1$ dimensions of the vector $\boldsymbol{v}\in\mathbb{R}^{p}$. We can rewrite the original model (1) as a partially linear quantile regression
$y_{i}^{(k)}=x_{i,1}^{(k)}\beta^{*}_{1}+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}+\epsilon_{i}^{(k)},\quad i=1,\ldots,n_{k},$
where $r_{i}^{(k)}=x_{i,1}^{(k)}\delta^{(k)}_{1}$ is regarded as an approximation error. A natural moment condition for identifying $\beta^{*}_{1}$ is to solve the equation (in $\alpha$)
$\mathbb{E}\sum_{i=1}^{n_{k}}\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}\right\}-\tau\right]x_{i,1}^{(k)}=0.$ (9)
The high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$ will bring trouble for identifying $\beta^{*}_{1}$, since the first-order derivative of (9) with respect to this nuisance parameter at $\alpha=\beta^{*}_{1}$ is non-zero:
$\left.\partial_{\boldsymbol{w}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\beta^{*}_{1}+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}+r_{i}^{(k)}\right\}-\tau\right]x_{i,1}^{(k)}\right|_{\boldsymbol{w}=\boldsymbol{w}_{-1}^{(k)}}=\sum_{i=1}^{n_{k}}f_{i}^{(k)}x_{i,1}^{(k)}\boldsymbol{x}_{i,-1}^{(k)}\neq 0.$
If conducting valid inference on $\beta^{*}_{1}$ directly with moment condition (9), the estimator for $\boldsymbol{w}_{-1}^{(k)}$ needs to converge faster than $1/\sqrt{n_{k}}$, which is difficult to achieve in theory. To address this problem, we adopt the Neyman orthogonality to make the moment condition immune to the first-order error from the nuisance estimator. The technique is based on the linear projection of the regressor of interest $x_{i,1}^{(k)}$ on the $\boldsymbol{x}^{(k)}_{i,-1}$ weighted by the $\tau$-th density $f_{i}^{(k)}$:
$f_{i}^{(k)}x_{i,1}^{(k)}=f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}_{0}^{(k)}+v_{0i}^{(k)},\quad i=1,\ldots,n_{k},$ (10)
where the coefficient $\boldsymbol{\theta}_{0}^{(k)}\in\arg\min_{\boldsymbol{\theta}}\mathbb{E}\sum_{i=1}^{n_{k}}[f_{i}^{(k)}x_{i,1}^{(k)}-f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}]^{2}$, satisfying the relationship $\mathbb{E}\sum_{i=1}^{n_{k}}[f_{i}^{(k)}\boldsymbol{x}^{(k)}_{i,-1}v_{0i}^{(k)}]=0$. We then construct the orthogonal score function
$\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}):=\left[\mathbb{I}\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}\}-\tau\right]\left[f_{i}^{(k)}x_{i,1}^{(k)}-f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}_{0}^{(k)}\right],$ (11)
and solve for $\mathbb{E}\sum_{i=1}^{n_{k}}[\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)}),f_{i}^{(k)}]=0$, where this moment condition can satisfy the orthogonality with respect to high-dimensional nuisance parameters $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$, i.e.
$\left.\partial_{\boldsymbol{w}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})\right]\right|_{\boldsymbol{w}=\boldsymbol{w}_{-1}^{(k)}}=\sum_{i=1}^{n_{k}}f_{i}^{(k)}\boldsymbol{x}_{i,-1}^{(k)}v_{0i}^{(k)}=0,$
$\left.\partial_{\boldsymbol{\theta}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta},f_{i}^{(k)})\right]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}^{(k)}}=\left.\partial_{\boldsymbol{\theta}}\mathbb{E}_{\boldsymbol{x}_{i}^{(k)}}[~{}0~{}]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}^{(k)}}=0.$
Consequently, the estimators for $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$ will only introduce a second-order bias, allowing the use of high-dimensional sparse estimates. It is important to note that none of the debiased approaches can be applied to the approximate error term, $r_{i}^{(k)}=x_{i,1}^{(k)}\delta^{(k)}_{1}$, because each $\delta^{(k)}_{1}$ shares the same covariate $x_{i,1}^{(k)}$ with the target parameter $\beta^{*}_{1}$. We will address the requirements and potential solutions for handling this term in the context of knowledge transfer later.
Now let’s discuss why this orthogonal approach can conduct knowledge transfer for statistical inference in the presence of distribution shift. The key point is that this approach can avoid the computation of precision matrices and free the identical covariance matrix assumption across studies, which exactly accommodates the circumstance under covariate shift. To tackle the parameter shift and residual shift, we only incorporate moment conditions of target and sources in the oracle transferable set $\mathcal{C}_{h}$, where the oracle moment condition is to solve
$\mathbb{E}\sum_{k\in\{0\}\cup\mathcal{C}_{h}}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})\right]=0.$ (12)
The parameter identified from (12) leverages information from informative source studies rather than only debiasing on the target, hence could achieve better inference performance than previous works.
We outline several points here for practical implementation. Firstly, we estimate the $\tau$-th density $f_{i}^{(k)}$ by
$\widehat{f}_{i}^{(k)}=\frac{2h_{k}}{(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau+h_{k}}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau-h_{k}}}\quad,$ (13)
where $\widehat{\boldsymbol{\beta}}^{(k)}_{\tau\pm h_{k}}$ are computed via $\ell_{1}$-QR at $(\tau\pm h_{k})$-th quantile with tuning parameters $\lambda_{\tau}^{(k)}$ on the $k$-th study respectively. The bandwidth parameter, $h_{k}$, is set to $\min\{n_{k}^{-1/6},\tau(1-\tau)/2\}$, as suggested by Koenker, (2005); Giessing and Wang, (2023). Next, each high-dimensional projection parameter $\boldsymbol{\theta}_{0}^{(k)}$ is determined by $\ell_{1}$-penalized weighted least squares
$\widehat{\boldsymbol{\theta}}^{(k)}\in\arg\min_{\boldsymbol{\theta}}\sum_{i=1}^{n_{k}}\left[\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}\right]^{2}+\lambda_{\boldsymbol{\theta}}^{(k)}\|\boldsymbol{\theta}\|_{1},$ (14)
where $\lambda_{\boldsymbol{\theta}}^{(k)}$ is a tuning parameter and its value will be shown in the simulation section. For each high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$, we calculate the post-selected QR estimator (denoted by $\widetilde{\boldsymbol{w}}_{-1}^{(k)}$), since the refitting step after feature screening can lead to better finite sample behaviors (Belloni et al.,, 2019). Specifically, we screen each source study’s covariates by setting threshold for each element of $\widehat{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)}$ obtained in Algorithm 2, then apply traditional QR on the selected features. The empirical score function corresponding to (11) is consequently defined as
$\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)}):=\left(\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\widetilde{\boldsymbol{w}}^{(k)}_{-1}\right\}-\tau\right)\left[\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\widehat{\boldsymbol{\theta}}^{(k)}\right].$ (15)
Finally, following the moment condition (12), we aggregate the empirical score functions within the set $\widehat{\mathcal{C}}$ from Algorithm 1 and compute the debiased estimate $\widehat{\alpha}$ for $\beta_{1}^{*}$ by
$\underset{\alpha\in\mathcal{A}_{\widehat{\mathcal{C}}}}{\arg\min}\left|\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)})\right|,$
where $\mathcal{A}_{\widehat{\mathcal{C}}}=\{\alpha\in\mathbb{R}:|\alpha-\widehat{\beta}_{1}|\leq 10[n_{\widehat{\mathcal{C}}}^{-1}\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}(x^{(k)}_{i,1})^{2}]^{-1/2}/\log n_{\widehat{\mathcal{C}}}\}$ is the search region for theoretical consideration, as suggested by Belloni et al., (2019). The whole inference procedure is summarized in Algorithm 3.
Tackling the trouble by approximation error $r_{i}^{(k)}$. As noted by Belloni et al., (2019), the orthogonal approach remains theoretically valid as long as $\mathbb{E}[(r_{i}^{(k)})^{2}]=(\mathbb{E}[x_{i1}^{(k)}]\delta_{1}^{(k)})^{2}$ stays below a certain threshold. In the transfer setting, introducing more external data increases the requirement for the contrast $\delta_{1}^{(k)}$, where a sufficient theoretical bound is $o(n_{\widehat{\mathcal{C}}}^{-1/2})$. However, constructing consistent selection procedures or hypothesis tests for this condition is nearly impossible, especially when dynamically selecting $\widehat{\mathcal{C}}$. Although there is no theoretical guarantee, we propose a practical quality control procedure to validate the transfer debiased estimator. We can use prior knowledge to exclude certain source studies with non-negligible $\widetilde{\delta}_{1}^{(k)}$ and perform a bootstrap test to check the normality of the debiased estimator. If the normality hypothesis is rejected, we revert to debiasing only the target study, yielding a result similar to previous approaches. The full procedure is outlined in Algorithm 4 in the supplementary materials.
Input: Detected transferable set $\widehat{\mathcal{C}}$ and estimators $\widehat{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k\in\widehat{\mathcal{C}}}$ from Algorithm 2, target study $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source studies $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\widehat{\mathcal{C}}}$, tuning parameters $\{\lambda^{(k)},\lambda_{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$, threshold parameter $\bar{\lambda}$.
  1. 1.
    For each study $k\in\{0\}\cup\widehat{\mathcal{C}}$, estimate the $\tau$-th densities $\widehat{f}_{i}^{(k)}$ by (13), conduct the projection (14), and calculate the residuals $\widehat{v}_{i}^{(k)}=\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\widehat{\boldsymbol{\theta}}^{(k)}$.
  2. 2.
    For each study $k\in\{0\}\cup\widehat{\mathcal{C}}$, compute $\widehat{\boldsymbol{w}}^{(k)}$ by traditional QR on selected data $\{((\boldsymbol{x}_{i,\widehat{S}^{(k)}}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}$, where the index set $\widehat{S}^{(k)}=\{j:|\widehat{\beta}_{j}+\widetilde{\delta}^{(k)}_{j}|>\bar{\lambda}\}$, then enlarge $\widehat{\boldsymbol{w}}^{(k)}$ (with zero) to the $p$-dimensional vector $\widetilde{\boldsymbol{w}}^{(k)}$.
  3. 3.
    Incorporate empirical score functions (15) within target and transferable source studies and solve for
    $\widehat{\alpha}\in\underset{\alpha\in\mathcal{A}_{\widehat{\mathcal{C}}}}{\arg\min}\left|\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)})\right|.$
Output: Debiased estimator $\widehat{\alpha}$.
Algorithm 3 Inference for $\beta^{*}_{1}$ with knowledge transfer
4.2 Asymptotic Normality for Debiased Estimator
Before examining the asymptotic normality of the estimator $\widehat{\alpha}$ in Algorithm 3, we introduce regularity conditions that ensure the validity of the inference results. As in Theorem 1, we assume an arbitrarily specified transferable set $\mathcal{C}$. Let $c$ and $C$ denote given positive constants. Although these could vary across studies, incorporating them into the growth conditions in Theorem 3 below would complicate the analysis significantly. Therefore, for simplicity, we treat them as universal fixed values.
Condition 7.
For each study $k\in\{0\}\cup\mathcal{C}$, (i) assume each $\|\boldsymbol{\theta}_{0}^{(k)}\|_{2}\leq C$ (ii) There exists a universal sparsity level $s^{*}$ and vectors $\{\boldsymbol{\theta}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$ such that $s^{*}=\max_{k\in\{0\}\cup\mathcal{C}}\{s+s_{k}+\|\boldsymbol{\theta}^{(k)}\|_{0}\}$, where each $\boldsymbol{\theta}^{(k)}$ satisfies $(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}_{0}^{(k)}=(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}^{(k)}+r_{i\boldsymbol{\theta}}^{(k)}$ with $\|\boldsymbol{\theta}_{0}^{(k)}-\boldsymbol{\theta}^{(k)}\|_{1}=Cs^{*}\sqrt{\log p/n_{k}}$, $\mathbb{E}[(r_{i\boldsymbol{\theta}}^{(k)})^{2}]\leq Cs^{*}/n_{k}$. (iii) We have $c\leq\min_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{2}]^{1/2}\leq\max_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{3}]^{1/3}\leq C$.
Condition 8.
For each study $k\in\{0\}\cup\mathcal{C}$, (i) $\mathbb{E}[(\xi^{\top}\boldsymbol{x}^{(k)}_{i,-1})^{2}(r_{i}^{(k)})^{2}]\leq C\|\xi\|_{2}^{2}\mathbb{E}[(r_{i}^{(k)})^{2}]$, for any $\xi\in\mathbb{R}^{p-1}$. (ii) $\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[(r_{i}^{(k)})^{2}]=o(1/\sqrt{n_{\mathcal{C}}})$. (iii) $\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}|\mathbb{E}[f_{i}^{(k)}v_{0i}^{(k)}r_{i}^{(k)}]|=o(1/\sqrt{n_{\mathcal{C}}})$.
Condition 9.
For each study $k\in\{0\}\cup\mathcal{C}$, (i) conditional density $f(\cdot|\boldsymbol{x}_{i}^{(k)})$ are continuously differentiable and uniformly lower bounded by a positive constant $\underline{f}$. (ii) For $u=\tau\pm h_{k}$, assume that $u$-quantile$(y_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})=(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{w}^{(k)}_{u}+\epsilon_{iu}^{(k)}$, where $\mathbb{E}[\epsilon_{iu}^{(k)}]\leq o(n_{k}^{-1/2}),|\epsilon_{iu}^{(k)}|\leq o(h_{k})$ and $|\boldsymbol{w}^{(k)}_{u}|_{0}\leq s^{*}$.
Condition 7-9 actually generalize the assumptions in Belloni et al., (2019) to the transfer setting. Condition 7 follows the original settings, providing the necessary requirements for establishing convergence rates of weighted Lasso estimators $\{\widehat{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$. Condition 7 (ii) is to assume the projection parameter $\boldsymbol{\theta}^{(k)}_{0}$ in (10) can be approximated by the sparse vector $\boldsymbol{\theta}^{(k)}$ up to a small error. while Condition 7 (iii) imposes necessary mild moment conditions on the variables $f_{i}^{(k)}\boldsymbol{x}^{(k)}_{i,j}v_{0i}^{(k)}$. Condition 8 set up the moment requirements for approximation error $r_{i}^{(k)}$. Since the debiased sample size $n_{\mathcal{C}}$ is larger than the target size $n_{0}$, Condition 8 (ii) and (iii) have to link each source’s quantity $r_{i}^{k}$ to the whole sample size $n_{\mathcal{C}}$. Admittedly, due to the need of approximate orthogonality, each $\mathbb{E}[f_{i}^{(k)}v_{0i}^{(k)}r_{i}^{(k)}]=\mathbb{E}[f_{i}^{(k)}x_{i1}^{(k)}v_{0i}^{(k)}\delta_{1}^{(k)}]=\mathbb{E}[(f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}_{0}^{(k)}+v_{0i}^{(k)})v_{0i}^{(k)}\delta_{1}^{(k)}]=\mathbb{E}[(v_{0i}^{(k)})^{2}]\delta_{1}^{(k)}$ needs to very small, where a sufficient condition is $\delta^{(k)}_{1}=o(n_{\mathcal{C}}^{-1/2})$. This requirement is understandable since we need smaller parameter contrast for better normality results, plus we don’t have any approaches to remove this error. Condition 9 is for the estimation of the conditional density at $(\tau\pm h_{k})$-th quantiles, where the approximately sparse QR model is assumed around the quantile index $\tau$. All of these conditions follow the corresponding version of single-study case established in Belloni et al., (2019).
Theorem 3.
Under Condition 1-9, if the parameters satisfy the following growth conditions
$\frac{s\log p}{n_{\mathcal{C}}}+\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}=o(1/\sqrt{n_{\mathcal{C}}}),$
$\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\left[\frac{1}{h_{k}}\sqrt{\frac{s^{*}\log p}{n_{k}}}+h_{k}^{2}+\frac{\lambda_{\boldsymbol{\theta}}^{(k)}s^{*}}{n_{k}}\right]=o(n_{\mathcal{C}}^{-1/4}),$
$\max_{k\in\{0\}\cup\mathcal{C}}\left\{s^{*}+\frac{n_{k}s^{*}\log p}{(h_{k}\lambda^{(k)}_{\boldsymbol{\theta}})^{2}}+\left(\frac{n_{k}h_{k}^{2}}{\lambda^{(k)}_{\boldsymbol{\theta}}}\right)^{2}\right\}\frac{\log p}{n_{\mathcal{C}}}=o(1/\sqrt{n_{\mathcal{C}}}).$
Then our estimator $\widehat{\alpha}$ based on Algorithm 3 satisfies
$\qquad\qquad\qquad\sigma_{\mathcal{C}}^{-1}\sqrt{n_{\mathcal{C}}}(\widehat{\alpha}-\beta_{1}^{*})=\mathbb{U}_{\mathcal{C}}(\tau)+o_{p}(1),\qquad\mathbb{U}_{\mathcal{C}}(\tau)\leadsto\mathcal{N}(0,1),$
$\text{where}\quad\mathbb{U}_{\mathcal{C}}(\tau)=\left(\tau(1-\tau)\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[(v_{0i}^{(k)})^{2}]\right)^{-1/2}\cdot\frac{1}{\sqrt{n_{\mathcal{C}}}}\sum_{k\in\{0\}\cup\mathcal{C}}\sum_{i=1}^{n_{k}}\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}),$
$\text{and}\quad\sigma_{\mathcal{C}}^{2}=\left(\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[f_{i}^{(k)}x_{i,1}^{(k)}v_{0i}^{(k)}]\right)^{-2}\cdot\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[\tau(1-\tau)(v_{0i}^{(k)})^{2}].$
Moreover, the estimator $\widehat{\sigma}_{\mathcal{C}}^{2}=\tau(1-\tau)n_{\mathcal{C}}[\sum_{k\in\{0\}\cup\mathcal{C}}\sum_{i=1}^{n_{k}}(\widehat{v}_{i}^{(k)})^{2}]^{-1}$ is consistent.
With the asymptotic normality of $\widehat{\alpha}$ and consistent variance estimators $\widehat{\sigma}_{\mathcal{C}}^{2}$ provided in Theorem 3, we can construct valid confidence intervals and hypothesis test procedures for the objective component $\beta^{*}_{1}$. Now we give some explanations for results in Theorem 3. Firstly, our estimator $\widehat{\alpha}$ enjoys $\sqrt{n_{\mathcal{C}}}$-normality, which is much sharper than the previous results, demonstrating the effectiveness of our debiased transfer methods. Secondly, conditions of Theorem 3 allow for nonidentical covariate distributions across studies, as well as robust to the divergence of covariance matrices between target and source studies, demonstrating our methods’ superiority under covariate shift. Moreover, note that the asymptotic variance $\sigma_{\mathcal{C}}^{2}$ includes $v_{0i}^{(k)}$ that is inversely proportional to the $\tau$-th density $f_{i}^{(k)}$, clearly revealing that residual shift may increase the variance and deteriorate the inference efficiency, and that’s why we choose to only debias on the detected transferable set in Algorithm 3. Although the transferable set and orthogonal technique can alleviate the bias a lot, leveraging source data still needs parameter shift small enough, especially for the approximation error $x_{i1}^{(k)}\delta_{1}^{(k)}$. In the following simulations we will show that transfer debiased estimators can enjoy sharper inference results if the contrast level $\delta_{1}^{(k)}$ is small, while may lose power when higher parameter shift is introduced.
5 Simulation
In this section, we conduct extensive numerical experiments to evaluate the empirical performance of our proposed methods. Before we implement the algorithms on the simulated data, we first explain how to choose the parameters for regularization and detection thresholds, which are very important in practice.
5.1 A Practical Guide for Tuning Parameters
There are several hyperparameters in our transfer procedures, including detection thresholds $t_{1},t_{2}$, tuning parameters $\{\lambda_{k}\}_{k=0}^{K},\lambda_{\boldsymbol{\beta}}$, bandwidth parameters $\{b_{k}\}_{k=0}^{K}$ in Algorithm 2, $\{\lambda_{\tau}^{(k)},\lambda_{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}},\bar{\lambda}$ in Algorithm 3. To attain better performance in the simulations, here we propose several ways to determine these hyperparameters.
To select the $\ell_{1}$-QR tuning parameters, we deviate from the vanilla penalized quantile regression and instead follow the $\lambda$-construction idea in Belloni and Chernozhukov, (2011); Giessing and Wang, (2023). Specifically, the regularization term changes from $\lambda\|\boldsymbol{\beta}\|_{1}$ to $\lambda\sqrt{\tau(1-\tau)}\sum_{j=1}^{p}\widehat{\sigma}_{k,j}|\beta_{j}|$, with $\widehat{\sigma}_{k,j}^{2}:=n_{k}^{-1}\sum_{i=1}^{n_{k}}(\boldsymbol{x}_{ij}^{(k)})^{2}$ and $\lambda=c^{*}\cdot\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$, where $\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$ is the $(1-\alpha^{*})$-th quantile of $\Lambda_{\tau}^{(k)}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$ and
$\Lambda_{\tau}^{(k)}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}:=\max_{1\leq j\leq p}\left|\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\frac{\left(\tau-1\left\{U_{i}\leq\tau\right\}\right)\boldsymbol{x}_{ij}^{(k)}}{\sqrt{\tau(1-\tau)\widehat{\sigma}_{k,j}^{2}}}\right|,$
with $U_{1},\ldots,U_{n_{k}}$ are i.i.d. Uniform$(0,1)$ random variables, independent of the covariates $\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$. Such a pilot construction can save the computation cost a lot compared to the cross-validation. $\alpha^{*}$ is set as 0.05 uniformly, while we set $c^{*}=1$ for $\lambda_{0},\lambda_{\boldsymbol{\beta^{*}}}$, $c^{*}=1.5$ for $\{\lambda_{k}\}_{k=1}^{K}$, $c^{*}=1$ for $\{\lambda_{0.5}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$, $c^{*}=0.95$ for $\{\lambda_{0.7}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$ and $c^{*}=0.9$ for $\{\lambda_{0.2}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$. This procedure can be executed through R functions LassoLambdaHat() in the quantreg package.
For the bandwidth parameters $\{b_{k}\}_{k=0}^{K}$, we follow the construction details in Huang et al., (2023) with estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ in Algorithm 2. We set the threshold parameter $t_{1}=5,t_{2}=0.3$ in Algorithm 2 for simplicity, although one can still choose $t_{1},t_{2}$ by cross-validation. In our simulations the performance of transfer estimators is not sensitive to the choice of $t_{1},t_{2}$ as long as they fall into a proper interval. We also tried other values and the results are similar. For simplicity, we set the post-selected threshold $\bar{\lambda}=0.01$. Then for the tuning parameter $\lambda_{\boldsymbol{\theta}}^{(k)}$ in the projection, we use the common high-dimensional choice $c_{\boldsymbol{\theta}}\sqrt{\log p/n_{k}}$ in the experiments, where the constant $c_{\boldsymbol{\theta}}$ is set as $0.1$.
5.2 Estimation Errors
We follow a similar setup as in Zhang and Zhu, (2023); Huang et al., (2023). Consider a target study with size $n_{0}=200$, dimension $p=500$, and sparsity level $s=10$. We simulate two types of target model:
[…]
where $\boldsymbol{\beta^{*}}=(\mathbf{1}_{s},\mathbf{0}_{p-s})^{\top}$, $\boldsymbol{x}_{i}^{(0)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$ with $\boldsymbol{\Sigma}_{\boldsymbol{x}}=\left(0.7^{|i-j|}\right)_{1\leq i,j\leq p}$, and $\epsilon_{i}^{(0)}$ follows the shifted standard normal distribution with quantile levels $\tau\in\{0.2,0.5,0.7\}$.
The source study number $K=5$ and follows the homogeneous/heterogeneous model type (16) with $(y_{i}^{(k)},\boldsymbol{x}_{i}^{(k)},\boldsymbol{w}^{(k)},\epsilon_{i}^{(k)})_{i=1}^{n_{k}}$. To simulate the distribution shift, we consider the following parameter configurations for the source studies.
  • Covariate shift. For each source study $k=1,\ldots,K$, we set each covariate $\boldsymbol{x}_{i}^{(k)}\sim\mathcal{N}\left(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}}+\boldsymbol{\epsilon}_{\boldsymbol{x}}\boldsymbol{\epsilon}_{\boldsymbol{x}}^{T}\right)$ with $\boldsymbol{\epsilon}_{\boldsymbol{x}}\sim\mathcal{N}\left(\mathbf{0}_{p},0.3^{2}\boldsymbol{I}_{p}\right)$.
  • Parameter shift. For each source study $k=1,\ldots,K$, denote $\zeta_{i}^{(k)}$ as independent Rademacher random variables (taking values in $\{1,-1\}$ with equal probability). Following the notation in Section 3, we denote the first oracle set $\mathcal{C}_{h_{1}}$, in which the $\ell_{1}$-norm of parameter contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}=\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}\|_{1}$ of each source study is relatively small. We set
    $w_{j}^{(k)}=\begin{cases}\beta^{*}_{j}+h_{1}/100\cdot r_{j}^{(k)}\cdot\mathbb{I}\left\{j\in H^{(k)}\cup[s/2]\right\},&k\in\mathcal{C}_{h_{1}},\\ h_{1}/10\cdot r_{j}^{(k)}\cdot\mathbb{I}\left\{j\in H^{(k)}\cup[s/2]\right\},&k\in\mathcal{C}_{h_{1}}^{c},\end{cases}$
    where $[s/2]=\{1,\ldots,s/2\}$, $H^{(k)}$ is a random subset chosen from $\{s/2+1,\ldots,p\}$ with $|H^{(k)}|=50$. It can be verified that $\max_{k\in\mathcal{C}_{h_{1}}}\|\boldsymbol{\delta}^{(k)}\|_{1}<h_{1}<\min_{k\in\mathcal{C}_{h_{1}}^{c}}\|\boldsymbol{\delta}^{(k)}\|_{1}$. In this setting, we consider $h_{1}=5$.
  • Residual shift. For each source study $k=1,\ldots,K$, the residual $r_{i}^{(k)}$ either follows the standard normal distribution, or it follows one type of the following abnormal distributions:
    1. 1.
      (Cauchy) cauchy distribution: $\mathcal{C}(0,3)$;
    2. 2.
      (Mixed) mixed gaussian: $z\mathcal{N}(-3,0.5)+(1-z)\mathcal{N}(3,0.5)$, where $z\sim\mathcal{B}ernoulli(\tau)$;
    3. 3.
      (Noisy) noisy setting: $\mathcal{N}(0,5^{2})$.
    All distributions have already been shifted to meet $\mathbb{P}(\epsilon_{i}^{(k)}\leq 0|\boldsymbol{x}_{i}^{(k)})=\tau$. We denote the set containing all standard normal source studies by $\mathcal{C}_{h_{2}}$, where we can let $h_{2}=4$ satisfy $\max_{k\in\mathcal{C}_{h_{2}}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]<h_{2}<\min_{k\in\mathcal{C}_{h_{2}}^{c}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]$. To clearly demonstrate the negative impact of residual shift, we set the sample size to $n_{k}=100\cdot\mathbb{I}\{k\in\mathcal{C}_{h_{2}}\}+200\cdot\mathbb{I}\{k\in\mathcal{C}_{h_{2}}^{c}\}$.
Based on the definition in (3), we define $\mathcal{C}_{h}=\mathcal{C}_{h_{1}}\cap\mathcal{C}_{h_{2}}$ as the transferable oracle set. To capture both parameter and residual shift in one setting, we vary the parametric transferable set $\mathcal{C}_{h_{1}}$ from 1 to 5, and assume that each source study follows a divergent distribution with probability 1/2. For example, in the “Cauchy” setting, each $\epsilon_{i}^{(k)}$ either follows a normal distribution like $\epsilon_{i}^{(0)}$ or a Cauchy distribution with equal probability. The same applies to the “Mixed” and “Noisy” settings. We compare six specific methods, which are: 1) L1-QR: conduct $\ell_{1}$ quantile regression only on the target data; 2) Pooling: Algorithm 1 on all source studies; 3) Oracle: Algorithm 1 on the oracle transferable set $\mathcal{C}_{h}$; 4) TransQR: Algorithm 2, our transfer approach; 5) Oracle_PS: Algorithm 1 on $\mathcal{C}_{h_{1}}$ (only considering parameter shift), the baseline in the previous literature; 6) TransQR_2step: two-step transfer approach proposed in Huang et al., (2023).
We evaluate the performance of each method in terms of the $\ell_{2}$-error $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{2}^{2}$, the $\ell_{1}$-error $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{1}$ and the average prediction error $n_{0}^{-1}\sum_{i=1}^{n_{0}}[(\boldsymbol{x}_{i}^{t})^{\top}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})]^{2}$, where the test samples $\{\boldsymbol{x}_{i}^{t}\}_{i=1}^{n_{0}}$ generate in the same way of $\{\boldsymbol{x}_{i}^{(0)}\}_{i=1}^{n_{0}}$. We repeat the experiment for 100 times and average the results for each setting respectively, where the average $\ell_{2}$-errors are shown in Figure 5.1 and 5.2. More simulation results are reported in the supplementary materials.
Figure 5.1: $\ell_{2}$-estimation errors of various methods under homogeneous model at quantile levels $\tau=\{0.2,0.5,0.7\}$.
Figure 5.2: $\ell_{2}$-estimation errors of various methods under heterogeneous model at quantile levels $\tau=\{0.2,0.5,0.7\}$
From Figure 5.1 and 5.2, we highlight three points.
  1. 1.
    As we expect, the Oracle approach has the best performance across all scenarios because it transfers knowledge exactly from the ideal source studies. Moreover, more transferable source studies can produce smaller estimation errors while they do not affect the performance of the single task L1-QR. This fact corresponds our motivation that one can benefit from transfer learning, and conforms with our theoretical results.
  2. 2.
    Similar to the previous work, the Pooling approach doesn’t perform well when there are scarce transferable source studies. This phenomenon strongly points out the importance of the source detection step before knowledge transfer. It needs to emphasize that the residual shift indeed can influence the transfer performance, since the Oracle_PS method performs much worser than the Oracle method. Note that Oracle_PS is also conducted on the prespecified studies, which is actually regarded as the oracle method in all previous literature. The only difference from the Oracle is that Oracle_PS only considers parameter shift, i.e., small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ for screening transferable sets, which demonstrates that our proposed $\tau$-th density ratio is indeed an important index for knowledge transfer under residual shift.
  3. 3.
    It is amazing that our proposed method TransQR almost match the oracle method, while the two-step approach TransQR_2step is not satisfactory. The failure of the two-step transfer may be due to its sensitivity to covariate shift, as well as its detection procedure’s inability to effectively screen out the source studies with residual shift. This fact reflects that our method is more effective for transferable set detection and parameter estimation under distribution shift.
5.3 Inference Results
To give some intuitive understandings of Algorithm 3, we conduct simulations for statistical inference on $\beta^{*}_{1}$ under the heterogeneous model setting in (16) with some adaptive changes. We let $|\mathcal{C}_{h_{1}}|=5,h_{1}=10$ with varying $|\delta_{1}^{(k)}|\in\{0.01,0.1\}$ (controlling approximation error $r_{i}^{(k)}$), also $n_{0}=n_{k|k\in\mathcal{C}_{h_{2}}}=300$ and $n_{k|k\in\mathcal{C}_{h_{2}}^{c}}=500$. In the simulations above we have verified the detection consistency, hence for computational simplicity we assume $\mathcal{C}_{h_{2}}:=\{1,2,3\}$ is known, and assume the set in $\mathcal{C}_{h_{2}}^{c}$ follows the mixed gaussian distribution.
We compare five specific methods, including:
  1. 1.
    Debias_tar: debiasing on the target study by Algorithm 3 with the L1-QR estimate (baseline);
  2. 2.
    Debias_trans_pool: debiasing on all studies by Algorithm 3 with the Pooling estimate;
  3. 3.
    Debias_trans_tar: debiasing on the target study by Algorithm 3 with the TransQR estimate;
  4. 4.
    Debias_trans: debiasing on the target and transferable source studies by Algorithm 3 with the TransQR estimate;
  5. 5.
    Debias_dl: debiased Lasso approach proposed in Huang et al., (2023) with the TransQR_2step estimate.
Note that Debias_trans is our primary approach for comparison. However, if the quality control step (Algorithm 4) rejects the normality test, we recommend using a more conservative approach, namely Debias_trans_tar. As will be demonstrated later, when $|\delta_{1}^{(k)}|=0.1$, Debias_trans fails due to the non-negligible approximation error $r_{i}^{(k)}$, whereas Debias_trans_tar proves to be the more robust option. We assess the inference efficiency of various methods across four dimensions: (a) coverage probability of 95% confidence intervals, (b) bias comparison, (c) average length of 95% confidence intervals, and (d) density plots of normalized estimates against $\mathcal{N}(0,1)$. The experiment was repeated 1,000 times using three methods at $\tau\in\{0.2,0.5,0.7\}$. The results (a) are summarized in Table 1-2. The results (b)-(d) are summarized in Figure 5.3-5.4.
Table 1: Coverage probability of $95\%$ confidence intervals of various methods with $\delta^{(k)}_{1}=0.01$
Method $\tau=0.2$ $\tau=0.5$ $\tau=0.7$
Debias_tar 0.953 0.939 0.951
Debias_trans_pool 0.917 0.934 0.909
Debias_trans_tar 0.955 0.950 0.962
Debias_trans 0.951 0.954 0.945
Debias_dl 0.938 0.954 0.961
Table 2: Coverage probability of $95\%$ confidence intervals of various methods with $\delta^{(k)}_{1}=0.1$
Method $\tau=0.2$ $\tau=0.5$ $\tau=0.7$
Debias_tar 0.953 0.939 0.951
Debias_trans_pool 0.789 0.766 0.756
Debias_trans_tar 0.96 0.96 0.96
Debias_trans 0.846 0.809 0.817
Debias_dl 0.951 0.962 0.961
Figure 5.3: Inference results of various methods with $\delta^{(k)}_{1}=0.01$
Figure 5.4: Inference results of various methods with $\delta^{(k)}_{1}=0.1$
From Table 1-2 and Figure 5.3-5.4, we highlight three points.
  1. 1.
    From Table 1, our method Debias_trans enjoys the better coverage performance, which is the nearest to $0.95$ among all methods. Since Debias_tar, Debias_trans_tar and Debias_dl only solve the target score functions approach, their coverage probabilities are somewhat fluctuant around $0.95$. The Debias_trans_pool approach has the worst coverage properties, which again demonstrates that introducing source stuides with large residual shift will bring negative effect to the transfer efficiency.
  2. 2.
    From the estimation bias and interval length in Figure 5.3, we can see that Debias_trans have less bias and shorter interval lengths than Debias_trans_tar and Debias_tar, collaborating to the theory that our approach has faster asymptotic rates by incorporating informative sources. Although Debias_trans_pool also has the comparable biases and intervals (utilizing the most studies), it sacrificed coverage accuracy due to residual shift. Moreover, since Debias_dl debiasing on the target study, it only obtains $\sqrt{n_{0}}$-rate and hence has relatively larger bias and longer interval lengths.
  3. 3.
    By comparing the density plots with the standard normal (orange dashed line) in Figure 5.3, it can be seen that normalized estimates of all these methods are close to the standard normal, except that Debias_trans_pool at $\tau=0.7$, Debias_tar, Debias_trans_tar at $\tau=0.2$ lack a little bit efficiency. Therefore, under small $\delta_{1}^{(k)}$ our transfer approach can explore source information validly and conduct more efficient inference than previous works, which will offer more insights for statistical inference under knowledge transfer.
  4. 4.
    From Table 2 and Figure 5.4, when $|\delta_{1}^{(k)}|$ is large, unfortunately our Debias_trans performs much worse than the target debiasing methods, which is in our expectation since Condition 8 is no longer satisfied. The density plots in Figure 5.4 show that the transfer debiased estimate is no longer normal, and the Shapiro-Wilk test rejects the normality hypothesis. In a similar way, the bootstrap approach in Algorithm 4 can also indicate us that debiasing on source studies will generate unreliable estimates. Therefore, we recommend only debiasing on the target study, namely the Debias_trans_tar approach, which can keep the coverage proportions with more stable interval lengths than Debias_dl.
6 Application
The proposed transfer learning algorithm is applied to the Genotype-Tissue Expression (GTEx) data, available at https://gtexportal.org/. This dataset contains gene expression levels from 49 tissues, involving 838 individuals, and includes 1,207,976 observations for 38,187 genes. Following the previous work (Li et al.,, 2022; Zhang and Zhu,, 2023; Zhang et al.,, 2024), our study examines gene regulation within the central nervous system (CNS) across various tissues. The CNS-related genes are grouped under MODULE_137, which includes 545 genes, along with an additional 1,632 genes that show significant enrichment in the same experiments as those in the module. Please refer to https://www.gsea-msigdb.org/gsea/msigdb/cards/MODULE_137.html for a detailed description of this module.
Our focus is on the expression levels of the genes JAM2 and SH2D2A at the lower quantile. Specifically, we aim to predict the expression levels of these two genes at $\tau=0.2$ within a target tissue by leveraging the expression of other central nervous system (CNS) genes. Recent research has identified that, JAM2 and SH2D2A play a significant role in various cellular functions and disease progressions, particularly in cancer and immune response.
  • JAM2 (Junctional Adhesion Molecule 2) is a gene that encodes a protein involved in tight junctions between epithelial and endothelial cells, playing a crucial role in cell adhesion and maintaining barrier functions. It is associated with various cellular processes such as cell proliferation and migration. Studies have shown that low expression of JAM2 is linked to poor prognosis in several cancers, including lung adenocarcinoma (LUAD), where reduced JAM2 levels correlate with tumor progression and metastasis. Investigating its low expression helps in understanding cancer progression and potentially identifying therapeutic targets for limiting tumor growth (Dong et al.,, 2024).
  • SH2D2A encodes the T-cell-specific adapter protein (TSAd), which is essential for T-cell signaling and activation. The expression of SH2D2A is tightly regulated at both the transcriptional and translational levels, with cAMP signaling shown to induce its mRNA expression in T cells. However, low levels of SH2D2A protein may affect immune responses. Notably, reduced SH2D2A expression can enhance T-cell-mediated anti-tumor immunity, making it an interesting target in cancer immunotherapy research (Berge et al.,, 2012).
We analyze 13 brain tissues as separate target studies, estimating the model for each tissue individually. The target data is divided into five folds, where each fold is used as a prediction set, while the remaining four folds are used for training. The methods L1-QR, Pooling, TransQR_2step, and TransQR from the simulation section are applied in this analysis. Detailed data pre-processing and tuning parameter selection can be found in the supplementary materials. Figure 6.1 shows the relative prediction errors of Pooling, TransQR_2step, and our approach, TransQR, compared to L1-QR across the target tissues, with the black horizontal line representing a ratio of 1. As expected, TransQR consistently achieves the lowest prediction errors, significantly improving the performance of the target task in nearly all cases. We observe that the two-step approach, TransQR_2step, tends to select more source studies than TransQR, indicating that our transferable set detection can effectively identify valuable studies even under parameter or residual shift. Moreover, in some instances, TransQR_2step does not provide substantial improvements over Pooling and L1-QR, further demonstrating the robustness of our transfer framework in addressing covariate shift. Compared to the mean regression results reported in Li et al., (2022); Zhao et al., (2023), our results offer richer insights for understanding the whole distribution of gene expression.
Figure 6.1: Relative prediction errors for the expression level of gene JAM2 and SH2D2A under quantile level $\tau=0.2$.
7 Discussion
To tackle the possible negative transfer caused by distribution shift, we introduce a novel transferable set with screening both $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and $\mathbb{E}[f_{i}^{(k)}]$ to address the parameter and residual shift, also we adopt the $\ell_{1}$-minimization framework to avoid the trouble from covariate shift. Along with the transferable set detection and parameter estimation, we propose an orthogonal debiased approach for statistical inference with leveraging informative sources. Theoretical results as well as numerical experiments validate the effectiveness of our proposed method. There are several avenues to generalize our method, such as
  • We can construct source weights proportional to their respective $\tau$-th densities to achieve sharper results. Leveraging the density at objective quantiles to enhance the efficiency of estimators has been a common approach in the literature on weighted quantile regression (Jiang et al.,, 2012; Huang et al.,, 2015; Lu and Fan,, 2015; Xiong and Tian,, 2022), as the Hessian matrices of check loss functions are directly proportional to $\tau$-th densities. Furthermore, abundant source information at multiple quantiles can be utilized to predict the target study’s objective quantile coefficient. For instance, we can extend panel quantile regression (Zhang et al.,, 2019) or GMM quantile regression (Firpo et al.,, 2022) to the transfer setting.
  • Our transferable set adopts a common all-in-or-all-out manner, thus in (3) we use the expectation of $\tau$-th densities as the transferability measure. In fact, we can generalize this measure to select partial informative data from each source study, especially when the density is strongly dependent on the covariates. To be specific, define the transferable dataset (not transferable set) $\mathcal{C}^{\prime}_{h}$ as
    $\mathcal{C}^{\prime}_{h}=\left\{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)}):\quad 1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{1}{f^{(k)}_{i}}\leq h_{2}\frac{1}{f^{(0)}_{i}}\right\}.$
    Then we use the data in $\mathcal{C}^{\prime}_{h}$ to conduct knowledge transfer, where we can leverage more information from related source studies.
  • Due to the privacy or data storage concern, we may not be able to combine the individual-level data of all studies in a single machine. Assume we can only transmit summary statistics between studies, and the final result is calculated on the machine with target data (referred as the target machine). By carefully observing our Algorithm 2, we can see that almost every step is based on the local data and the summary statistics from other studies. Briefly speaking, we just need to transmit the estimators like $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ between each source machine and the target machine. The target machine can then determine the transferable set $\widehat{\mathcal{C}}$. The only step needed to pool the data from different studies is the last step of Algorithm 2. To accommodate this step to the distributed setting, we can use existing algorithms (for example Algorithm 1 in Chen et al., (2020)) to conduct distributed high-dimensional quantile regression. Therefore, our algorithm can be easily generalized to the communication-efficient version with multiple data storage nodes.
There are many interesting topics of transfer learning for future research. For example, distributionally robust transfer learning aims to obtain an optimized target distribution in a prespecified set with feasible distributions related to the source distributions. The present works mainly focus on the feasible distribution set constructed by the convex hull of source distributions (Xiong et al.,, 2023; Wang et al.,, 2023). We can extend the feasible set in terms of the Wasserstein distance metric (Blanchet et al.,, 2019; Gao and Kleywegt,, 2023), which may help us to select source domains more effectively, with learning a target model more robust to various distribution shifts.
Acknowledgment
The research of Zhu is partially supported by the National Natural Science Foundation of China 12071087, 12331009.