### v2

# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

Abstract
Information from related source studies can often enhance the findings of a target study. However, the distribution shift between target and source studies can severely impact the efficiency of knowledge transfer. In the high-dimensional regression setting, existing transfer approaches mainly focus on the parameter shift. In this paper, we focus on the high-dimensional quantile regression with knowledge transfer under three types of distribution shift: parameter shift, covariate shift, and residual shift. We propose a novel transferable set and a new transfer framework to address the above three discrepancies. Non-asymptotic estimation error bounds and source detection consistency are established to validate the availability and superiority of our method in the presence of distribution shift. Additionally, an orthogonal debiased approach is proposed for statistical inference with knowledge transfer, leading to sharper asymptotic results. Extensive simulation results as well as real data applications further demonstrate the effectiveness of our proposed procedure.
Keywords: Knowledge transfer; High-dimensional quantile regression; Distribution shift; Transferable set; Orthogonal debiasing.
1 Introduction
Previous experiences can offer valuable insights for learning new tasks. Transfer learning improves the performance of target learners by leveraging knowledge from different but related sources [zhuang2020comprehensive], has achieved success across various areas, including natural language processing [ruder2019transfer], protein representation [fenoy2022transfer], and drug discovery [turki2017transfer]. Despite its broad applications, statistical understanding of transfer learning remains incomplete, with the fundamental challenge being the detection and integration of valuable information from heterogeneous sources.
In this paper, we focus on the transfer learning for quantile regression (QR) [koenker2017quantile] in the context of a high-dimensional setting. One of QR’s most appealing features is its ability to describe the impact of covariates not only on the mean but also on the tails of the response distribution with robustness to heteroscedastic errors. Given potential non-negligible divergence between target and source distributions, QR can provide a more comprehensive assessment of what can be transferred from the source studies.
To set up the framework, suppose we have a target dataset $((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}$ and $K$ independent source datasets $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$. Consider the linear QR models
$y_{i}^{(k)}=(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{w}^{(k)}+\epsilon_{i}^{(k)},\qquad i=1,\dots,n_{k},\quad k=0,\dots,K,$ (1)
where $\boldsymbol{w}^{(k)}\in\mathbb{R}^{p}$ is the quantile coefficient of the response $y_{i}^{(k)}$ conditional on covariates $\boldsymbol{x}_{i}^{(k)}\in\mathbb{R}^{p}$ at a given quantile level $\tau\in(0,1)$, and $\epsilon_{i}^{(k)}$ is the error term satisfying $\mathbb{P}(\epsilon_{i}^{(k)}\leq 0|\boldsymbol{x}_{i}^{(k)})=\tau$ with conditional density function $f^{(k)}(\cdot|\boldsymbol{x}_{i}^{(k)})$. The parameter of interest is $\boldsymbol{\beta^{*}}:=\boldsymbol{w}^{(0)}$ on the target data, which is assumed to be sparse, i.e., $\|\boldsymbol{\beta^{*}}\|_{0}=s\ll\min\{n_{0},p\}$. In the high-dimensional setting, the number of covariates $p$ is very large, possibly larger than the overall sample size $N:=\sum_{k=0}^{K}n_{k}$. Our goal is to estimate the target parameter $\boldsymbol{\beta^{*}}$, with the help of abundant external data.
Three types of distribution shift. For the QR model (1) specified in all studies, there may exist three types of distribution shift between the target and source domains, which are: 1) parameter shift: $\boldsymbol{\beta^{*}}\neq\boldsymbol{w}^{(k)}$, i.e., the model coefficients differ; 2) residual shift: $P(\epsilon_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})\neq P(\epsilon_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$, i.e., the conditional distributions of model residuals differ; 3) covariate shift: $P(\boldsymbol{x}_{i}^{(0)})\neq P(\boldsymbol{x}_{i}^{(k)})$, i.e., the marginal covariate distributions differ. Both parameter and residual shift can be seen as label shift, meaning that the conditional label distributions $P(y_{i}^{(0)}|\boldsymbol{x}_{i}^{(0)})$ and $P(y_{i}^{(k)}|\boldsymbol{x}_{i}^{(k)})$ differ. Many examples have revealed that blindly transferring sources under distribution shift may pose challenges to the final performance and lead to the negative transfer [wang2019characterizing, zhang2022survey]. Hence, the core issue of this paper is:
How to identify the transferable information from different source studies under distribution (covariate/parameter/residual) shift? How to conduct valid prediction and inference for high-dimensional quantile regression with knowledge transfer?
Closely related to transfer learning, numerous approaches have been developed for multi-source data analysis, including federated learning [mcmahan2017communication], distributed learning [jordan2018communication, fan2023communication], fusion learning [shen2020fusion, li2023efficient], and multi-task learning [duan2023adaptive, knight2024multi]. However, distinguished from multi-source settings where each dataset is equally important, transfer learning aims to leverage information from source domains to improve the performance of the target task. Statistical investigations into transfer learning have spanned diverse contexts, including linear regression [li2022transfer, tian2022transfer], nonparametric classification [cai2021transfer], nonparametric regression [cai2024transfer], graphical models [li2023transfer], and causal inference [wei2023transfer].
To address parameter shift, [li2022transfer] introduced the contrast vector $\boldsymbol{\delta}^{(k)}:=\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$ and defined the transferable set $A_{h}:=\{\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h\}$, where source studies with a small contrast of parameters are considered informative. They proposed a two-step transfer procedure to first obtain an integrated estimate based on informative sources, then debias this estimate on the target data. The two-step framework is further extended to generalized linear regression [tian2022transfer] and quantile regression [zhang2022transfer, huang2023estimation]. In addition, several studies have investigated covariate shift and proposed remedies in high-dimensional linear regression [li2023estimation, zhao2023residual, he2024transfusion].
To the best of our knowledge, none have explicitly accounted for residual shift. However, many applications highlight the necessity of addressing residual shift in heterogeneous settings, such as in image recognition [park2023label] and disease diagnosis [chen2022estimating]. Here a motivating example is provided to illustrate the potential negative transfer caused by residual shift. We consider a similar setting to the simulation in [zhang2022transfer] with no parameter or covariate shift. Specifically, we assume only one source study with different residual distributions. Figure 1 compares the average $\ell_{2}$-errors of four methods for estimating the target parameter at $0.2$-th quantile. More details can be found in the supplementary materials.
Figure 1: Average $\ell_{2}$-errors at $0.2$-th quantile, where source residual distributions are: 1) Normal: $\mathcal{N}(0,1)$; 2) Cauchy: $\mathcal{C}(0,5)$; 3) Mixed: mixed $\mathcal{N}(-3,0.5)$ and $\mathcal{N}(3,0.5)$; 4) Noisy: $\mathcal{N}(0,5^{2})$.
As shown in Figure 1, apart from the normal setting that demonstrates the anticipated effect, all other settings reveal significant negative transfer, strongly suggesting that assessing transferability based solely on parameter similarity is insufficient in cases involving residual shift. Notably, residual shift fundamentally reflects the information quality of a source study around the objective quantile, where this difference can be characterized by the local density and sample size. Accordingly, we introduce a novel transferable set that simultaneously accounts for parameter and residual shift,
$\mathcal{C}_{h}=\left\{k:1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{n_{0}\mathbb{E}[f^{(0)}(0|\boldsymbol{x}_{i}^{(0)})]}{n_{k}\mathbb{E}[f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})]}\leq h_{2}\right\},$
where $h_{1},h_{2}$ represent oracle yet unknown transferability levels. This set is designed to filter out source studies with large parameter discrepancies or limited information at the $\tau$-th quantile. We further elaborate on the rationale behind this set in Section 2. With this transferable set, we develop a comprehensive transfer learning framework to address parameter/covariate/residual shift concurrently for high-dimensional quantile regression.
Our contributions are threefold. First, we introduce a new transferable set to handle both parameter and residual shift, offering deeper insights into managing distribution shift under knowledge transfer. For parameter estimation, we extend the constrained $\ell_{1}$-minimization [li2023estimation] to quantile regression, ensuring robustness in the presence of covariate shift. We further develop a corresponding source detection procedure and integrate it into a unified transfer estimation algorithm. Second, in theory, we establish non-asymptotic upper error bounds for our transfer estimators, broadening existing results to account for residual shift in addition to parameter shift. We also derive the minimax lower error bound to demonstrate the rationale of our approach, and provide the detection consistency of the source screening procedure. Third, a debiased transfer approach based on Neyman orthogonality [belloni2019valid] is proposed for statistical inference on quantile coefficients, which to our knowledge, is the first to debias on informative sources to enhance inference efficiency under distribution shift. By integrating valuable source studies, our approach achieves sharper asymptotic rates than existing results with $\sqrt{n_{0}}$-normality [tian2022transfer, huang2023estimation], leading to more efficient debiased estimates and shorter confidence intervals.
The rest of the article is organized as follows. In Section 2, we introduce our transfer framework under distribution shift. Section 3 describes the source detection procedure. In Section 4, we present the orthogonal debiased approach for statistical inference with knowledge transfer. Simulation studies and an application to Genotype-Tissue Expression (GTEx) dataset are provided in Section 5 and 6, respectively.
2 Transfer Learning under Distribution Shift
We begin this section with some necessary notations. For two sequences of positive numbers $\left\{a_{n}\right\}$ and $\left\{b_{n}\right\}$, we write $a_{n}\lesssim b_{n}$ if $a_{n}\leq cb_{n}$ for some universal constant $c\in(0,\infty)$, and $a_{n}\gtrsim b_{n}$ if $a_{n}\geq c^{\prime}b_{n}$ for some universal constant $c^{\prime}\in(0,\infty)$. We say $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. For any $q\in[0,\infty]$ and vector $\boldsymbol{x}=(x_{1},\cdots,x_{p})\in\mathbb{R}^{p}$, we write $\|\boldsymbol{x}\|_{q}$ for its $l_{q}$-norm. Let $a\wedge b$ and $a\vee b$ denote the minimum and maximum of $a$ and $b$, respectively. For an index subset $S\subseteq\{1,2,\dots,p\}$, let $|S|$ denote its cardinality, and let $S^{c}$ denote its complement, i.e., $S^{c}=\{1,2,\dots,p\}\setminus S$. For a vector $\boldsymbol{v}\in\mathbb{R}^{p}$, we use $\boldsymbol{v}_{S}$ to denote the restriction of vector $\boldsymbol{v}$ to the index set $S$. Denote $\boldsymbol{e}_{j}=(\boldsymbol{0}_{j-1}^{\top},1,\boldsymbol{0}_{p-j}^{\top})^{\top}$ as the unit vector. The sub-Gaussian norm of a random variable $u\in\mathbb{R}$ is $\|u\|_{\psi_{2}}=\inf\{t>0:\mathbb{E}[\exp(u^{2}/t^{2})]\leq 2\}$ and the sub-Gaussian norm of a random vector $\boldsymbol{U}\in\mathbb{R}^{n}$ is $\|\boldsymbol{U}\|_{\psi_{2}}=\sup_{\|\boldsymbol{v}\|_{2}=1,\boldsymbol{v}\in\mathbb{R}^{n}}\|\langle\boldsymbol{U},\boldsymbol{v}\rangle\|_{\psi_{2}}$. The expectation symbol $\mathbb{E}[(\cdot)_{i}^{(k)}]:=\mathbb{E}_{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})}[(\cdot)_{i}^{(k)}]$ means obtaining expectation with the joint distribution $P(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)})$.
2.1 Methodology of Tackling Three Types of Distribution Shift
Parameter Shift. Naturally we should consider a source study as informative if its parameter contrast $\boldsymbol{\delta}^{(k)}$ is relatively small. We choose the $\ell_{1}$-sparsity to quantify the informative level, as in many practical applications, parameter shift typically occurs across multiple dimensions without the overall magnitude growing too fast. Formally, the screening criterion is to reject source studies with $\|\boldsymbol{\delta}^{(k)}\|_{1}$ larger than an oracle transferring level, which has been widely adopted in the previous transfer work [li2022transfer, tian2022transfer].
Residual Shift. We start from the simplest low-dimensional linear regression to understand how to tackle the residual shift. Suppose the target and source model as
$\left\{\begin{array}[]{l}\text{target model: }~y_{i}^{(0)}=(\boldsymbol{x}_{i}^{(0)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(0)},~i=1,\dots,n_{0},~\epsilon_{i}^{(0)}\sim\mathcal{N}(0,\sigma_{0}^{2}),\\ \text{source model: }~y_{i}^{(1)}=(\boldsymbol{x}_{i}^{(1)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(1)},~i=1,\dots,n_{1},~\epsilon_{i}^{(1)}\sim\mathcal{N}(0,\sigma_{1}^{2}).\end{array}\right$.
Denote the covariance matrix $\Sigma_{x}:=\mathbb{E}[\boldsymbol{x}_{i}^{(0)}(\boldsymbol{x}_{i}^{(0)})^{\top}]=\mathbb{E}[\boldsymbol{x}_{i}^{(1)}(\boldsymbol{x}_{i}^{(1)})^{\top}]$. Based on the knowledge of linear models, the target and pooling OLS estimator $\widehat{\boldsymbol{\beta}}_{tar},\widehat{\boldsymbol{\beta}}_{pool}$ satisfy:
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{\sigma_{0}^{2}}{n_{0}}\Sigma_{x}^{-1}\right),\qquad\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\sim\mathcal{N}\left(0,\frac{n_{0}\sigma_{0}^{2}+n_{1}\sigma_{1}^{2}}{(n_{0}+n_{1})^{2}}\Sigma_{x}^{-1}\right).$
If $\sigma_{1}^{2}$ is much larger than $\sigma_{0}^{2}$, i.e., the source study is much more noisy, then the pooling estimator will be much less efficient compared to the target one.
Moving to the quantile regression with traditional loss $\rho_{\tau}(x):=x(\tau-\mathbb{I}(x\leq 0))$, assume $\epsilon_{i}^{(0)}\perp\boldsymbol{x}_{i}^{(0)},\epsilon_{i}^{(1)}\perp\boldsymbol{x}_{i}^{(1)}$ and $\mathbb{P}(\epsilon_{i}^{(0)}\leq 0)=\mathbb{P}(\epsilon_{i}^{(1)}\leq 0)=\tau$, denote the density functions of $\epsilon_{i}^{(0)},\epsilon_{i}^{(1)}$ as $f_{0}(\cdot),f_{1}(\cdot)$, respectively. Under some regular conditions, it is easy to verify that
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\tau(1-\tau)}{n_{0}(f^{(0)}(0))^{2}}\Sigma_{x}^{-1}\right),~\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\tau(1-\tau)(n_{0}+n_{1})}{(n_{0}f^{(0)}(0)+n_{1}f^{(1)}(0))^{2}}\Sigma_{x}^{-1}\right).$
If $f^{(0)}(0)$ is much smaller than $f^{(1)}(0)$, for example, when $n_{1}f^{(1)}(0)\leq(\sqrt{1+n_{1}/n_{0}}-1)n_{0}f^{(0)}(0)$, $\widehat{\boldsymbol{\beta}}_{pool}$ will be less efficient than $\widehat{\boldsymbol{\beta}}_{tar}$. The residual density values $f^{(0)}(0)$ and $f^{(1)}(0)$ serve a similar role for QR estimators as residual variances $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ do for OLS estimators, both indicating the quality of information available for estimation. Also, information quality depends on the sample size, since given the density function, a larger sample size provides richer quantile information. Consequently, another screening criterion should exclude source studies with small sample sizes $n_{k}$ and low densities at the $\tau$-th quantile $f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$, as they are unlikely to contribute valuable information beyond introducing noise. For simplicity, we denote $f^{(k)}_{i}:=f^{(k)}(0|\boldsymbol{x}_{i}^{(k)})$ as the conditional density of residual $\epsilon_{i}^{(k)}$ at zero, hereinafter referred to as the $\tau$-th density.
Remark 1.
Rejecting sources with lower $\tau$-th density $f^{(k)}_{i}$, i.e., leveraging sources with higher $\tau$-th density, can also be explained as reducing external variance. If a source study has a negligible variance term, i.e., $\sigma_{1}^{2}=0$, then it means accurately identifying the target parameter, leading to a wonderful transferable study. From the perspective of minimizing the mean square error (MSE), the transfer estimator benefits from a smaller contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ to reduce bias, while a larger $\tau$-th density helps to control variance. We further extend this motivating case to the covariate shift setting, where similar insights hold; please see Section 8.1.1 of supplementary materials for details.
Building on the previous analysis, we propose the oracle transferable set $\mathcal{C}_{h}$ as:
$\mathcal{C}_{h}=\left\{k:\quad 1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{n_{0}\mathbb{E}[f^{(0)}_{i}]}{n_{k}\mathbb{E}[f^{(k)}_{i}]}\leq h_{2}\right\},$ (2)
where $h_{1}$ and $h_{2}$ are oracle transferability levels characterizing the magnitude of parameter and residual shift, respectively. For each source study, the term $\|\boldsymbol{\delta}^{(k)}\|_{1}$ reflects parameter similarity, while the term $n_{k}\mathbb{E}[f^{(k)}_{i}$] captures the information quality around the objective quantile. The multiplicative form of $n_{k}\mathbb{E}[f^{(k)}_{i}$] is motivated by the structure $\sum_{k\in{0}\cup\mathcal{C}_{h}}n_{k}\mathbb{E}[f_{i}^{(k)}]$ appearing in Theorem 1 in the next subsection. To guarantee the improvement of transfer learning, we only leverage the information from source studies included in this set $\mathcal{C}_{h}$. It is important to note that, unlike $h_{1}$, which is expected to converge to zero as in the previous literature [li2022transfer, tian2022transfer], $h_{2}$ depends on the sample sizes of the target and source studies, thus is not necessary to converge to zero or diverges to infinity.
Covariate Shift. The negative effect of covariate shift is mainly brought by the divergence of covariance matrices $\boldsymbol{\Sigma}^{(k)}:=\mathbb{E}[\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]$ between target and source studies. As mentioned by [li2022transfer, he2024transfusion], in the two-step transfer approach with transferable set $\mathcal{C}_{h}$, the pooling estimator in the first step makes the contrast $\boldsymbol{\delta}^{(k)}$ amplified by the factor $C_{\boldsymbol{\Sigma}}:=1+|\mathcal{C}_{h}|\max_{k\in\mathcal{C}_{h}}\sup_{1\leq j\leq p}\|\boldsymbol{e}_{j}^{\top}(\boldsymbol{\Sigma}^{(k)}-\boldsymbol{\Sigma}^{(0)})(\sum_{k\in\mathcal{C}_{h}}\boldsymbol{\Sigma}^{(k)})^{-1}\|_{1}$, which may diverge when source covariance matrices $\boldsymbol{\Sigma}^{(k)}$ are dissimilar to the target $\boldsymbol{\Sigma}^{(0)}$. To address this problem, here we generalize the constrained optimization algorithm proposed by [li2023estimation] to the field of quantile regression, for jointly estimating the target parameter $\boldsymbol{\beta^{*}}$ and contrast vectors $\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}$. Our transfer framework is formalized as:
$\left(\widehat{\boldsymbol{\beta}},\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}\right)=\underset{\boldsymbol{\beta},\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}}{\arg\min}\lambda_{\boldsymbol{\beta}}\left\|\boldsymbol{\beta}\right\|_{1}+\sum_{k\in\mathcal{C}_{h}}\lambda_{k}\|\boldsymbol{\delta}^{(k)}\|_{1},$ (3)
$\text{subject to }\left\{\begin{array}[]{l}\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k},\qquad\text{for }k\in\{0\}\cup\mathcal{C}_{h},\\ \|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}},\end{array}\right$.
where $S_{n_{k}}(\boldsymbol{\beta}):=\partial L^{(k)}(\boldsymbol{\beta})/\partial\boldsymbol{\beta}$ is the subgradient of the quantile loss function $L^{(k)}(\boldsymbol{\beta}):=\sum_{i=1}^{n_{k}}\rho_{\tau}(y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\boldsymbol{\beta})$, and $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$ are tuning parameters. To solve this constrained $\ell_{1}$-minimization problem (3), we can follow the iterative algorithm adopted in [li2023estimation], where the detailed procedure is summarized in Section 8.1.2 of supplementary materials. Notably, the $\ell_{1}$-sparsity regularizer for each contrast $\boldsymbol{\delta}^{(k)}$ allows us to be free from the similarity restriction of covariance matrices, hence removing the $C_{\boldsymbol{\Sigma}}$ factor and achieving robustness under covariance heterogeneity [li2023estimation, he2024transfusion].
Remark 2.
In our transfer framework (3), the objective function encourages sparse solutions of $\widehat{\boldsymbol{\beta}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}$. The constraint $\|S_{n_{0}}(\boldsymbol{\beta})\|_{\infty}\leq\lambda_{0}$ is inherited from the target model, imposing that $\boldsymbol{\beta}$ should be identified as the true parameter $\boldsymbol{\beta^{*}}$ in the target task. The constraint $\|S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{k}$ comes from the $k$-th source model, imposing that $\boldsymbol{\delta}^{(k)}$ should be identified as the true contrast $\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}$. The last constraint $\|\sum_{k\in\{0\}\cup\mathcal{C}_{h}}S_{n_{k}}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})\|_{\infty}\leq\lambda_{\boldsymbol{\beta}}$ aggregates target study and studies in the oracle transferable set $\mathcal{C}_{h}$, revealing that our estimation leverages information from valuable sources. In contrast to the two-step transfer framework [li2022transfer, tian2022transfer], our algorithm can simultaneously estimate $\boldsymbol{\beta^{*}}$ and $\boldsymbol{\delta}^{(k)}$, providing more information for the downstream inference task.
2.2 Theory on Estimation Error
Now we study the non-asymptotic convergence rate of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ obtained by the transfer approach (3) with transferable set $\mathcal{C}$, where $\mathcal{C}$ is an arbitrary prespecified set for theoretical generality. Under the high-dimensional sparse setting, we characterize the geometric structure of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ and $\{\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\}_{k=1}^{K}$ through restrictive sets [belloni2011l1, wang2024analysis], denoted by
for $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ with $\ell_{0}$-sparsity: $\Gamma_{H}=\left\{v\in\mathbb{R}^{p}:\left\|v_{S^{c}}\right\|_{1}\leq\left\|v_{S}\right\|_{1}\right\},$
for $\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}$ with $\ell_{1}$-sparsity: $\Gamma_{W}^{(k)}=\{v\in\mathbb{R}^{p}:\|v_{S_{k}^{c}}\|_{1}\leq\|v_{S_{k}}\|_{1}+2\|\boldsymbol{\delta}^{(k)}_{S_{k}^{c}}\|_{1}\},$
where $S=\{j:\boldsymbol{\beta}^{*}_{j}\neq 0,1\leq j\leq p\}$, $S_{k}=\{j:|\boldsymbol{\delta}^{(k)}_{j}|>a_{k},1\leq j\leq p\}$ with a threshold $a_{k}>0$. Furthermore, for each contrast vector $\boldsymbol{\delta}^{(k)}$, we denote $s_{k}=\|\boldsymbol{\delta}^{(k)}_{S_{k}}\|_{0},\eta_{k}=\|\boldsymbol{\delta}^{(k)}_{S_{k}^{c}}\|_{1}$. As noted in [wang2024analysis], $a_{k}$ controls the approximate sparsity level of the contrast vector $\boldsymbol{\delta}^{(k)}$. Although the exact value of each $a_{k}$ is not specified, the constraints on $s_{k}$ and $\eta_{k}$ in our subsequent theorem impose implicit requirements on $a_{k}$. For simplicity, we omit the dependence on $a_{k}$ in the notations $\Gamma_{W}^{(k)},s_{k},\eta_{k}$. We introduce three conditions below, where the index $k$ refers to each study $k\in\{0\}\cup\mathcal{C}$.
Condition 1.
There exist universal positive constants $m_{1},m_{2},m_{u}$ such that
$\min\left\{\frac{v^{T}\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v}{\|v\|_{2}^{2}},~\frac{\|\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v\|_{\infty}}{\|v\|_{\infty}}\right\}$ $\geq m_{1}\mathbb{E}[f_{i}^{(k)}],$
$\max\left\{\frac{v^{T}\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v}{\|v\|_{2}^{2}},~\frac{\|\mathbb{E}[f_{i}^{(k)}\boldsymbol{x}_{i}^{(k)}(\boldsymbol{x}_{i}^{(k)})^{\top}]v\|_{\infty}}{\|v\|_{\infty}}\right\}$ $\leq m_{2}\mathbb{E}[f_{i}^{(k)}],$
for any non-zero vector $v\in\Gamma_{W}^{(k)}$ ($v\in\Gamma_{H}$ when $k=0$). Also, the spectral norm of covariance matrix satisfies $\lambda_{\max}(\Sigma^{(k)})\leq m_{u}$.
Condition 2.
Each covariate $\boldsymbol{x}_{i}^{(k)}$ is a mean-zero $m_{g}$-sub-Gaussian random vector, where the sub-Gaussian norm $m_{g}$ is a universal positive constant.
Condition 3.
The conditional density function $f^{(k)}(\cdot|\boldsymbol{x}_{i}^{(k)})$ is positive, continuously differentiable, and uniformly upper bounded. Also, there exist a universal positive constant $b_{0}$ such that the derivative $f_{i}^{(k)^{\prime}}(t|\boldsymbol{x}_{i}^{(k)})$ is uniformly bounded for $|t|\leq b_{0}$.
Condition 1 assumes a modified version of restricted eigenvalue condition (RSC) with incorporating the $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$. The reason for placing $\mathbb{E}[f_{i}^{(k)}]$ on the right-hand side is to technically derive error bounds involving each study’s $\tau$-th density. Note that if the study model is homoscedastic, then this condition will return to the original RSC form. Condition 2 needs covariates following sub-Gaussian distributions, which relaxes the former requirement of bounded design [tian2022transfer, huang2023estimation]. Condition 3 requires the boundness and smoothness of conditional density functions. Here a uniform lower bound of $f_{i}^{(k)}$ is not assumed, since we want to reveal the effect of residual shift in the error bounds. All three conditions are very common in the high-dimensional QR literature [belloni2011l1, wang2024analysis].
Theorem 1.
(Convergence rate for $\widehat{\boldsymbol{\beta}}$ with known transferable set $\mathcal{C}$)
Assume Conditions 1-3 are satisfied. Define $n_{\mathcal{C}}=n_{0}+\sum_{k\in\mathcal{C}}n_{k}$. Let $\widehat{\boldsymbol{\beta}}$ be obtained from the transfer framework (3) with the transferable set $\mathcal{C}$. For each study $k\in\{0\}\cup\mathcal{C}$, let $\lambda_{\boldsymbol{\beta}}=c_{\lambda}\sqrt{n_{\mathcal{C}}\log p},\lambda_{k}=c_{\lambda}n_{k}\sqrt{\log p/(n_{k}\wedge n_{0})}$, where $c_{\lambda}$ is a large enough constant. Let $r_{k}=s_{k}\sqrt{\log p/(n_{k}\wedge n_{0})}+\eta_{k}$, if
$\frac{s\log p}{n_{0}^{1/4}\sqrt{n_{k}}}+\frac{s(\log p)^{3/2}}{n_{k}}+r_{k}\sqrt{\frac{\log p}{\sqrt{n_{k}\wedge n_{0}}}}=o\left(\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\right)$ (4)
holds for each source study $k\in\mathcal{C}$, then for any constant $\varepsilon>0$, with probability at least $1-\varepsilon-(|\mathcal{C}|+1)A_{1}\exp(-A_{2}\log p)$, we have that
$\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{2}$ $\leq C_{1}\frac{\sqrt{n_{\mathcal{C}}s\log p}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}+C_{2}\sqrt{\frac{\sum_{k\in\mathcal{C}}n_{k}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}},$ (5)
$\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{1}$ $\leq C_{1}\frac{s\sqrt{n_{\mathcal{C}}\log p}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}+C_{2}\sqrt{\frac{\sum_{k\in\mathcal{C}}sn_{k}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}}{\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]}},$ (6)
where $A_{1},A_{2},C_{1},C_{2}$ are positive constants and their detailed formation are shown in the supplementary materials.
Remark 3.
The assumption (4) is a technical requirement for parameter growth, where $r_{k}$ is actually the $\ell_{1}$-error bound of each $\widehat{\boldsymbol{\delta}}^{(k)}$ under soft sparsity. A sufficient condition for (4) is to let $s^{6}(\log p)^{2}=o(n_{k})$ and $r_{k}=o(n_{k}^{-1/4})$ for each $k\in\mathcal{C}$, which is not harsh in the high-dimensional setting. Compared to Theorem 1 in [li2023estimation] under the similar transfer framework, we relaxed their target size assumption $|\mathcal{C}|n_{0}\ll n_{\mathcal{C}}$, and our choice for $\lambda_{\boldsymbol{\beta}}$ frees the dependence on the unknown parameters $s$ and $h_{1}$. Moreover, our analysis allows the size of transferable set $|\mathcal{C}|$ goes to infinity.
The error bounds of $\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}$ in (5)/(6) contain two parts. The first part can be seen as the inherent estimation error $\sqrt{s\log p/n_{\mathcal{C}}}$ weighted by the variance term $\sum_{k\in\{0\}\cup\mathcal{C}}n_{k}\mathbb{E}[f_{i}^{(k)}]$ from residual shift. The second part represents the bias term involving each contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ from parameter shift weighted by the variance term. The error bounds reveal that to improve the transfer performance, source studies need to possess small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ as well as large $n_{k}\mathbb{E}[f_{i}^{(k)}]$, strongly confirming the rationality of our screening criteria (2). Here we give some quantitative explanations under simplified settings, taking the $\ell_{2}$-error bound (5) as the example.
  • (Parameter Shift) If the $\tau$-th densities $\mathbb{E}[f_{i}^{(k)}]$ are neglected and $n_{0}=O(n_{k}),k\in\mathcal{C}$, then the error bound can be simplified to $\sqrt{s\log p/n_{\mathcal{C}}}+(\log p/n_{0})^{1/4}\sqrt{\max_{k\in\mathcal{C}}\|\boldsymbol{\delta}^{(k)}\|_{1}}$. Parallel to the previous transfer theory, the first term $\sqrt{s\log p/n_{\mathcal{C}}}$ enjoys a sharper rate than the single-task $\ell_{2}$-error $\sqrt{s\log p/n_{0}}$ [belloni2011l1, wang2024analysis], which indeed reflects the superiority of transfer learning. The second term is a bias term by parameter shift, corresponding to the result in Theorem 1 of [li2023estimation].
  • (Residual Shift) To our knowledge, this non-asymptotic result containing $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ is new to the transfer literature, since we keep the density items along our proving process. For both terms, we can observe clearly from the denominator that the contribution of each study is represented by the product $n_{k}\mathbb{E}[f_{i}^{(k)}]$, which reflects the information quality around the objective quantile and reinforces the reasonability of our transferability level $h_{2}$ defined in (2). For example, consider the case with only one source study that exhibits substantial residual shift such that $n_{1}\mathbb{E}[f_{i}^{(1)}]<(\sqrt{1+n_{1}/n_{0}}-1)n_{0}\mathbb{E}[f_{i}^{(0)}]$, then incorporating this source study will lead to negative transfer.
  • (Covariate Shift) Our framework can relax the elusive conditions about the covariance matrices across studies (Assumption 4 in [tian2022transfer], Condition 4 in [zhang2022transfer], Assumption 4 in [huang2023estimation]). Specifically, our results neither assume the similarity of $\boldsymbol{\Sigma}^{(k)}$, nor involve the factor $C_{\boldsymbol{\Sigma}}$ mentioned above, hence achieving better performance under covariate shift.
Corollary 1.
Under Conditions 1-3 and assumptions in Theorem 1, denote by $\widehat{\boldsymbol{\beta}}_{ora}$ the estimator obtained from (3) with the oracle transferable set $\mathcal{C}_{h}$. Then for any constant $\varepsilon>0$,
$\|\widehat{\boldsymbol{\beta}}_{ora}-\boldsymbol{\beta^{*}}\|_{2}\lesssim\sqrt{\frac{s\log p}{n_{\mathcal{C}_{h}}}}\frac{n_{\mathcal{C}_{h}}h_{2}}{|\mathcal{C}_{h}|n_{0}}+\sqrt{h_{1}}\left(\frac{\log p}{n_{0}}\right)^{\frac{1}{4}}\sqrt{\frac{n_{\mathcal{C}_{h}}h_{2}}{|\mathcal{C}_{h}|n_{0}}}.$
with probability at least $1-\varepsilon-(|\mathcal{C}_{h}|+1)A_{1}\exp(-A_{2}\log p)$.
As shown in Corollary 1, the terms $h_{1}$ and $h_{2}$ regulate the degree of parameter and residual shift, respectively, striking a balance between performance improvement and the risk of negative transfer. For the first term, a sufficient condition to avoid negative transfer, i.e, sharper than $\sqrt{s\log p/n_{0}}$, is to ensure $h_{2}=o(|\mathcal{C}_{h}|\sqrt{n_{0}/n_{\mathcal{C}_{h}}})$, which indicates that the information quality $n_{k}\mathbb{E}[f_{i}^{(k)}]$ in the oracle transferable set should exceed the order $\sqrt{n_{\mathcal{C}_{h}}n_{0}}\mathbb{E}[f_{i}^{(0)}]/|\mathcal{C}_{h}|$. The second term highlights the nuanced relationship between the shift levels $h_{1}$ and $h_{2}$, with a sufficient condition to avoid negative transfer being $h_{1}=o(s\sqrt{\log p/n_{0}}),h_{2}=o(|\mathcal{C}_{h}|\sqrt{n_{0}/n_{\mathcal{C}_{h}}})$. This generalizes the previous result [li2023estimation], allowing for better error control by managing both parameter and residual shift. To further clarify why incorporating $\tau$-th densities into the upper error bound is meaningful, we provide the following lower bound of $\ell_{2}$-estimation error in the minimax sense.
Condition 4.
Assume the second derivative of the logarithm of the conditional density functions $f^{(k)}(\cdot|\boldsymbol{x})$ exist and are uniformly bounded, i.e, $|[\log f_{i}^{(k)}(\cdot|\boldsymbol{x})]^{\prime\prime}|\leq L$ for any $\boldsymbol{x}$ and $k\in\{0\}\cup\mathcal{C}_{h_{1}}$ with the positive constant $L$.
Remark 4.
Condition 4 imposes a mild smoothness requirement on the conditional density function, which can be satisfied by many common distributions such as Gaussian, Logistic, and $t$ distributions.
Theorem 2.
Assume Condition 2 and 4 hold. Suppose the parameter space
$\boldsymbol{\Theta}(s,h_{1})=\{\boldsymbol{B}=(\boldsymbol{\beta},\boldsymbol{\delta}^{(1)},\dots,\boldsymbol{\delta}^{(K)}):\|\boldsymbol{\beta}\|_{0}\leq s,\max_{k\in\mathcal{C}_{h_{1}}}\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1}\}.$
If $\max\{s\log p/n_{\mathcal{C}_{h_{1}}},h_{1}(\log p/n_{0})^{1/2}\}=o(1)$, then
$\inf_{\widehat{\boldsymbol{\beta}}}\sup_{\boldsymbol{B}\in\boldsymbol{\Theta}(s,h_{1})}\mathbb{P}\left(\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta^{*}}\|_{2}\gtrsim\sqrt{\frac{s\log p}{n_{\mathcal{C}_{h_{1}}}}}+\sqrt{\frac{s\log p}{n_{0}}}\wedge\left(\frac{h_{1}^{2}\log p}{n_{0}}\right)^{1/4}\wedge h_{1}\right)\geq\frac{1}{2},$
where $\widehat{\boldsymbol{\beta}}$ is an arbitrary estimator obtained from the data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\{0\}\cup\mathcal{C}_{h_{1}}}$.
To understand the lower bound in Theorem 2, we highlight two points. First, when the residual shift is ignored and the transfer procedure (3) is applied to the transferable set $\mathcal{C}_{h_{1}}$, the $l_{2}$-error bound $\sqrt{s\log p/n_{\mathcal{C}_{h_{1}}}}+\sqrt{h_{1}}(\log p/n_{0})^{1/4}$ established in Corollary 1 is nearly minimax-optimal. The first term attains the minimax-optimal rate, whereas the second term becomes slower when $h_{1}=o(\log p/n_{0})$ or $h_{1}\gg s\sqrt{\log p/n_{0}}$, reflecting the limitations of current transfer theory for high-dimensional quantile regression. Second, since the upper bounds are essentially sharp, when the parameter shift within the transferable set is already well controlled (e.g., $h_{1}=O(\sqrt{\log p/n_{0}})$), it remains worthwhile to improve estimation efficiency by further controlling the residual shift. This improvement is particularly meaningful when the sample sizes of source studies are moderate, which underscores the necessity of introducing the additional transferability measure $h_{2}$ in (2).
3 Transferable Set Detection
Since the oracle transferable set $\mathcal{C}_{h}$ may be unknown in practice, here we propose a detection procedure to screen out the source studies satisfying the requirements in (2). We can first get an initial estimate $\widetilde{\boldsymbol{\beta}}$ through $\ell_{1}$-QR on the target study, then remove the part accounted by $\widetilde{\boldsymbol{\beta}}$ on each source study and conduct $\ell_{1}$-QR to obtain the contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$. With contrast estimators $\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$, screening studies with small $\|\boldsymbol{\delta}^{(k)}\|_{1}$ is relatively simple. We can directly calculate each $\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}$ and set a threshold to reject the dissimilar studies. As implied by Corollary 1, $h_{1}$ should be no more than the single-task error [zhang2022transfer, huang2023estimation], here we set the threshold as $t_{1}\sqrt{\log p/n_{0}}$, where $t_{1}$ is a thresholding parameter. The first screening set is then defined as $\widehat{\mathcal{C}}_{1}=\{k:\|\widetilde{\boldsymbol{\delta}}^{(k)}\|_{1}\leq t_{1}\sqrt{\log p/n_{0}}\}$. We will show that under some regularity conditions, appropriate choice of $t_{1}$ will lead to consistent selection conforming with the requirement $\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1}$.
The second requirement of the oracle transferable set $\mathcal{C}_{h}$ is to reject source studies with tiny information $n_{k}\mathbb{E}[f_{i}^{(k)}]$ compared to the target $n_{0}\mathbb{E}[f_{i}^{(0)}]$ around the $\tau$-th quantile. We employ a kernel-based approximation using the Powell bandwidth $b_{k}$ [dai2021inference, huang2023estimation, koenker2005quantile] to estimate each $\mathbb{E}[f_{i}^{(k)}]$, with the choice of $b_{k}$ detailed in the supplementary materials. Specifically, for each study we calculate
$\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})=\frac{1}{2b_{k}n_{k}}\sum_{i=1}^{n_{k}}\mathbb{I}\left(|y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})|\leq b_{k}\right).$ (7)
[…]
Note that the calculation of estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ is exactly the first round of iterative procedure for solving the $\ell_{1}$-minimization transfer framework (3). Hence here, we propose a unified one-shot algorithm for transferable set detection and parameter prediction, which is summarized in Algorithm 1. Note that for complex practical scenarios, such as strong distribution shift or heavy-tailed distributions, the procedure need to be iterated multiple times until the detected set has stabilized.
Input: Target data $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k=1}^{K}$, thresholds $t_{1}$ and $t_{2}$, tuning parameters $\lambda_{\boldsymbol{\beta}}$ and $\{\lambda_{k}\}_{0\leq k\leq K}$, bandwidths $\{b_{k}\}_{0\leq k\leq K}$.
  1. 1.
    Compute $\widetilde{\boldsymbol{\beta}}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}},\lambda_{0}]$.
  2. 2.
    Compute $\widetilde{\boldsymbol{\delta}}^{(k)}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widetilde{\boldsymbol{\beta}})_{i=1}^{n_{k}},\lambda_{k}],1\leq k\leq K$.
  3. 3.
    Compute $\{\widetilde{f}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})\}_{0\leq k\leq K}$ with bandwidths $\{b_{k}\}_{0\leq k\leq K}$ via (7).
  4. 4.
[…]
  5. 5.
    Determine the transferable set $\widehat{\mathcal{C}}=\widehat{\mathcal{C}}_{1}\cap\widehat{\mathcal{C}}_{2}$, then solve the pooled $\ell_{1}$-QR
    $\widehat{\boldsymbol{\beta}}=\underset{\boldsymbol{\beta}}{\arg\min}\left\{\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}L^{(k)}(\boldsymbol{\beta}+\widetilde{\boldsymbol{\delta}}^{(k)})+\lambda_{\boldsymbol{\beta}}\left\|\boldsymbol{\beta}\right\|_{1}\right\}.$
Output: $\widehat{\boldsymbol{\beta}}$.
Algorithm 1 Parameter prediction with transferable set detection
We next establish the detection consistency of $\widehat{\mathcal{C}}$ towards the oracle but unknown transferable set $\mathcal{C}_{h}$. To ensure the identifiability of source studies satisfying the oracle requirement in (2), we need to assume a significant gap of quantile coefficients and $\tau$-th densities, where similar conditions are also assumed in [tian2022transfer, zhang2022transfer, huang2023estimation].
Condition 5.
For each source study $k$, let $\Omega_{1}^{(k)}:=C_{3}\frac{s_{k}}{\mathbb{E}[f_{i}^{(k)}]}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}+2\eta_{k}$, where $C_{3}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{1}}:=\{k\in\{1,\dots,K\}:\|\boldsymbol{\delta}^{(k)}\|\leq h_{1}\}$. $\mathcal{C}_{h_{1}}$ satisfies that
$\inf_{k\in\mathcal{C}_{h_{1}}^{c}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}-\Omega_{1}^{(k)}\right)>\sup_{k\in\mathcal{C}_{h_{1}}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}+\Omega_{1}^{(k)}\right).$
Condition 6.
For each study $k$, let
$\Omega_{2}^{(k)}:=C_{4}\left(\sqrt{\frac{s+s_{k}}{b_{k}n_{k}}}+\sqrt{\frac{\eta_{k}}{b_{k}n_{k}}}(\frac{\log p}{n_{k}\wedge n_{0}})^{-1/4}+\sqrt{\frac{\log p}{n_{k}}}+s\sqrt{\frac{\log p}{n_{0}}}+r_{k}\right),$
where $r_{k}$ is defined in Theorem 1 ($r_{0}=0$), and $C_{4}$ is a positive constant specified in the supplementary materials. Denote the set $\mathcal{C}_{h_{2}}:=\{k\in\{1,\dots,K\}:(n_{0}\mathbb{E}[f_{i}^{(0)}])/(n_{k}\mathbb{E}[f_{i}^{(k)}])\leq h_{2}\}$. $\mathcal{C}_{h_{2}}$ satisfies that
$\inf_{k\in\mathcal{C}_{h_{2}}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)})}>\sup_{k\in\mathcal{C}_{h_{2}}^{c}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)})}.$
Remark 5.
Similar to previous transfer literature [tian2022transfer, zhang2022transfer], Conditions 5 and 6 assume the parameter contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and the population $\tau$-th density $\mathbb{E}[f_{i}^{(k)}]$ exhibit sufficient theoretical separation between transferable and non-transferable source studies, which exactly satisfies the requirement of the oracle transferable set (2). To be more specific, the value $\Omega_{1}^{(k)}$ is the upper error bound of $\|\widehat{\boldsymbol{\delta}}^{(k)}-\boldsymbol{\delta}^{(k)}\|_{1}$, while the value $\Omega_{2}^{(k)}$ is the upper error bound of $|\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\widetilde{f}_{i}^{(k)}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})-\mathbb{E}[f_{i}^{(k)}]|$.
Theorem 3.
Assuming Conditions 1-6 and assumptions in Theorem 1 for all source studies, choose the thresholding parameters $t_{1},t_{2}$ in Algorithm 1 as
$\inf_{k\in\mathcal{C}_{h_{1}}^{c}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}-\Omega_{1}^{(k)}\right)/\sqrt{\frac{\log p}{n_{0}}}$ $\geq t_{1}\geq\sup_{k\in\mathcal{C}_{h_{1}}}\left(\|\boldsymbol{\delta}^{(k)}\|_{1}+\Omega_{1}^{(k)}\right)/\sqrt{\frac{\log p}{n_{0}}},$
$\inf_{k\in\mathcal{C}_{h_{2}}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]-\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]+\Omega_{2}^{(0)})}$ $\geq t_{2}\geq\sup_{k\in\mathcal{C}_{h_{2}}^{c}}\frac{n_{k}(\mathbb{E}[f_{i}^{(k)}]+\Omega_{2}^{(k)})}{n_{0}(\mathbb{E}[f_{i}^{(0)}]-\Omega_{2}^{(0)})}.$
Then for the detected transferable set $\widehat{\mathcal{C}}$ in Algorithm 1, for any $\varepsilon>0$, there exists a positive constant $N(\varepsilon)$ such that when $\min_{k}n_{k}>N(\varepsilon)$, $\mathbb{P}(\widehat{\mathcal{C}}=\mathcal{C}_{h})\geq 1-\varepsilon$. Consequently, our practical $\widehat{\boldsymbol{\beta}}$ obtained from Algorithm 1 can enjoy the same $\ell_{1}/\ell_{2}$-convergence rates as the oracle estimator $\widehat{\boldsymbol{\beta}}_{ora}$ in Corollary 1.
4 Statistical Inference with Knowledge Transfer
4.1 Debiased Estimator based on Neyman Orthogonality
We now turn to the statistical inference for the target parameter, $\boldsymbol{\beta^{*}}$, focusing specifically on the first component, $\beta^{*}_{1}$, without loss of generality. Previous studies [tian2022transfer, li2023estimation, huang2023estimation] addressed this inference task building on the principles of debiased lasso [zhang2014confidence, van2014asymptotically]. Unfortunately, these methods only achieve $\sqrt{n_{0}}$-normality for the debiased estimators, offering little improvement over debiasing based solely on the target study. The key reason for this limitation is that their debiasing procedures need the covariance matrix to be identical for all samples, a condition that is violated when covariate shift occurs. Also, they do not offer any measures to adjust for the parameter contrast $\boldsymbol{\delta}^{(k)}$. The $\sqrt{n_{0}}$-normality is suboptimal, especially given the wealth of transferable information from source studies. The challenge, then, is how to effectively leverage the source information to enhance inference efficiency, particularly under distributional shift.
To address this challenge, we propose a novel approach motivated by the Neyman orthogonal idea [belloni2019valid]. Denote $\boldsymbol{v}_{-1}$ as the last (p1)(p-1)-dimensions of the vector $\boldsymbol{v}\in\mathbb{R}^{p}$. We can rewrite the original model (1) as a partially linear quantile regression
$y_{i}^{(k)}=x_{i,1}^{(k)}\beta^{*}_{1}+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}+\epsilon_{i}^{(k)},\quad i=1,\ldots,n_{k},$
where $r_{i}^{(k)}=x_{i,1}^{(k)}\delta^{(k)}_{1}$ is regarded as an approximation error. A natural moment condition for identifying $\beta^{*}_{1}$ is to solve the equation (in $\alpha$)
$\mathbb{E}\sum_{i=1}^{n_{k}}\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}\right\}-\tau\right]x_{i,1}^{(k)}=0.$ (8)
The high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$ will bring trouble for identifying $\beta^{*}_{1}$, since the first-order derivative of (8) with respect to this nuisance parameter at $\alpha=\beta^{*}_{1}$ is non-zero. If conducting valid inference on $\beta^{*}_{1}$ directly with moment condition (8), the estimator for $\boldsymbol{w}_{-1}^{(k)}$ needs to converge faster than $1/\sqrt{n_{k}}$, which is difficult to achieve. To address this problem, we adopt the Neyman orthogonality to make the moment condition immune to the first-order error from the nuisance estimator. The technique is based on the linear projection of the regressor of interest $x_{i,1}^{(k)}$ on the $\boldsymbol{x}^{(k)}_{i,-1}$ weighted by the $\tau$-th density $f_{i}^{(k)}$:
$f_{i}^{(k)}x_{i,1}^{(k)}=f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}_{0}^{(k)}+v_{0i}^{(k)},\quad i=1,\ldots,n_{k},$ (9)
where the coefficient $\boldsymbol{\theta}_{0}^{(k)}\in\arg\min_{\boldsymbol{\theta}}\mathbb{E}\sum_{i=1}^{n_{k}}[f_{i}^{(k)}x_{i,1}^{(k)}-f_{i}^{(k)}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}]^{2}$, satisfying the relationship $\mathbb{E}\sum_{i=1}^{n_{k}}[f_{i}^{(k)}\boldsymbol{x}^{(k)}_{i,-1}v_{0i}^{(k)}]=0$. We then construct the orthogonal score function
$\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}):=\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{w}_{-1}^{(k)}+r_{i}^{(k)}\right\}-\tau\right]v_{0i}^{(k)},$ (10)
and solve for $\mathbb{E}\sum_{i=1}^{n_{k}}[\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})]=0$, where this moment condition can satisfy the orthogonality with respect to high-dimensional nuisance parameters $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$, i.e.
$\left.\partial_{\boldsymbol{w}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)})\right]\right|_{\boldsymbol{w}=\boldsymbol{w}_{-1}^{(k)}}=\left.\partial_{\boldsymbol{\theta}}\mathbb{E}\sum_{i=1}^{n_{k}}\left[\psi_{i}^{(k)}(\beta_{1}^{*},\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta},f_{i}^{(k)})\right]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}^{(k)}}=0.$
Consequently, the estimators for $\boldsymbol{w}_{-1}^{(k)}$ and $\boldsymbol{\theta}_{0}^{(k)}$ will only introduce a second-order bias, allowing the use of high-dimensional sparse estimates. Note that this approach can avoid the computation of precision matrices and free the identical covariance matrix assumption across studies, which exactly accommodates the circumstance under distribution shift.
We outline several points here for practical implementation. Firstly, we estimate the $\tau$-th density $f_{i}^{(k)}$ by $\widehat{f}_{i}^{(k)}=2h^{(k)}/[(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau+h^{(k)}}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}^{(k)}_{\tau-h^{(k)}}]$, where $\widehat{\boldsymbol{\beta}}^{(k)}_{\tau\pm h^{(k)}}$ are computed via $\ell_{1}$-QR at the $(\tau\pm h^{(k)})$-th quantiles with tuning parameters $\lambda_{\tau}^{(k)}$ on the $k$-th study respectively. The bandwidth parameter, $h^{(k)}$, is set to $\min\{n_{k}^{-1/6},\tau(1-\tau)/2\}$, as suggested by [koenker2005quantile, giessing2023debiased]. Next, each high-dimensional projection parameter $\boldsymbol{\theta}_{0}^{(k)}$ is determined by $\ell_{1}$-penalized weighted least squares with tuning parameters $\lambda_{\boldsymbol{\theta}}^{(k)}$
$\widehat{\boldsymbol{\theta}}^{(k)}\in\arg\min_{\boldsymbol{\theta}}\sum_{i=1}^{n_{k}}\left[\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\boldsymbol{\theta}\right]^{2}+\lambda_{\boldsymbol{\theta}}^{(k)}\|\boldsymbol{\theta}\|_{1}.$ (11)
For each high-dimensional nuisance parameter $\boldsymbol{w}_{-1}^{(k)}$, we calculate the post-selected QR estimator (denoted by $\widetilde{\boldsymbol{w}}_{-1}^{(k)}$), since the refitting step after feature screening can lead to better finite sample behaviors [belloni2019valid]. Specifically, we screen each source study’s covariates by setting threshold for each element of $\widehat{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)}$ obtained in Algorithm 1, then apply traditional QR on the selected features. Denote the empirical residuals $\widehat{v}_{i}^{(k)}=\widehat{f}^{(k)}_{i}x_{i,1}^{(k)}-\widehat{f}^{(k)}_{i}(\boldsymbol{x}^{(k)}_{i,-1})^{\top}\widehat{\boldsymbol{\theta}}^{(k)}$. The empirical score function corresponding to (10) is defined as
$\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)}):=\left[\mathbb{I}\left\{y_{i}^{(k)}\leq x_{i,1}^{(k)}\alpha+(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\widetilde{\boldsymbol{w}}^{(k)}_{-1}\right\}-\tau\right]\widehat{v}_{i}^{(k)}.$ (12)
To tackle parameter and residual shift, we only incorporate the score functions of sources within the set $\widehat{\mathcal{C}}$ from Algorithm 1. Finally, we set the search region $\widehat{\alpha}\in\mathcal{A}_{\widehat{\mathcal{C}}}:=\{\alpha\in\mathbb{R}:|\alpha-\widehat{\beta}_{1}|\leq 10[n_{\widehat{\mathcal{C}}}^{-1}\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}(x^{(k)}_{i,1})^{2}]^{-1/2}/\log n_{\widehat{\mathcal{C}}}\}$, as suggested by [belloni2019valid]. The whole inference procedure is summarized in Algorithm 2.
Input: Detected transferable set $\widehat{\mathcal{C}}$ and estimators $\widehat{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k\in\widehat{\mathcal{C}}}$ from Algorithm 1,    target study $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source studies $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\widehat{\mathcal{C}}}$,    tuning parameters $\{\lambda_{\tau}^{(k)},\lambda_{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$, threshold parameter $\bar{\lambda}$.
  1. 1.
    For $k\in\{0\}\cup\widehat{\mathcal{C}}$, obtain the density estimate $\widehat{f}_{i}^{(k)}$ and conduct the projection (11).
  2. 2.
    For $k\in\{0\}\cup\widehat{\mathcal{C}}$, compute $\widehat{\boldsymbol{w}}^{(k)}$ by traditional QR on $\{((\boldsymbol{x}_{i,\widehat{S}^{(k)}}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}$, where the index set $\widehat{S}^{(k)}=\{j:|\widehat{\beta}_{j}+\widetilde{\delta}^{(k)}_{j}|>\bar{\lambda}\}$, then enlarge $\widehat{\boldsymbol{w}}^{(k)}$ (with zero) to $\widetilde{\boldsymbol{w}}^{(k)}\in\mathbb{R}^{p}$.
  3. 3.
    Incorporate empirical score functions (12) and solve for
    $\widehat{\alpha}\in\underset{\alpha\in\mathcal{A}_{\widehat{\mathcal{C}}}}{\arg\min}\left|\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}\widehat{\psi}_{i}^{(k)}(\alpha,\widetilde{\boldsymbol{w}}_{-1}^{(k)},\widehat{\boldsymbol{\theta}}^{(k)},\widehat{f}_{i}^{(k)})\right|.$
Output: Debiased estimator $\widehat{\alpha}$.
Algorithm 2 Inference for $\beta^{*}_{1}$ with knowledge transfer
4.2 Asymptotic Normality for Debiased Estimator
As in Theorem 1, we assume an arbitrarily specified transferable set $\mathcal{C}$. Let $c,C$ be generic constants which can be different at different places.
Condition 7.
For each study $k\in\{0\}\cup\mathcal{C}$, (i) $\mathbb{E}[(\xi^{\top}\boldsymbol{x}^{(k)}_{i,-1})^{2}(r_{i}^{(k)})^{2}]\leq C\|\xi\|_{2}^{2}\mathbb{E}[(r_{i}^{(k)})^{2}]$, for any $\xi\in\mathbb{R}^{p-1}$. (ii) $\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[(r_{i}^{(k)})^{2}]=o(1/\sqrt{n_{\mathcal{C}}})$. (iii) $\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}|\mathbb{E}[f_{i}^{(k)}v_{0i}^{(k)}r_{i}^{(k)}]|=o(1/\sqrt{n_{\mathcal{C}}})$.
Condition 8.
For each study $k\in\{0\}\cup\mathcal{C}$, (i) $\|\boldsymbol{\theta}_{0}^{(k)}\|_{2}\leq C$; (ii) There exist sparsity levels $\{s_{k}^{*}\}_{k\in\{0\}\cup\mathcal{C}}$ and vectors $\{\boldsymbol{\theta}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$ such that $s_{k}^{*}=\max\{s+s_{k},\|\boldsymbol{\theta}^{(k)}\|_{0}\}$, where each $\boldsymbol{\theta}^{(k)}$ satisfies $(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}_{0}^{(k)}=(\boldsymbol{x}_{i,-1}^{(k)})^{\top}\boldsymbol{\theta}^{(k)}+r_{\boldsymbol{\theta}i}^{(k)}$ with $\|\boldsymbol{\theta}_{0}^{(k)}-\boldsymbol{\theta}^{(k)}\|_{1}=Cs_{k}^{*}\sqrt{\log p/n_{k}}$, $\mathbb{E}[(r_{\boldsymbol{\theta}i}^{(k)})^{2}]\leq Cs_{k}^{*}/n_{k}$; (iii) $c\leq\min_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{2}]^{1/2}\leq\max_{1<j\leq p}\mathbb{E}[(f_{i}^{(k)}x^{(k)}_{i,j}v_{0i}^{(k)})^{3}]^{1/3}\leq C$.
Condition 9.
[…]
Conditions 7-9 generalize the assumptions of single-study version in [belloni2019valid] to the transfer setting. Condition 7 set up the moment requirements for approximation error $r_{i}^{(k)}$, where each source’s quantity $r_{i}^{(k)}$ is linked to the whole sample size $n_{\mathcal{C}}$. Condition 8 follows the original settings, providing the necessary requirements for establishing convergence rates of weighted Lasso estimators $\{\widehat{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\mathcal{C}}$. Condition 9 is for the estimation of the conditional density at $(\tau\pm h^{(k)})$-th quantiles, where the approximately sparse QR model is assumed around the quantile index $\tau$.
Remark 6.
Since our inference framework perform debiasing on the transferable source studies, this naturally requires the debiased component across studies to be very similar, which is relatively strict than the similarity requirement for estimation in Theorem 1. As in Condition 7 (iii), each term $\mathbb{E}[f_{i}^{(k)}v_{0i}^{(k)}r_{i}^{(k)}]=\mathbb{E}[(v_{0i}^{(k)})^{2}]\delta_{1}^{(k)}$ needs be sufficient small with order $o(1/\sqrt{n_{\mathcal{C}}})$. Nevertheless, this assumption is reasonable to hold in many practical scenarios. For instance, in multi-source genetic studies, researchers are often interested in estimating the common effect of a specific genetic variant on a phenotype across different sources [michailidou2017association, kunkle2019genetic]. In practice, we further propose a quality control procedure in Section 8.1.4 of supplementary materials. The procedure ensures that if the similarity of the coefficient of interest is not well supported, one should instead debias on the target study only.
Theorem 4.
Assume Conditions 7-9 and all conditions required by Theorem 1. For each study $k\in\{0\}\cup\mathcal{C}$, let $\lambda_{\tau}^{(k)}=c_{\lambda}\sqrt{n_{k}\log p}$, where $c_{\lambda}$ is a large enough constant. If the parameters satisfy the following growth conditions
$\frac{s\log p}{n_{\mathcal{C}}}+\sum_{k\in\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\sqrt{\frac{\log p}{n_{k}\wedge n_{0}}}\|\boldsymbol{\delta}^{(k)}\|_{1}=o(1/\sqrt{n_{\mathcal{C}}}),$ (13)
$\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\left[\frac{1}{h^{(k)}}\sqrt{\frac{s_{k}^{*}\log p}{n_{k}}}+(h^{(k)})^{2}+\frac{\lambda_{\boldsymbol{\theta}}^{(k)}s_{k}^{*}}{n_{k}}\right]=o(n_{\mathcal{C}}^{-1/4}),$
$\max_{k\in\{0\}\cup\mathcal{C}}\left\{s_{k}^{*}+\frac{n_{k}s_{k}^{*}\log p}{(h^{(k)}\lambda^{(k)}_{\boldsymbol{\theta}})^{2}}+\left(\frac{n_{k}(h^{(k)})^{2}}{\lambda^{(k)}_{\boldsymbol{\theta}}}\right)^{2}\right\}\frac{\log p}{n_{\mathcal{C}}}=o(1/\sqrt{n_{\mathcal{C}}}),$
then for the estimator $\widehat{\alpha}$ obtained from Algorithm 2 with the transferable set $\mathcal{C}$, we have
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\sigma_{\mathcal{C}}^{-1}\sqrt{n_{\mathcal{C}}}(\widehat{\alpha}-\beta_{1}^{*})=\mathbb{U}_{\mathcal{C}}(\tau)+o_{p}(1),\hskip 18.49988pt\mathbb{U}_{\mathcal{C}}(\tau)\leadsto\mathcal{N}(0,1),$
$\text{where}\qquad\mathbb{U}_{\mathcal{C}}(\tau)=\left(\tau(1-\tau)\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[(v_{0i}^{(k)})^{2}]\right)^{-1/2}\cdot\frac{1}{\sqrt{n_{\mathcal{C}}}}\sum_{k\in\{0\}\cup\mathcal{C}}\sum_{i=1}^{n_{k}}\psi_{i}^{(k)}(\alpha,\boldsymbol{w}_{-1}^{(k)},\boldsymbol{\theta}_{0}^{(k)},f_{i}^{(k)}),$
$\text{and}\qquad\sigma_{\mathcal{C}}^{2}=\left(\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[f_{i}^{(k)}x_{i,1}^{(k)}v_{0i}^{(k)}]\right)^{-2}\cdot\sum_{k\in\{0\}\cup\mathcal{C}}\frac{n_{k}}{n_{\mathcal{C}}}\mathbb{E}[\tau(1-\tau)(v_{0i}^{(k)})^{2}].$
Moreover, the estimator $\widehat{\sigma}_{\mathcal{C}}^{2}=\tau(1-\tau)n_{\mathcal{C}}[\sum_{k\in\{0\}\cup\mathcal{C}}\sum_{i=1}^{n_{k}}(\widehat{v}_{i}^{(k)})^{2}]^{-1}$ is consistent.
Remark 7.
As implied by the minimax lower bound in [li2023transfer], the $\sqrt{n_{\mathcal{C}}}$-rate becomes theoretically possible only when the similarity conditions are linked to the total sample size $n_{\mathcal{C}}$ instead of $n_{0}$ only. As shown in the parameter growth conditions stated in (13), we relate the parameter growth of each individual study to the total sample size $n_{\mathcal{C}}$. Owing to the Neyman orthogonal technique we employ, these requirements are not stringent, where a sufficient condition is to let $h^{(k)}\asymp n_{k}^{-1/6},\lambda_{\boldsymbol{\theta}}^{(k)}\asymp\sqrt{n_{k}\log p},s_{k}^{*}\log p=o(n_{k}^{2/3}/\sqrt{n_{\mathcal{C}}}),\|\boldsymbol{\delta}^{(k)}\|_{1}\log p=o(\sqrt{n_{k}\wedge n_{0}}/\sqrt{n_{\mathcal{C}}})$ for each study $k\in\{0\}\cup\mathcal{C}$.
With the asymptotic normality of $\widehat{\alpha}$ and consistent variance estimators $\widehat{\sigma}_{\mathcal{C}}^{2}$ provided in Theorem 4, we can construct valid confidence intervals and hypothesis test procedures for the objective component $\beta^{*}_{1}$. Now we give some explanations for results in Theorem 4. Firstly, our estimator $\widehat{\alpha}$ enjoys $\sqrt{n_{\mathcal{C}}}$-normality, which is much sharper than the previous results, demonstrating the effectiveness of our debiased transfer methods. Secondly, conditions of Theorem 4 allow for nonidentical covariate distributions across studies, as well as robust to the divergence of covariance matrices between target and source studies, demonstrating our methods’ superiority under covariate shift. Moreover, note that the asymptotic variance $\sigma_{\mathcal{C}}^{2}$ includes $v_{0i}^{(k)}$ that is inversely proportional to the $\tau$-th density $f_{i}^{(k)}$, clearly revealing that residual shift may increase the variance and deteriorate the inference efficiency, and that’s why we choose to only debias on the detected transferable set.
5 Simulation
5.1 Estimation Errors
We follow a similar setup as in [zhang2022transfer, huang2023estimation]. Consider a target study with size $n_{0}=200$, dimension $p=500$, and sparsity level $s=10$. We simulate two types of target model:
$\text{Homoscedastic: }y_{i}^{(0)}=(\boldsymbol{x}_{i}^{(0)})^{\top}\boldsymbol{\beta^{*}}+\epsilon_{i}^{(0)},~\text{Heteroscedastic: }y_{i}^{(0)}=(\boldsymbol{x}_{i}^{(0)})^{\top}\boldsymbol{\beta^{*}}+|x_{i,1}^{(0)}|\epsilon_{i}^{(0)},$ (14)
where $\boldsymbol{\beta^{*}}=(\mathbf{1}_{s},\mathbf{0}_{p-s})^{\top}$, $\boldsymbol{x}_{i}^{(0)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$ with $\boldsymbol{\Sigma}_{\boldsymbol{x}}=(0.7^{|i-j|})_{1\leq i,j\leq p}$, and $\epsilon_{i}^{(0)}$ follows the shifted standard normal distribution with quantile levels $\tau\in\{0.2,0.5,0.7\}$. The source study number $K=5$ and follows the homoscedastic/heteroscedastic model type (14) with $(y_{i}^{(k)},\boldsymbol{x}_{i}^{(k)},\boldsymbol{w}^{(k)},\epsilon_{i}^{(k)})_{i=1}^{n_{k}}$. To simulate the distribution shift, we consider the following parameter configurations for the source studies.
  • Covariate shift. For each source study, we set the covariates $\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$ follow $\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}}+\boldsymbol{\epsilon}_{\boldsymbol{x}}^{(k)}\boldsymbol{\epsilon}_{\boldsymbol{x}}^{(k)\top})$ with $\boldsymbol{\epsilon}_{\boldsymbol{x}}^{(k)}\sim\mathcal{N}(\mathbf{0}_{p},0.3^{2}\boldsymbol{I}_{p})$, then normalize their variance to one.
  • Parameter shift. For each source study, denote $\{\zeta_{j}^{(k)}\}_{j=1}^{p}$ as independent Rademacher random variables (taking values in $\{1,-1\}$ with equal probability). Following the notation in Section 3, we denote the first oracle set $\mathcal{C}_{h_{1}}$, in which the $\ell_{1}$-norm of parameter contrast $\|\boldsymbol{\delta}^{(k)}\|_{1}=\|\boldsymbol{w}^{(k)}-\boldsymbol{\beta^{*}}\|_{1}$ of each source study is relatively small. We set
    $w_{j}^{(k)}=\begin{cases}\beta^{*}_{j}+h_{1}/100\cdot\zeta_{j}^{(k)}\cdot\mathbb{I}\left\{j\in H^{(k)}\cup[s/2]\right\},&k\in\mathcal{C}_{h_{1}},\\ h_{1}/10\cdot\zeta_{j}^{(k)}\cdot\mathbb{I}\left\{j\in H^{(k)}\cup[s/2]\right\},&k\in\mathcal{C}_{h_{1}}^{c},\end{cases}$
    where $s/2]=\{1,\ldots,s/2\}$, $H^{(k)}$ is a random subset chosen from $\{s/2+1,\ldots,p\}$ with $|H^{(k)}|=50$. It can be verified that $\max_{k\in\mathcal{C}_{h_{1}}}\|\boldsymbol{\delta}^{(k)}\|_{1}<h_{1}<\min_{k\in\mathcal{C}_{h_{1}}^{c}}\|\boldsymbol{\delta}^{(k)}\|_{1}$. In this setting, we consider $h_{1}=5$.
  • Residual shift. For each source study $k=1,\ldots,K$, the residual $\epsilon_{i}^{(k)}$ either follows the standard normal distribution, or it follows one type of the following abnormal distributions: 1) (Cauchy) cauchy distribution: $\mathcal{C}(0,3)$; 2) (Mixed) mixed gaussian: $z\mathcal{N}(-3,0.5)+(1-z)\mathcal{N}(3,0.5)$, where $z\sim\mathcal{B}ernoulli(\tau)$; 3) (Noisy) noisy setting: $\mathcal{N}(0,5^{2})$. All distributions have already been shifted to meet $\mathbb{P}(\epsilon_{i}^{(k)}\leq 0|\boldsymbol{x}_{i}^{(k)})=\tau$. We denote the set containing all standard normal source studies by $\mathcal{C}_{h_{2}}$, where we can let $h_{2}=4$ satisfy $\max_{k\in\mathcal{C}_{h_{2}}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]<h_{2}<\min_{k\in\mathcal{C}_{h_{2}}^{c}}\mathbb{E}[f_{i}^{(0)}]/\mathbb{E}[f_{i}^{(k)}]$. To clearly demonstrate the negative impact of residual shift, we set the sample size to $n_{k}=100\cdot\mathbb{I}\{k\in\mathcal{C}_{h_{2}}\}+200\cdot\mathbb{I}\{k\in\mathcal{C}_{h_{2}}^{c}\}$.
Based on the definition in (2), we define $\mathcal{C}_{h}=\mathcal{C}_{h_{1}}\cap\mathcal{C}_{h_{2}}$ as the transferable oracle set. To capture both parameter and residual shift in one setting, we vary the parametric transferable set $\mathcal{C}_{h_{1}}$ from 1 to 5, and assume that each source study follows a divergent distribution with probability 1/2. For example, in the “Cauchy” setting, each $\epsilon_{i}^{(k)}$ either follows a normal distribution like $\epsilon_{i}^{(0)}$ or a Cauchy distribution with equal probability. The same applies to the “Mixed” and “Noisy” settings. We compare six specific methods, which are: 1) L1-QR: conduct $\ell_{1}$ quantile regression only on the target data (baseline); 2) Pooling: Algorithm 1 (without detection) on all source studies; 3) Oracle: Algorithm 1 (without detection) on the oracle transferable set $\mathcal{C}_{h}$; 4) TransQR: Algorithm 1, our transfer approach; 5) Oracle_PS: Algorithm 1 (without detection) on $\mathcal{C}_{h_{1}}$, the oracle choice in the previous literature; 6) TransQR_2step: two-step transfer approach proposed in [huang2023estimation]. The choices of all tuning parameters are provided in the supplementary materials.
We evaluate the performance of each method in terms of the $\ell_{2}$-error $\|\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{2}^{2}$. We repeat the experiment for 100 times and average the results for each setting respectively. The average $\ell_{2}$-error of six methods under homoscedastic model are shown in Figure 2. More extensive simulation results, including scenarios with strong covariate shift and heavy-tailed target distributions, are provided in Section 8.2 of supplementary materials.
Figure 2: Average $\ell_{2}$-error of various methods under homoscedastic model.
From Figure 2, we highlight three points. First, as expected, the Oracle approach achieves the best performance across all scenarios, whereas the Pooling approach performs the worst when transferable source studies are scarce. This finding aligns with our theoretical motivation that knowledge transfer should rely on informative sources; otherwise, negative transfer may occur. Second, the Oracle_PS method underperforms relative to the Oracle method across all settings, indicating that residual shift indeed affects transfer performance. The only difference between the two methods is that Oracle_PS accounts solely for parameter shift and includes sources with heterogeneous residual distributions. The superior performance of Oracle confirms that our proposed metric $n_{k}\mathbb{E}[f_{i}^{(k)}]$ serves as a crucial measure for controlling residual shift. Finally, it is noteworthy that our proposed method, TransQR, performs almost identically to the oracle benchmark, demonstrating the effectiveness of our transferable set detection under distribution shift. In contrast, the two-step approach TransQR_2step performs unsatisfactorily, revealing its limited ability to handle residual shift and its sensitivity to covariate shift.
5.2 Inference Results
To give some intuitive understandings of Algorithm 2, we conduct simulations for statistical inference on $\beta^{*}_{1}$ under the heteroscedastic model setting in (14) with some adaptive changes. We let $|\mathcal{C}_{h_{1}}|=5,h_{1}=10$ with $|\delta_{1}^{(k)}|=0.01$ (controlling approximation error $r_{i}^{(k)}$), also let $n_{0}=n_{k|k\in\mathcal{C}_{h_{2}}}=300$ and $n_{k|k\in\mathcal{C}_{h_{2}}^{c}}=500$. In the simulations above we have verified the detection consistency, hence for computational simplicity we assume $\mathcal{C}_{h_{2}}:=\{1,2,3\}$ is known, and assume the set in $\mathcal{C}_{h_{2}}^{c}$ follows the “Mixed” setting.
We compare five specific methods, including: 1) Debias_tar: debiasing on the target study by Algorithm 2 with the L1-QR estimate (baseline); 2) Debias_trans_pool: debiasing on all studies by Algorithm 2 with the Pooling estimate; 3) Debias_trans_tar: debiasing on the target study by Algorithm 2 with the TransQR estimate; 4) Debias_trans: debiasing on the target and transferable source studies by Algorithm 2 with the TransQR estimate; 5) Debias_dl: debiased Lasso approach proposed in [huang2023estimation] with the TransQR_2step estimate. Note that Debias_trans is our primary approach for comparison. The experiment was repeated 1,000 times at $\tau\in\{0.2,0.5,0.7\}$. The choices of all tuning parameters are provided in the supplementary materials. We evaluate the inference efficiency via multiple dimensions. Table 1 reports the coverage probabilities of 95% confidence intervals, the average absolute bias, the empirical standard error, and the average estimated standard error. Figure 3 further illustrates these results through boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared with $\mathcal{N}(0,1)$.
Table 1: Coverage probability of $95\%$ confidence intervals (CP), average absolute bias (Bias), standard error (SE), and average estimated standard error (ESE) of various methods.
Method $\tau=0.2$ $\tau=0.5$ $\tau=0.7$
CP Bias SE ESE CP Bias SE ESE CP Bias SE ESE
Debias_tar 0.965 0.101 0.123 0.138 0.954 0.0857 0.108 0.110 0.953 0.0880 0.110 0.117
Debias_trans_pool 0.906 0.0654 0.0812 0.0735 0.930 0.0543 0.0666 0.0627 0.913 0.0598 0.0704 0.0657
Debias_trans_tar 0.960 0.0966 0.123 0.138 0.961 0.0830 0.103 0.110 0.957 0.0867 0.109 0.117
Debias_trans 0.948 0.0625 0.0755 0.0804 0.960 0.0522 0.0642 0.0677 0.948 0.0572 0.0683 0.0710
Debias_dl 0.936 0.0865 0.109 0.104 0.939 0.0695 0.0874 0.0876 0.936 0.0756 0.0942 0.0910
Figure 3: Inference performance of various methods, including boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared to $\mathcal{N}(0,1)$ with black dotted line.
From Table 1 and Figure 3, we highlight three points. Firstly, our method Debias_trans achieves the better coverage performance, with coverage probabilities closest to $0.95$ among all methods. Because Debias_tar, Debias_trans_tar and Debias_dl only debias on the target study, their coverage probabilities are somewhat a little higher/lower than $0.95$. Moreover, Debias_trans exhibits smaller estimation bias and shorter confidence intervals than these target-only methods, consistent with our theory that incorporating informative sources yields faster asymptotic rates. Secondly, the Debias_trans_pool approach exhibits the poorest coverage performance. In principle, Debias_trans_pool should enjoy reduced bias and tighter intervals due to its larger effective sample size (1000 more observations than Debias_trans in our simulation setting). However, residual shift undermines these potential gains: both its bias and interval lengths are only comparable to those of Debias_trans, while its coverage is noticeably lower. A likely explanation is the downward bias in its estimated standard errors. When Debias_trans_pool includes source studies with large residual shift, the individual density $f_{i}^{(k)}$ for $k\in\mathcal{C}_{h_{2}}^{c}$ is very small, which causes instability in estimation (often upward empirically). Since the standard error estimator has $\widehat{f}_{i}^{(k)}$ in the denominator, this results in underestimation of the variance. Lastly, comparing the density plots with the standard normal curve, the normalized estimates from most methods align well with the standard normal distribution. In contrast, Debias_trans_pool shows a slight but obvious deviation, further indicating the abnormality induced by pooling source studies under residual shift.
6 Application
The proposed transfer learning algorithm is applied to the Genotype-Tissue Expression (GTEx) data, available at https://gtexportal.org/. This dataset contains gene expression levels from 49 tissues, involving 838 individuals, and includes 1,207,976 observations for 38,187 genes. Following the previous work [li2022transfer, zhang2022transfer, zhang2024concert], our study examines gene regulation within the central nervous system (CNS) across various tissues. The CNS-related genes are grouped under MODULE_137, which includes 545 genes, along with an additional 1,632 genes that show significant enrichment in the same experiments as those in the module. Please refer to https://www.gsea-msigdb.org/gsea/msigdb/cards/MODULE_137.html for a detailed description of this module.
Our analysis focuses on the lower-quantile expression levels of the genes JAM2 and SH2D2A. Specifically, we aim to predict their expression at $\tau=0.2$ within a given target tissue by leveraging the expression of other central nervous system (CNS) genes. Both JAM2 and SH2D2A are known to play important roles in cellular functions and disease progression, particularly in cancer and immune response. We consider 11 brain tissues as separate target studies, estimating the model for each tissue individually. For each tissue, the data are divided into five folds: one fold is used as the prediction set, while the remaining four folds are used for training. Details of the dataset, gene functions, and data analysis procedures are provided in the supplementary materials. Figure 4 reports the relative prediction errors of Pooling, TransQR_2step, and our proposed method TransQR, all compared with the baseline L1-QR. The black horizontal line marks a ratio of 1. Consistent with the simulation results, the Pooling approach achieves the least improvement across all tissues for both genes, indicating the presence of non-transferable tissues whose inclusion may deteriorate prediction performance. By contrast, TransQR achieves the lowest prediction errors in almost every tissue for both genes (except C.hemisphere and N.C.A.B.ganglia for JAM2), demonstrating that our transferable set detection can effectively identify informative studies under various types of distribution shift. The two-step transfer method, TransQR_2step, generally lies between Pooling and TransQR, further underscoring that it is not sufficient to account only for parameter shift; a robust and efficient transfer framework must also address covariate shift and residual shift. Compared with mean regression results reported in [li2022transfer, zhao2023residual], our findings provide richer insights into the full distribution of gene expression, particularly at the lower quantile.
Figure 4: Relative prediction errors for the gene expression levels of JAM2 and SH2D2A at $\tau=0.2$.
7 Discussion
There are several avenues to generalize our method. Firstly, we can construct source weights proportional to their respective $\tau$-th densities to achieve sharper results, since leveraging the density at objective quantiles to enhance efficiency is a common approach in the weighted quantile regression [jiang2012oracle]. Also, we can integrate source information at multiple quantiles to predict the target study’s objective quantile coefficient [firpo2022gmm].
Second, we can generalize the all-in-or-all-out source selection manner to select partial informative samples from each source study, especially when the density is strongly dependent on the covariates. To be specific, define the transferable dataset (not transferable set) $\mathcal{C}^{\prime}_{h}=\{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)}):1\leq k\leq K,\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},1/{f^{(k)}_{i}}\leq h_{2}/{f^{(0)}_{i}}\}$, then we can use the data in $\mathcal{C}^{\prime}_{h}$ to leverage more information from related source studies.
Third, our algorithm can be easily generalized to the communication-efficient version with multiple data storage nodes [chen2020distributed]. By carefully observing our Algorithm 1, almost every step is based on the local data and the summary statistics from other studies. We just need to transmit the estimators like $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ between studies, and conduct distributed high-dimensional quantile regression [chen2020distributed].
Lastly, distributionally robust optimization [gao2023distributionally] aims to obtain an optimized target distribution within a feasible distribution set. We can generalize this idea to transfer learning [xiong2023distributionally], which may help us to select source domains more effectively and learn a robust target model under distribution shift. We leave these directions for future work.
8 Supplementary Materials
Outline
This appendix is organized into four subsections. Section 8.1 offers additional explanations for the main text, including the reasonability of our detection rules under covariate shift, the algorithm of the constrained $\ell_{1}$-minimization problem (3), the practical guide for selecting hyperparameters and the discussion of similarity conditions for statistical inference. Section 8.2 presents extensive simulation results under various settings. Section 8.3 provides details on the real data application, including the dataset description, the biological context and the implementation procedure.
8.1 Additional Explanations for the Main Text
8.1.1 Reasonability of Detection Rules under Covariate Shift
In Section 2.1, we introduced the basic idea of excluding source studies with small $\tau$-th densities under residual shift using a simplified setting of low-dimensional linear/quantile regression with identical covariance matrices. However, one may naturally question whether our proposed definition of an oracle transferable set remains valid under covariate shift—that is, when covariance matrices differ across studies. To address this concern clearly, we now present a more sophisticated example to illustrate the underlying rationale.
Compared to the low-dimensional linear setting in Section 2.1, here we assume different covariance matrices $\Sigma_{0}:=\mathbb{E}[\boldsymbol{x}_{i}^{(0)}(\boldsymbol{x}_{i}^{(0)})^{\top}],\Sigma_{1}:=\mathbb{E}[\boldsymbol{x}_{i}^{(1)}(\boldsymbol{x}_{i}^{(1)})^{\top}]$ for the target and source study. Based on the knowledge of linear models, the target and pooling OLS estimator $\widehat{\boldsymbol{\beta}}_{tar},\widehat{\boldsymbol{\beta}}_{pool}$ satisfy:
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\sigma_{0}^{2}}{n_{0}}\Sigma_{0}^{-1}\right),\qquad\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\bar{\Sigma}^{-1}\bar{\Sigma}_{\sigma}\bar{\Sigma}^{-1}}{n_{0}+n_{1}}\right),$
where $\bar{\Sigma}=(n_{0}\Sigma_{0}+n_{1}\Sigma_{1})/(n_{0}+n_{1}),\bar{\Sigma}_{\sigma}=(n_{0}\sigma_{0}^{2}\Sigma_{0}+n_{1}\sigma_{1}^{2}\Sigma_{1})/(n_{0}+n_{1})$. Suppose we are interested in the asymptotic variance of the $j$-th component (using symbol $\mathrm{Avar}_{j}[\cdot]$), we have that
$\mathrm{Avar}_{j}[\widehat{\boldsymbol{\beta}}_{tar}]=\frac{\sigma_{0}^{2}}{n_{0}}(\Sigma_{0}^{-1})_{jj},\qquad\mathrm{Avar}_{j}[\widehat{\boldsymbol{\beta}}_{pool}]=\frac{\sigma_{0}^{2}}{n_{0}+n_{1}}(\bar{\Sigma}^{-1})_{jj}+\frac{n_{1}(\sigma_{1}^{2}-\sigma_{0}^{2})}{(n_{0}+n_{1})^{2}}(\bar{\Sigma}^{-1}\Sigma_{1}\bar{\Sigma}^{-1})_{jj}.$
If the residual variance $\sigma_{1}^{2}$ is much larger than $\sigma_{0}^{2}$, for example when $\sigma_{1}^{2}\geq\sigma_{0}^{2}+\frac{\sigma_{0}^{2}(n_{0}+n_{1})^{2}(\Sigma_{0}^{-1})_{jj}}{n_{0}n_{1}(\bar{\Sigma}^{-1}\Sigma_{1}\bar{\Sigma}^{-1})_{jj}}$, the pooling estimator will be much less efficient compared to the target one. Moving to the quantile regression, we can similarly get that
$\left(\widehat{\boldsymbol{\beta}}_{tar}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\tau(1-\tau)}{n_{0}(f^{(0)}(0))^{2}}\Sigma_{0}^{-1}\right),\quad\left(\widehat{\boldsymbol{\beta}}_{pool}-\boldsymbol{\beta^{*}}\right)\xrightarrow{d}\mathcal{N}\left(0,\frac{\tau(1-\tau)}{n_{0}+n_{1}}\Sigma_{f}^{-1}\bar{\Sigma}\Sigma_{f}^{-1}\right),$
$\mathrm{Avar}_{j}[\widehat{\boldsymbol{\beta}}_{tar}]=\frac{\tau(1-\tau)}{n_{0}(f^{(0)}(0))^{2}}(\Sigma_{0}^{-1})_{jj},$
$\mathrm{Avar}_{j}[\widehat{\boldsymbol{\beta}}_{pool}]=\frac{\tau(1-\tau)}{f^{(0)}(0)}\left[\frac{(\bar{\Sigma}_{f}^{-1})_{jj}}{n_{0}+n_{1}}+\frac{n_{1}(f^{(0)}(0)-f^{(1)}(0))}{(n_{0}+n_{1})^{2}}(\bar{\Sigma}_{f}^{-1}\Sigma_{1}\bar{\Sigma}_{f}^{-1})_{jj}\right],$
where $\bar{\Sigma}_{f}=(n_{0}f^{(0)}(0)\Sigma_{0}+n_{1}f^{(1)}(0)\Sigma_{1})/(n_{0}+n_{1})$. If $f^{(1)}(0)\leq f^{(0)}(0)-\frac{(n_{0}+n_{1})^{2}(\Sigma_{0}^{-1})_{jj}}{f^{(0)}(0)n_{0}n_{1}(\bar{\Sigma}_{f}^{-1}\Sigma_{1}\bar{\Sigma}_{f}^{-1})_{jj}}$, then again the pooling estimator will lose efficiency.
From the above derivation, we also notice that the asymptotic variance is related to the composite matrices $\bar{\Sigma}$, $\bar{\Sigma}_{\sigma}$, and $\bar{\Sigma}_{f}$, which are weighted sums (rather than differences) of the covariance matrices $\Sigma_{0}$ and $\Sigma_{1}$. By utilizing the matrix Loewner order, we have that
$\Sigma_{1}\preceq\frac{n_{0}+n_{1}}{n_{1}}\bar{\Sigma},\quad\bar{\Sigma}^{-1}\preceq\frac{n_{0}}{n_{0}+n_{1}}\Sigma_{0}^{-1}+\frac{n_{1}}{n_{0}+n_{1}}\Sigma_{1}^{-1},$
$\Sigma_{1}\preceq\frac{n_{0}+n_{1}}{n_{1}f^{(1)}(0)}\bar{\Sigma}_{f},\quad\bar{\Sigma}_{f}^{-1}\preceq\frac{n_{0}f^{(0)}(0)}{n_{0}+n_{1}}\Sigma_{0}^{-1}+\frac{n_{1}f^{(1)}(0)}{n_{0}+n_{1}}\Sigma_{1}^{-1}.$
We can then get that
$(\bar{\Sigma}^{-1})_{jj}\leq\frac{n_{0}}{n_{0}+n_{1}}(\Sigma_{0}^{-1})_{jj}+\frac{n_{1}}{n_{0}+n_{1}}(\Sigma_{1}^{-1})_{jj},$
$(\bar{\Sigma}^{-1}\Sigma_{1}\bar{\Sigma}^{-1})_{jj}\leq\frac{n_{0}}{n_{1}}(\Sigma_{0}^{-1})_{jj}+(\Sigma_{1}^{-1})_{jj},$
$(\bar{\Sigma}_{f}^{-1}\Sigma_{1}\bar{\Sigma}_{f}^{-1})_{jj}\leq\frac{n_{0}f^{(0)}(0)}{n_{1}f^{(1)}(0)}(\Sigma_{0}^{-1})_{jj}+(\Sigma_{1}^{-1})_{jj}.$
From these expressions, we can see that for both OLS and QR pooling estimators, the component-wise asymptotic variance can be well controlled by the corresponding diagonal element of matrices $\Sigma_{0}^{-1}$ and $\Sigma_{1}^{-1}$, which is not overly sensitive to the difference of covariance matrices $(\Sigma_{1}-\Sigma_{0})$. Instead, the primary drivers affecting estimation performance are the residual shift, i.e., $\sigma_{1}^{2}/\sigma_{0}^{2}$ and $f^{(1)}(0)/f^{(0)}(0)$. Moreover, as the source sample size $n_{1}$ increases, the condition required to improve estimation efficiency becomes less restrictive. Consequently, our detection rule $n_{0}\mathbb{E}[f_{i}^{(0)}]\leq h_{2}n_{k}\mathbb{E}[f_{i}^{(k)}]$ remains reasonable and robust under covariate shift scenarios.
To further validate our detection rule under strong covariate shift, we conduct additional simulations following the setting of [he2024transfusion]. The covariates of each source study now follow $\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}}+\boldsymbol{A}_{\boldsymbol{x}}^{(k)\top}\boldsymbol{A}_{\boldsymbol{x}}^{(k)})$, where $\boldsymbol{A}_{\boldsymbol{x}}^{(k)}$ is a random matrix whose entries are $0.3$ with probability $0.3$ and 0 with probability $0.7$. The simulation results are shown in Figure 10, 11, 12, 16 and 17. All figures convey similar conclusions to those in the main text, demonstrating the robustness of our proposed methods under strong covariate shift. Specifically, our method performs almost as well as the oracle benchmark in the parameter estimation, whereas the two-step approach exhibits noticeably higher $\ell_{2}$-error, underscoring its vulnerability to heterogeneous covariance structures.
8.1.2 Algorithm of the Constrained L1-minimization Problem
In Section 2.1, we introduce the transfer framework (3), which can simultaneously tackle the parameter/residual/residual shift when utilizing the information from source studies. Here we introduce the iterative algorithm to solve the constrained $\ell_{1}$-minimization problem (3) when the oracle transferable set $\mathcal{C}_{h}$ is known, which is also applied in [li2023estimation]. We can approximate (3) with a joint $\ell_{1}$-penalized quantile regression ($\ell_{1}$-QR) form, leveraging the equivalence between Lasso and Dantzig selectors [bickel2009simultaneous]:
$\left(\widehat{\boldsymbol{\beta}},\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}\right)=\underset{\boldsymbol{\beta},\{\boldsymbol{\delta}^{(k)}\}_{k\in\mathcal{C}_{h}}}{\arg\min}\left\{\sum_{k\in\{0\}\cup\mathcal{C}_{h}}L^{(k)}(\boldsymbol{\beta}+\boldsymbol{\delta}^{(k)})+\lambda_{\boldsymbol{\beta}}\|\boldsymbol{\beta}\|_{1}+\sum_{k\in\mathcal{C}_{h}}\lambda_{k}\|\boldsymbol{\delta}^{(k)}\|_{1}\right\}.$ (15)
This formulation is similar to the co-training step in [he2024transfusion], with replacing $\boldsymbol{\delta}^{(k)}$ by $\boldsymbol{w}^{(k)}-\boldsymbol{\beta}$. As suggested by [li2023estimation], the algorithm for solving (15) proceeds naturally using iterative alternating direction methods. At a high level, our method works as follows. We first obtain an initial estimator $\widehat{\boldsymbol{\beta}}_{0}$ by $\ell_{1}$-QR solely on the target data (or the estimator $\widehat{\boldsymbol{\beta}}_{t-1}$ from the last iteration). Then for each source $k\in\mathcal{C}_{h}$ we subtract $(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}_{t-1}$ from the response $y_{i}^{(k)}$ and obtain the estimator $\widehat{\boldsymbol{\delta}}^{(k)}_{t}$ by $\ell_{1}$-QR on the refined source dataset. After that, we take $\{\widehat{\boldsymbol{\delta}}^{(k)}_{t}\}_{k\in\mathcal{C}_{h}}$ into the objective function (15), solve for the estimator $\widehat{\boldsymbol{\beta}}_{t}$, and begin the next round. The iteration stops when estimators converge, where the detailed procedure is summarized in Algorithm 3.
Input: Target data $\{((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}}\}$, source data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\mathcal{C}_{h}}$,   tuning parameters $\lambda_{\boldsymbol{\beta}},\{\lambda_{k}\}_{k\in\{0\}\cup\mathcal{C}_{h}}$.
Compute the initial estimate $\widehat{\boldsymbol{\beta}}_{0}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(0)})^{\top},y_{i}^{(0)})_{i=1}^{n_{0}},\lambda_{0}]$.
for $t=1,2,\dots$ do
  1. 1.
    Compute $\widehat{\boldsymbol{\delta}}^{(k)}_{t}\leftarrow\ell_{1}\text{-QR }\text{ on }[((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}\widehat{\boldsymbol{\beta}}_{t-1})_{i=1}^{n_{k}},\lambda_{k}],k\in\mathcal{C}_{h}$.
  2. 2.
    Solve the pooled $\ell_{1}$-QR problem
    $\widehat{\boldsymbol{\beta}}_{t}=\underset{\boldsymbol{\beta}}{\arg\min}\left\{\sum_{k\in\{0\}\cup\mathcal{C}_{h}}L^{(k)}(\boldsymbol{\beta}+\widehat{\boldsymbol{\delta}}^{(k)}_{t})+\lambda_{\boldsymbol{\beta}}\left\|\boldsymbol{\beta}\right\|_{1}\right\}.$
  3. 3.
    When $\widehat{\boldsymbol{\beta}}_{t}$ converges at step $T$, record the present result and exit the iteration.
Output: $\{\widehat{\boldsymbol{\beta}},\{\widehat{\boldsymbol{\delta}}^{(k)}\}_{k\in\mathcal{C}_{h}}\}:=\{\widehat{\boldsymbol{\beta}}_{T},\{\widehat{\boldsymbol{\delta}}^{(k)}_{T}\}_{k\in\mathcal{C}_{h}}\}$.
Algorithm 3 Parameter estimation with oracle transferable set $\mathcal{C}_{h}$
8.1.3 A Practical Guide for Tuning Parameters
There are several hyperparameters in our transfer procedures, including detection thresholds $t_{1},t_{2}$, tuning parameters $\{\lambda_{k}\}_{k=0}^{K},\lambda_{\boldsymbol{\beta}}$, bandwidth parameters $\{b_{k}\}_{k=0}^{K}$ in Algorithm 1, $\{\lambda_{\tau}^{(k)},\lambda_{\boldsymbol{\theta}}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}},\bar{\lambda}$ in Algorithm 2. To attain better performance in the simulations, here we propose several ways to determine these hyperparameters.
  • To select the $\ell_{1}$-QR tuning parameters, we deviate from the vanilla penalized quantile regression and instead follow the $\lambda$-construction idea in [belloni2011l1, giessing2023debiased]. Specifically, the regularization term changes from $\lambda\|\boldsymbol{\beta}\|_{1}$ to $\lambda\sqrt{\tau(1-\tau)}\sum_{j=1}^{p}\widehat{\sigma}_{k,j}|\beta_{j}|$, with $\widehat{\sigma}_{k,j}^{2}:=n_{k}^{-1}\sum_{i=1}^{n_{k}}(\boldsymbol{x}_{ij}^{(k)})^{2}$ and $\lambda=c^{*}\cdot\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$, where $\Lambda_{\tau}^{(k)}(\alpha^{*}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}})$ is the $(1-\alpha^{*})$-th quantile of $\Lambda_{\tau}^{(k)}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$ and
    $\Lambda_{\tau}^{(k)}|\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}:=\max_{1\leq j\leq p}\left|\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\frac{\left(\tau-1\left\{U_{i}\leq\tau\right\}\right)\boldsymbol{x}_{ij}^{(k)}}{\sqrt{\tau(1-\tau)\widehat{\sigma}_{k,j}^{2}}}\right|,$
    with $U_{1},\ldots,U_{n_{k}}$ are i.i.d. Uniform(0,1)(0,1) random variables, independent of the covariates $\{\boldsymbol{x}_{i}^{(k)}\}_{i=1}^{n_{k}}$. Such a pilot construction can save the computation cost a lot compared to the cross-validation. $\alpha^{*}$ is set as 0.05 uniformly, while we set $c^{*}=1$ for $\lambda_{0},\lambda_{\boldsymbol{\beta^{*}}}$, $c^{*}=1.5$ for $\{\lambda_{k}\}_{k=1}^{K}$, $c^{*}=2\sqrt{\tau(1-\tau)}$ for $\{\lambda_{\tau}^{(k)}\}_{k\in\{0\}\cup\widehat{\mathcal{C}}}$. This procedure can be executed through R functions LassoLambdaHat() in the quantreg package.
  • For the bandwidth parameters $\{b_{k}\}_{k=0}^{K}$, we follow the Powell bandwidth construction [dai2021inference, huang2023estimation] with estimators $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ in Algorithm 1. Specifically,
    $b_{k}=\left(\Phi^{-1}(\tau+\widetilde{b}_{k})-\Phi^{-1}(\tau-\widetilde{b}_{k})\right)\min\left\{\sqrt{\widehat{var}\left[\widehat{\boldsymbol{e}}^{(k)}_{i}\right]},\frac{\widehat{Q}_{0.75}\left[\widehat{\boldsymbol{e}}^{(k)}_{i}\right]-\widehat{Q}_{0.25}\left[\widehat{\boldsymbol{e}}^{(k)}_{i}\right]}{1.34}\right\},$
    $\widetilde{b}_{k}=n_{k}^{-1/3}\left\{\Phi^{-1}(0.975)\right\}^{2/3}\left(\frac{1.5\phi^{2}(\Phi^{-1}(\tau))}{2\Phi^{-2}(\tau)+1}\right)^{1/3},$
    where $\widehat{\boldsymbol{e}}^{(k)}_{i}=y_{i}^{(k)}-(\boldsymbol{x}_{i}^{(k)})^{\top}(\widetilde{\boldsymbol{\beta}}+\widetilde{\boldsymbol{\delta}}^{(k)})$, $\widehat{var}[\widehat{\boldsymbol{e}}^{(k)}_{i}$] and $\widehat{Q}_{\tau}[\widehat{\boldsymbol{e}}^{(k)}_{i}$] denote the sample variance and the $\tau$-th sample quantile of $\{\widehat{\boldsymbol{e}}^{(k)}_{i}\}_{i=1}^{n_{k}}$.
  • For the threshold parameter $t_{1},t_{2}$ in Algorithm 1, one practical strategy is to conduct the cross-validation procedure [tian2022transfer, zhang2022transfer] for a target study and several candidate source studies, where we evaluate a grid of threshold values and choose the thresholds that yield the best prediction performance (least quantile loss on the validation set). The chosen thresholds can then be applied across all related transfer learning tasks. For moderately complex scenarios, this procedure is not computationally expensive and provides sufficiently accurate threshold choices. To further examine the impact of threshold selection, we conducted sensitivity analyses by varying the values of $t_{1}$ and $t_{2}$ in the simulations, as shown in Figures 13 and 14. The results indicate that as long as the thresholds lie within a reasonable interval, our method’s performance remains close to the oracle benchmark, consistent with the detection consistency established in Theorem 3. For all simulations in this paper, we set $t_{1}=5$ and $t_{2}=0.3$.
  • For the post-selected threshold $\bar{\lambda}$, we directly let it be 0.01 for simplicity.
  • For the parameters $\lambda_{\boldsymbol{\theta}}^{(k)}$ in the projection, we use the common pivotal choice $c_{\boldsymbol{\theta}}\sqrt{n_{k}\log p}$ in the high-dimensional settings. For different simulations (with settings described in Section 8.2), the constant $c_{\boldsymbol{\theta}}$ is set as $0.1$ for the target study and source studies under normal distributions without strong covariate shift, $0.05$ for source studies under normal distributions with strong covariate shift, $0.08$ for the target study and source studies under $t$ distributions without strong covariate shift, $0.03$ for source studies under $t$ distributions with strong covariate shift.
8.1.4 Quality Control for Statistical Inference
In Section 4, we introduce the Neyman orthogonal debiased approach to incorporate informative source studies for statistical inference. when the source studies in the oracle transferable set $\mathcal{C}_{h}$ satisfy the requirement for Theorem 4, we can get the asymptotic result of Algorithm 2 by combining Theorem 3 and 4.
Corollary 2.
Assume all conditions required by Theorem 1, 3 and 4 hold for the oracle transferable set $\mathcal{C}_{h}$. For the estimator $\widehat{\alpha}$ obtained from Algorithm 2, we have
$\sigma_{\mathcal{C}_{h}}^{-1}\sqrt{n_{\mathcal{C}_{h}}}(\widehat{\alpha}-\beta_{1}^{*})=\mathbb{U}_{\mathcal{C}_{h}}(\tau)+o_{p}(1),\hskip 18.49988pt\mathbb{U}_{\mathcal{C}_{h}}(\tau)\leadsto\mathcal{N}(0,1),$
where $\sigma_{\mathcal{C}_{h}},\mathbb{U}_{\mathcal{C}_{h}}$ follow the definitions of $\sigma_{\mathcal{C}},\mathbb{U}_{\mathcal{C}}$ in Theorem 4 with $\mathcal{C}=\mathcal{C}_{h}$. Moreover, the estimator $\widehat{\sigma}_{\widehat{\mathcal{C}}}^{2}=\tau(1-\tau)n_{\widehat{\mathcal{C}}}[\sum_{k\in\{0\}\cup\widehat{\mathcal{C}}}\sum_{i=1}^{n_{k}}(\widehat{v}_{i}^{(k)})^{2}]^{-1}$ is consistent.
Although this corollary seems to provide the theoretical guarantee for the post-detection inference procedure, leveraging source data still needs parameter shift small enough, especially for the approximation error $r_{i}^{(k)}=x_{i,1}^{(k)}\delta^{(k)}_{1},k\in\mathcal{C}_{h}$. In fact, none of the existing debiased approaches can be applied to the approximate error term, because each $\delta^{(k)}_{1}$ shares the same covariate $x_{i,1}^{(k)}$ with the target parameter $\beta^{*}_{1}$. As noted by [belloni2019valid], the orthogonal approach remains theoretically valid as long as $\mathbb{E}[(r_{i}^{(k)})^{2}]=(\mathbb{E}[x_{i1}^{(k)}]\delta_{1}^{(k)})^{2}$ stays below a certain threshold. In the transfer setting, introducing more external data increases the requirement for the contrast $\delta_{1}^{(k)}$, where Condition 7 showed that a sufficient theoretical bound is $o(n_{\widehat{\mathcal{C}}}^{-1/2})$. This requirement is understandable since we need smaller parameter contrast for better normality results. However, constructing consistent selection procedures or hypothesis tests for this condition is nearly impossible, especially when dynamically selecting $\widehat{\mathcal{C}}$. Although there is no theoretical guarantee, we propose a practical quality control procedure to validate the transfer debiased estimator. We can use prior knowledge to exclude certain source studies with non-negligible $\widetilde{\delta}_{1}^{(k)}$ and perform a bootstrap test to check the normality of the debiased estimator. If the normality hypothesis is rejected, we revert to debiasing only the target study, yielding a result similar to previous approaches. The full procedure is outlined in Algorithm 4.
Input: Same as Algorithm 2.
  1. A)
    Pre-control: for each source study $k\in\widehat{\mathcal{C}}$, if there has prior knowledge that the parameter contrast $\delta_{1}^{(k)}$ is non-negligible, or $|\widetilde{\delta}_{1}^{(k)}|$ is larger than a prespecified constant, then we remove this study. Denote the new detected set by $\widetilde{\mathcal{C}}$.
  2. B)
    Post-control: normality test with quantile bootstrap.
    1. (1)
      Conduct the quantile pair bootstrap [hahn1995bootstrapping] $R$ times on the data $\{((\boldsymbol{x}_{i}^{(k)})^{\top},y_{i}^{(k)})_{i=1}^{n_{k}}\}_{k\in\{0\}\cup\widetilde{\mathcal{C}}}$ to obtain the debiased estimates $\{\widehat{\alpha}_{r}\}_{r=1}^{R}$.
    2. (2)
      Check the normality of $\{\widehat{\alpha}_{r}\}_{r=1}^{R}$ by Shapiro-Wilk test [shapiro1965analysis].
- If rejecting null hypothesis ($p\text{-value}<0.05$), we choose to only debias on the target study by $\arg\min_{\alpha\in\mathcal{A}_{0}}|\sum_{i=1}^{n_{0}}\widehat{\psi}_{i}^{(0)}(\alpha)|$, where the search region $\mathcal{A}_{0}=\{\alpha\in\mathbb{R}:|\alpha-\widehat{\beta}_{1}|\leq 10[n_{0}^{-1}\sum_{i=1}^{n_{0}}(x^{(0)}_{i,1})^{2}]^{-1/2}/\log n_{0}\}$, which is exactly the approach in [belloni2019valid].
- If not rejecting null hypothesis, conduct Algorithm 2 with $\widetilde{\mathcal{C}}$. Output: Debiased estimate $\widehat{\alpha}$.
Algorithm 4 Quality control for Algorithm 2
In Section 5.2, we have shown that transfer debiased estimators can enjoy sharper inference results if the contrast level $\delta_{1}^{(k)}$ is small. Now with other settings unchanged, we let $|\delta_{1}^{(k)}|=0.1$ and still compare the methods conducted in 5.2, where the results are shown in Table 2 and Figure 15.
From Table 2 and Figure 15, when $|\delta_{1}^{(k)}|$ is large, unfortunately our Debias_trans performs much worse than the target debiasing methods, which is in our expectation since Condition 7 is no longer satisfied. The density plots in Figure 15 show that the transfer debiased estimate is no longer normal, and the Shapiro-Wilk test rejects the normality hypothesis. In a similar way, the bootstrap approach in Algorithm 4 can also indicate us that debiasing on source studies will generate unreliable estimates. Therefore, we recommend only debiasing on the target study, namely the Debias_trans_tar approach, which can keep the coverage proportions with more stable interval lengths than Debias_dl.
8.2 Additional Simulation Results
We first introduce the simulation details for the motivating example in Section 1. The setting is similar to the simulation in [zhang2022transfer]. We set $p=500,s=10,n_{0}=200,\boldsymbol{\beta^{*}}=(\mathbf{0.5}_{s},\mathbf{0}_{p-s})^{\top},\boldsymbol{x}_{i}^{(0)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$ with $\boldsymbol{\Sigma}_{\boldsymbol{x}}=(0.7^{|i-j|})_{1\leq i,j\leq p}$. We consider only one source study with sample size $n_{1}$ varying from 100 to 500, and we assume no parameter or covariate shift, i.e., $\boldsymbol{\delta}^{(1)}=0,\boldsymbol{x}_{i}^{(1)}\sim\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}})$. The key change lies in the residual distributions, where $\epsilon_{i}^{(0)}$ is standard normal, but $\epsilon_{i}^{(1)}$ follows one of the four types of distributions: 1) standard normal: $\mathcal{N}(0,1)$; 2) Cauchy distribution: $\mathcal{C}(0,3)$; 3) mixed Gaussian: $z\mathcal{N}(-3,0.5)+(1-z)\mathcal{N}(3,0.5)$, $z\sim\mathcal{B}ernoulli(\tau)$; 4) noisy normal: $\mathcal{N}(0,5^{2})$. It is worth noting that in this setting, the source study is included in the oracle transferable set by the previous definition [li2022transfer, tian2022transfer]. We compare four methods for estimating the target parameter $\boldsymbol{\beta^{*}}$ at the $0.2$-th quantile, which are: 1) L1-QR: $\ell_{1}$-penalized quantile regression [belloni2011l1] on the target data; 2) Pooling: $\ell_{1}$-penalized quantile regression on the entire dataset; 3) TransQR_2step: two-step transfer for quantile regression [huang2023estimation]; 4) TransSQR_2step: two-step transfer for smoothed quantile regression [zhang2022transfer]. We repeat these methods 50 times and report the average of $\ell_{2}$-estimation errors in Figure 1.
We conduct extensive experiments for parameter estimation and statistical inference under several circumstances. In particular, we consider the cases with strong covariate shift and $t$-distribution. Strong covariate shift means the covariates of each source study now follow $\mathcal{N}(\mathbf{0}_{p},\boldsymbol{\Sigma}_{\boldsymbol{x}}+\boldsymbol{A}_{\boldsymbol{x}}^{(k)\top}\boldsymbol{A}_{\boldsymbol{x}}^{(k)})$, where $\boldsymbol{A}_{\boldsymbol{x}}^{(k)}$ is a random matrix whose entries are $0.3$ with probability $0.3$ and 0 with probability $0.7$. $t$-distribution means we set $\epsilon_{i}^{(k)},k\in\{0\}\cup\mathcal{C}_{h_{2}}$ follow the shifted $t(3)$ distribution. Except the changes we mentioned, all other simulation settings are kept the same as those used in Figure 2 and 3 of the main manuscript.
Parameter estimation
  • Figure 5: heteroscedastic model, $K=5$, $h_{1}=5$.
  • Figure 6: homoscedastic model, $K=20$, $h_{1}=5$.
  • Figure 7: heteroscedastic model, $K=20$, $h_{1}=5$.
  • Figure 8: homoscedastic model, $K=5$, $h_{1}=15$.
  • Figure 9: heteroscedastic model, $K=5$, $h_{1}=15$.
  • Figure 10: homoscedastic model, $K=5$, $h_{1}=5$, strong covariate shift.
  • Figure 11: homoscedastic model, $K=5$, $h_{1}=5$, strong covariate shift, $t$-distribution.
  • Figure 12: heteroscedastic model, $K=5$, $h_{1}=5$, strong covariate shift, $t$-distribution.
  • Figure 13: homoscedastic model with varying $t_{1}$ values for TransQR. We set $\mathcal{C}_{h_{1}}=\{1,2\},\mathcal{C}_{h_{2}}=\{1,2,3,4,5\}$. We use five-fold cross-validation on the target study to compute the prediction error (quantile loss). The results are averaged over $20$ repetitions.
  • Figure 14: homoscedastic model with varying $t_{2}$ values for TransQR. We set $\mathcal{C}_{h_{1}}=\{1,2,3,4,5\},\mathcal{C}_{h_{2}}=\{1,2\}$. We use five-fold cross-validation on the target study to compute the prediction error (quantile loss). The results are averaged over $20$ repetitions.
Statistical inference
  • Figure 15 and Table 2: $|\delta_{1}^{(k)}|=0.1$.
  • Figure 16 and Table 3: strong covariate shift.
  • Figure 17 and Table 4: strong covariate shift.
Figure 5: Average $\ell_{2}$-error of various methods under heteroscedastic model with $K=5$.
Figure 6: Average $\ell_{2}$-error of various methods under homoscedastic model with $K=20$.
Figure 7: Average $\ell_{2}$-error of various methods under heteroscedastic model with $K=20$.
Figure 8: Average $\ell_{2}$-error of various methods under homoscedastic model with $h_{1}=15$.
Figure 9: Average $\ell_{2}$-error of various methods under heteroscedastic model with $h_{1}=15$.
Figure 10: Average $\ell_{2}$-error of various methods under homoscedastic model with strong covariate shift.
Figure 11: Average $\ell_{2}$-error of various methods under homoscedastic model and $t$-distribution with strong covariate shift.
Figure 12: Average $\ell_{2}$-error of various methods under heteroscedastic model and $t$-distribution with strong covariate shift.
Figure 13: Average prediction errors under different values of the threshold $t_{1}$ in TransQR.
Figure 14: Average prediction errors under different values of the threshold $t_{2}$ in TransQR.
Table 2: Coverage probability of $95\%$ confidence intervals (CP), average absolute bias (Bias), standard error (SE), and average estimated standard error (ESE) of various methods with $|\delta^{(k)}_{1}|=0.1$.
Method $\tau=0.2$ $\tau=0.5$ $\tau=0.7$
CP Bias SE ESE CP Bias SE ESE CP Bias SE ESE
Debias_tar 0.965 0.101 0.123 0.138 0.954 0.0857 0.108 0.110 0.953 0.0880 0.110 0.117
Debias_trans_pool 0.809 0.0892 0.107 0.0738 0.788 0.0792 0.0955 0.0627 0.781 0.0823 0.0998 0.0657
Debias_trans_tar 0.971 0.0958 0.125 0.138 0.950 0.0844 0.105 0.110 0.954 0.0875 0.110 0.117
Debias_trans 0.866 0.0861 0.104 0.0807 0.836 0.0772 0.0937 0.0677 0.853 0.0791 0.0953 0.0711
Debias_dl 0.928 0.0876 0.109 0.104 0.935 0.0714 0.0893 0.0854 0.936 0.0782 0.0964 0.0910
Figure 15: Inference results of various methods with $|\delta^{(k)}_{1}|=0.1$, including boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared to $\mathcal{N}(0,1)$ with black dotted line.
Table 3: Coverage probability of $95\%$ confidence intervals (CP), average absolute bias (Bias), standard error (SE), and average estimated standard error (ESE) of various methods with strong covariate shift.
Method $\tau=0.2$ $\tau=0.5$ $\tau=0.7$
CP Bias SE ESE CP Bias SE ESE CP Bias SE ESE
Debias_tar 0.963 0.103 0.137 0.138 0.941 0.0853 0.108 0.110 0.954 0.0934 0.114 0.118
Debias_trans_pool 0.927 0.0620 0.0778 0.0696 0.944 0.0556 0.0687 0.0656 0.918 0.0594 0.0774 0.0655
Debias_trans_tar 0.979 0.0964 0.134 0.138 0.952 0.0826 0.105 0.110 0.963 0.0879 0.110 0.118
Debias_trans 0.961 0.0592 0.0755 0.0769 0.963 0.0515 0.0652 0.0710 0.948 0.0565 0.0784 0.0713
Debias_dl 0.966 0.0784 0.0986 0.110 0.945 0.0677 0.0868 0.0884 0.951 0.0724 0.0914 0.0962
Figure 16: Inference results of various methods with strong covariate shift, including boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared to $\mathcal{N}(0,1)$ with black dotted line.
Table 4: Coverage probability of $95\%$ confidence intervals (CP), average absolute bias (Bias), standard error (SE), and average estimated standard error (ESE) of various methods under $t$-distribution with strong covariate shift.
Method $\tau=0.2$ $\tau=0.5$ $\tau=0.7$
CP Bias SE ESE CP Bias SE ESE CP Bias SE ESE
Debias_tar 0.929 0.140 0.166 0.165 0.957 0.100 0.124 0.131 0.926 0.116 0.162 0.137
Debias_trans_pool 0.905 0.0795 0.107 0.0833 0.933 0.0624 0.0802 0.0801 0.906 0.0689 0.0893 0.0770
Debias_trans_tar 0.945 0.128 0.161 0.165 0.966 0.0943 0.119 0.131 0.946 0.107 0.135 0.137
Debias_trans 0.944 0.0753 0.105 0.0935 0.978 0.0588 0.0745 0.0874 0.948 0.0648 0.0836 0.0844
Debias_dl 0.940 0.110 0.135 0.138 0.956 0.0787 0.0996 0.101 0.952 0.0857 0.108 0.113
Figure 17: Inference results of various methods under $t$-distribution with strong covariate shift, including boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared to $\mathcal{N}(0,1)$ with black dotted line.
8.3 Details of Real Data Application
We first give a brief introduction for the two genes studied in Section 6.
  • JAM2 (Junctional Adhesion Molecule 2) is a gene that encodes a protein involved in tight junctions between epithelial and endothelial cells, playing a crucial role in cell adhesion and maintaining barrier functions. It is associated with various cellular processes such as cell proliferation and migration. Studies have shown that low expression of JAM2 is linked to poor prognosis in several cancers, including lung adenocarcinoma (LUAD), where reduced JAM2 levels correlate with tumor progression and metastasis. Investigating its low expression helps in understanding cancer progression and identifying therapeutic targets for limiting tumor growth [dong2024jam2].
  • SH2D2A encodes the T-cell-specific adapter protein (TSAd), which is essential for T-cell signaling and activation. The expression of SH2D2A is tightly regulated at both the transcriptional and translational levels, with cAMP signaling shown to induce its mRNA expression in T cells. However, low levels of SH2D2A protein may affect immune responses. Notably, reduced SH2D2A expression can enhance T-cell-mediated anti-tumor immunity, making it an interesting target in cancer immunotherapy research [berge2012sh2d2a].
Then we provide the complete name for each target issue, which is shown in Table 5.
Table 5: Complete Name for 11 Target Brain Issues in the Application
Tissue Abbreviation
brain amygdala Amygdala
brain hypothalamus Hypothalamus
brain caudate basal ganglia C.B.ganglia
brain cerebellar hemisphere C.hemisphere
brain cerebellum Cerebellum
brain cortex Cortex
brain frontal cortex ba9 F.cortex
brain anterior cingulate cortex ba24 A.C.cortex
brain nucleus accumbens basal ganglia N.C.A.B.ganglia
brain putamen basal ganglia P.B.ganglia
brain spinal cord cervical c-1 S.C.cervical
There are some additional points in our practical implementation:
  • The gene expression data is pre-processed by filtering out genes with constant expression levels and applying standard normalization.
  • We removed the source tissues with sample sizes less than $100$.
  • We performed threshold value searching for the “brain cortex” tissue as the target study in real data applications. The five-fold cross-validation relative prediction errors on the “brain cortex” tissue, shown in Figure 18, guide our choice of $t_{1}=7$ and $t_{2}=0.8$ for all tasks. The other regularization parameters follow the guidance in Section 8.1.3.
  • We observe that source data appears less beneficial at $\tau=0.5$ but seems to be quite helpful at $\tau=0.2$. This suggests that the distribution shift between the target and source studies may vary with the quantile level of model responses.
Figure 18: Transfer performance under different values for the thresholds $t_{1},t_{2}$ for predicting the cortex tissue at quantile level $\tau=0.2$.